### v2

# Rethinking Oversmoothing in Graph Neural Networks: A Rank-Based Perspective

Abstract
Oversmoothing is a fundamental challenge in graph neural networks (GNNs): as the number of layers increases, node embeddings become increasingly similar, and model performance drops sharply. Traditionally, oversmoothing has been quantified using metrics that measure the similarity of neighbouring node features, such as the Dirichlet energy. While these metrics are related to oversmoothing, we argue they have critical limitations and fail to reliably capture oversmoothing in realistic scenarios. For instance, they provide meaningful insights only for very deep networks and under somewhat strict conditions on the norm of network weights and feature representations. As an alternative, we propose measuring oversmoothing by examining the numerical or effective rank of the feature representations. We provide theoretical support for this approach, demonstrating that the numerical rank of feature representations converges to one for a broad family of nonlinear activation functions under the assumption of nonnegative trained weights. To the best of our knowledge, this is the first result that proves the occurrence of oversmoothing without assumptions on the boundedness of the weight matrices. Along with the theoretical findings, we provide extensive numerical evaluation across diverse graph architectures. Our results show that rank-based metrics consistently capture oversmoothing, whereas energy-based metrics often fail. Notably, we reveal that a significant drop in the rank aligns closely with performance degradation, even in scenarios where energy metrics remain unchanged.
Machine Learning, ICML
1 Introduction
Graph neural networks (GNNs) have emerged as a powerful framework for learning representations from graph-structured data, with applications spanning knowledge retrieval and reasoning (Peng et al., 2023; Tian et al., 2022), personalised recommendation systems (Damianou et al., 2024; Peng et al., 2022), social network analysis (Fan et al., 2019), and 3D mesh classification (Shi & Rajkumar, 2020). Central to most GNN architectures is the message-passing paradigm, where node features are iteratively aggregated from their neighbours and transformed using learned functions, such as multi-layer perceptrons or graph-attention mechanisms.
However, the performance of message-passing-based GNNs is known to deteriorate after only a few layers, essentially placing a soft limit on the depth of GNNs. This is often linked to the observation of increasing similarity between learned features as GNNs deepen and is named oversmoothing (Li et al., 2018; Nt & Maehara, 2019).
Over the years, oversmoothing in GNNs, as well as methods to alleviate it, have been studied based on the decay of some node feature similarity metrics, such as the Dirichlet energy and its variants (Oono & Suzuki, 2019; Cai & Wang, 2020; Bodnar et al., 2022; Nguyen et al., 2022; Di Giovanni et al., 2023; Wu et al., 2023; Roth & Liebig, 2023). At a high level, most of these metrics directly measure the norm of the absolute deviation from the dominant eigenspace of the message-passing matrix. In linear GNNs without bias terms, this eigenspace is often known and easily computable via e.g. the power method. However, when nonlinear activation functions or biases are used, the dominant eigenspace may change, causing these oversmoothing metrics to fail and give false negative signals about the oversmoothing state of the learned features.
Therefore, these metrics are often considered as providing sufficient but not necessary evidence for oversmoothing (Rusch et al., 2023a). Despite this, there is a considerable body of literature using these somewhat unreliable metrics as part of their evidence for non-occurrence of oversmoothing in GNNs (Zhou et al., 2021; Chen et al., 2022; Rusch et al., 2022; Wang et al., 2022; Maskey et al., 2023; Nguyen et al., 2023; Rusch et al., 2023b; Epping et al., 2024; Scholkemper et al., 2024; Wang & Cho, 2024).
As we show in Section 6, the performance degradation of GNNs trained on real datasets often happens well before any noticeable decay in these oversmoothing metrics can be observed. Most empirical studies in the literature that observe the decay of the Dirichlet-energy-like metrics are conducted over the layers of the same very deep untrained (with randomly sampled weights) or effectively untrained [^1] Deep networks (with, say, over 100 layers) that are trained but whose loss and accuracy remain far from acceptable. GNNs (Rusch et al., 2022; Wang et al., 2022; Rusch et al., 2023b; Wu et al., 2023), where the decay of the metrics is likely driven by the small weight initializations reducing all features to zero. Instead, we observe that when GNNs of different depths are separately trained, these metrics do not correlate well with their performance degradation.
Furthermore, we note that these metrics can only indicate oversmoothing when their values converge exactly to zero, corresponding to either an exact alignment to the dominant eigenspace or to the feature representation matrix collapsing to the all-zero matrix. This double implication presents an issue: in realistic settings with a large but not excessively large number of layers, we may observe the decay of the oversmoothing metric by, say, two orders of magnitude while still being far from zero. In such cases, it is unclear whether the features are aligning with the dominant eigenspace, simply decreasing in magnitude, or exhibiting neither of the two behaviours. As a result, these types of metrics provide little to no explanation for the degradation of GNN performance.
As an alternative to address these shortcomings, we advocate for the use of a continuous approximation of the rank of the network’s feature representations to measure oversmoothing. Our experimental evaluation across various GNN architectures trained for node classification demonstrates that continuous rank relaxations, such as the numerical rank and the effective rank, correlate strongly with performance degradation in independently trained GNNs—even in settings where popular energy-like metrics show little to no correlation.
Overall, the main contributions of this paper are as follows:
  • We review popular oversmoothing metrics in the current literature and provide a novel perspective on their theoretical analysis based on nonlinear activation eigenvectors.
  • Based on this analysis and observations on the linear case, we propose the rank as a better metric for quantifying oversmoothing, and thereby re-defining oversmoothing in GNNs as the convergence towards a low-rank matrix rather than to a matrix of exactly rank one.
  • We prove the convergence of the numerical rank towards one for linear GNNs and nonlinear GNNs where the eigenvector of the message-passing matrix is also the eigenvector of the nonlinear activation function under the assumption of non-negative weights. To our knowledge, this is the first theoretical result proving that oversmoothing may occur independently of the weights’ magnitude with nonlinear activation functions.
  • We provide extensive numerical evidence that continuous rank relaxation functions better capture oversmoothing than commonly used Dirichlet-like metrics.
2 Background
2.1 Graph Convolutional Network
Let $\mathcal{G}=(\mathcal{V},\mathcal{E})$ be an undirected graph with $\mathcal{V}$ denoting its set of vertices and $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$ its set of edges. Let $\widetilde{A}\in\mathbb{R}^{N\times N}$ be the unweighted adjacency matrix with $N=|\mathcal{V}|$ being the total number of nodes, $|\mathcal{E}|$ being the total number of edges of $\mathcal{G}$ and $A$ the corresponding symmetric adjacency matrix normalized by the node degrees:
$A=\widetilde{D}^{-1/2}\widetilde{A}\widetilde{D}^{-1/2},$ (1)
where $\widetilde{D}=D+I$, $D$ is the diagonal degree matrix of the graph $\mathcal{G}$, and $I$ is the identity matrix. The rows of the feature matrix $X\in\mathbb{R}^{N\times d}$ are the concatenation of the $d$-dimensional feature vectors of all nodes in the graph. At each layer $l$, the GCN updates the node features as follows
$X^{(l+1)}=\sigma(AX^{(l)}W^{(l)})$ (2)
where $\sigma$ is a nonlinear activation function, applied componentwise, and $W^{(l)}$ is a trainable weight matrix.
2.2 Graph Attention Network
While GCNs use a fixed normalized adjacency matrix to perform graph convolutions at each layer, Graph Attention Networks (GATs) (Veličković et al., 2017; Brody et al., 2021) perform graph convolution through a layer-dependent message-passing matrix $A^{(l)}$ learned through an attention mechanism as follows
$A^{(l)}_{ij}=\text{softmax}_{j}(\sigma_{a}(p_{1}^{(l)\top}W^{(l)\top}X_{i}+p_{2}^{(l)\top}W^{(l)\top}X_{j}))$ (3)
where $p_{i}^{(l)}$ are learnable parameter vectors, $X_{i},X_{j}$ denote the feature of the $i$th and $j$th nodes respectively, the activation $\sigma_{a}$ is typically chosen to be LeakyReLU, and $\text{softmax}_{j}$ corresponds to the row-wise normalization
$\text{softmax}_{j}(A_{ij})=\frac{\exp(A_{ij})}{\sum_{j^{\prime}}\exp(A_{ij^{\prime}})}.$ (4)
The corresponding feature update is
$X^{(l+1)}=\sigma(A^{(l)}X^{(l)}W^{(l)})\,.$ (5)
3 Oversmoothing
Oversmoothing can be broadly understood as an increase in similarity between node features as inputs are propagated through an increasing number of message-passing layers, leading to a noticeable decline in GNN performance. However, the precise definition of this phenomenon varies across different sources. Some works define oversmoothing more rigorously as the alignment of all feature vectors with each other. This definition is motivated by the behaviour of a linear GCN:
$X^{(l+1)}=A\cdots AX^{(0)}W^{(0)}\dots W^{(l)}.$ (6)
Indeed if $\widetilde{A}$ is the adjacency matrix of a fully connected graph $\mathcal{G}$, $A$ will have a spectral radius equal to $1$ with multiplicity $1$, and $A^{l}$ will thus converge towards the eigenspace spanned by the dominant eigenvector. Precisely, we have
$A^{l}\to uv^{\top}\quad\text{as}\quad l\rightarrow\infty,$ (7)
where $Au=u$, and $A^{\top}v=v$, see e.g. (Tudisco et al., 2015).
As a consequence, if the product of the weight matrices $W^{(0)}\cdots W^{(l)}$ converges in the limit $l\rightarrow\infty$, then the features degenerate to a matrix having rank at most one, where all the features are aligned with the dominant eigenvector $u$. Mathematically, if we assume $u$ to be such that $\|u\|=1$, this alignment can be expressed by stating that the difference between the features and their projection onto $u$, given by $\|X^{(l)}-uu^{T}X^{(l)}\|$, converges to zero.
3.1 Existing Oversmoothing Metrics
Motivated by the discussion about the linear case, oversmoothing is thus quantified and analysed in terms of the convergence of some node similarity metrics towards zero. In particular, in most cases, it is measured exactly by the alignment of the features with the dominant eigenvector of the matrix $A$. The most prominent metric that has been used to quantify oversmoothing is the Dirichlet energy, which measures the norm of the difference between the degree-normalized neighbouring node features (Cai & Wang, 2020; Rusch et al., 2023a)
$E_{\mathrm{Dir}}(X)=\sum_{(i,j)\in\mathcal{E}}\left\|\frac{X_{i}}{u_{i}}-\frac{X_{j}}{u_{j}}\right\|^{2}_{2},$ (8)
where $u_{i}$ is the $i$-th entry of the dominant eigenvector of the message-passing matrix in (1). It thus immediately follows from our discussion on the linear setting that $E_{\mathrm{Dir}}(X^{(l)})$ converges to zero as $l\to\infty$ for a linear GCN with converging weights product $W^{(0)}\cdots W^{(l)}$. This intuition suggests that a similar behaviour may occur for “smooth-enough” nonlinearities.
In particular, in the case of a GCN, the dominant eigenvector $u$ is defined by $u_{i}=\sqrt{1+d_{i}}$ and (Cai & Wang, 2020) have proved that, using LeakyReLU activation functions $E_{\mathrm{Dir}}(X^{(l+1)})\leq s_{l}\bar{\lambda}E_{\mathrm{Dir}}(X^{(l)})$, where $s_{l}=\|W^{(l)}\|_{2}$ is the largest singular value of the weight matrix $W^{(l)}$, and $\bar{\lambda}=(1-\min_{i}\lambda_{i})^{2}$, where $\lambda_{i}\in(0,2]$ varies among the nonzero eigenvalues of the normalized graph Laplacian $\widetilde{\Delta}=I-A=I-\widetilde{D}^{-\frac{1}{2}}\widetilde{A}\widetilde{D}^{-\frac{1}{2}}$.
Similarly, in the case of GATs, the matrices $A_{i}$ are all row stochastic, meaning that $u_{i}=1$ for all $i$. In this case, in (Wu et al., 2023), it has been proved that whenever the product of the entry-wise absolute value of the weights is bounded, that is $\|\Pi_{k=1}^{\infty}|W^{(k)}|\|<\infty$, then the following variant of the Dirichlet energy decays to zero
$E_{\mathrm{Proj}}(X)=\|X-\mathcal{P}X\|^{2}_{F}$ (9)
where $\mathcal{P}=uu^{\top}$ is the projection matrix on the space spanned by the dominant eigenvector $u$ of the matrices $A^{(l)}$. In particular, this metric can be used only if the dominant eigenvector $u$ is the same for all $l$; this is, for example, the case with row stochastic matrices or when $A^{(l)}=A$ for all $l$.
3.2 A unifying perspective based on the eigenvectors of nonlinear activations
In the following discussion, we present a unified and more general perspective of the necessary conditions to have oversmoothing in the sense of the classical metrics, based on the eigenvectors of a nonlinear activation function. In the interest of space, longer proofs for this and the subsequent sections are moved to Appendix A.
Definition 3.1.
We say that a vector $u\in\mathbb{R}^{N}\setminus\{0\}$ is an eigenvector of the (nonlinear) activation function $\sigma:\mathbb{R}^{N}\to\mathbb{R}^{n}$ if for any $t\in\mathbb{R}\setminus\{0\}$, there exists $\mu_{t}$ such that $\sigma(tu)=\mu_{t}u$.
With this definition, we can now provide a unifying characterization of message-passing operators $A^{(l)}$ and activation functions $\sigma$ that guarantee the convergence of the Dirichlet-like energy metric $E_{\mathrm{Proj}}$ to zero for the feature representation sequence defined by $X^{(l+1)}=\sigma(A^{(l)}X^{(l)}W^{(l)})$. Specifically, Theorem 3.3 shows that this holds provided all matrices $A^{(l)}$ share a common dominant eigenvector $u$, which is also an eigenvector of $\sigma$.
This assumption recurs throughout our theoretical analysis, aligning with existing results in the literature. For example, (a) the proof by Cai & Wang (2020) applies to GCNs with LeakyReLU, where the dominant eigenvector of $A$ is nonnegative by the Perron-Frobenius theorem and, therefore, an eigenvector of LeakyReLU; and (b) the proof by Wu et al. (2023) holds for stochastic message-passing matrices $A^{(l)}$, which inherently share a common dominant eigenvector with constant entries, that is an eigenvector for any pointwise nonlinear activation function.
Lemma 3.2.
Let $X^{(l+1)}=\sigma(A^{(l)}X^{(l)}W^{(l)})$ be the $l$-th layer of a GNN. Assume that $A^{(l)}$ has a simple dominant eigenvalue with corresponding eigenvector $u$ that is also an eigenvector of $\sigma$. If $\sigma$ is $1$-Lipschitz, namely $\|\sigma(x)-\sigma(y)\|\leq\|x-y\|$ for any $x,y$, then
$\|(I-\mathcal{P})X^{(l+1)}\|_{F}\leq\|(I-\mathcal{P})A^{(l)}X^{(l)}W^{(l)}\|_{F},$
where $\mathcal{P}=uu^{T}/\|u\|^{2}$ is the projection matrix on the linear space spanned by $u$.
The following theorem follows as a consequence.
Theorem 3.3.
With the same notation of Lemma 3.2, if
$\lim_{L\to\infty}\prod_{l=0}^{L}\|(I-\mathcal{P})A^{(l)}\|_{2}\|W^{(l)}\|_{2}=0,$
then $E_{\mathrm{Proj}}(X^{(l)})\to 0$ as $l\to\infty$.
Proof.
The proof follows observing that in the decomposition $(I-\mathcal{P})A^{(l)}=(I-\mathcal{P})A^{(l)}\mathcal{P}+(I-\mathcal{P})A^{(l)}(I-\mathcal{P})$ the matrix $(I-\mathcal{P})A^{(l)}\mathcal{P}$ is zero because $A^{(l)}u=\lambda_{1}^{l}u$ for any $l$. Thus
$(I-\mathcal{P})A^{(l)}=(I-\mathcal{P})A^{(l)}(I-\mathcal{P}),$
and, from Lemma 3.2 and the inequality $\|AB\|_{F}\leq\|A\|_{2}\|B\|_{F}$, we have $\|(I-\mathcal{P})X^{(L)}\|_{F}\leq\Big{(}\Pi_{l=0}^{L-1}\|(I-\mathcal{P})A^{(l)}\|_{2}\|W^{(l)}\|_{2}\Big{)}\|X^{(0)}\|_{F}$. ∎
Note, in particular, that in the case of GCNs, the matrix $A^{(l)}=A$ is symmetric, and thus $\|I-\mathcal{P}A^{(l)}\|_{2}=\lambda_{2}$. Therefore, when $\sigma=\text{LeakyReLU}$, we obtain the result by Cai & Wang (2020) as convergence to zero is guaranteed if $\|W^{(l)}\|_{2}\leq\lambda_{2}$. Recall that, as discussed above, the choice $\sigma=\text{LeakyReLU}$ satisfies our eigenvector assumption since $u\geq 0$ by the Perron-Frobenius theorem, and thus LeakyReLU$(tu)=\alpha tu$ with $\alpha$ depending only on the sign of $t$. Similarly, in the case of GATs, the matrices $A^{(l)}$ are stochastic for all $l$, implying that $u=\mathbbm{1}$ is the constant vector with $(u)_{i}=1$ for all $i$. If $\sigma=\otimes\psi$ is a nonlinear activation function acting entrywise through $\psi$, then $\sigma(t\mathbbm{1})=\psi(t)\mathbbm{1}$. Therefore, Theorem 3.3 implies that if the weights are sufficiently small, the features align independently of the activation function used. This is consistent with the results in (Wu et al., 2023). However, we note that the bounds on the weights required by Theorem 3.3 and those in (Wu et al., 2023) on the weights $W^{(l)}$ are not identical, and it is unclear which of the two is more significant. Nonetheless, in both cases, having bounded weights along with any $1$-Lipschitz pointwise activation function is a sufficient condition for observing oversmoothing via $E_{\mathrm{Proj}}$ in GATs. In addition to offering a different and unifying theoretical perspective on the results in (Cai & Wang, 2020; Wu et al., 2023), we highlight the simplicity of our eigenvector-based proof, which we hope provides added clarity on this phenomenon.
4 Energy-like metrics: what can go wrong
Energy-like metrics such as $E_{\mathrm{Dir}}$ and $E_{\mathrm{Proj}}$ are among the most commonly used oversmoothing metrics. However, they suffer from inherent limitations that hinder their practical usability and informational content.
One important limitation of these metrics is that they indicate oversmoothing only in the limit of infinitely many layers, when their values converge exactly to zero. Since they measure a form of absolute distance, a small but nonzero value does not provide any meaningful information. On the other hand, convergence to zero corresponds to either perfect alignment with the dominant eigenspace or the collapse of the feature representation matrix to the all-zero matrix. While the former is a symptom of oversmoothing, the latter does not necessarily imply oversmoothing. Moreover, this convergence property requires the weights to be bounded. However, in practical cases, performance degradation is observed even in relatively shallow networks, far from being infinitely deep, and with arbitrarily large (or small) weight magnitudes. This aligns with our intuition and what occurs in the linear case. Indeed, for a linear GCN, even when the features $X^{(l)}$ grow to infinity as $l\to\infty$, one observes that $X^{(l)}$ becomes dominated by the dominant eigenspace of $A$, even for finite and possibly small values of $l$, depending on the spectral gap of the graph.
More precisely, the following theorem holds:
Theorem 4.1.
Let $X^{(l+1)}=AX^{(l)}W^{(l)}$ be a linear GCN. Let $\lambda_{1},\lambda_{2}$ be the largest and second-largest eigenvalues (in modulus) of $A$, respectively. Assume the weights $\{W^{(l)}\}_{l=1}^{\infty}$ are randomly sampled from i.i.d. random variables with distribution $\nu$ such that
[…]
$\lim_{l\to\infty}\frac{\|(I-\mathcal{P})X^{(l)}\|_{F}}{\|\mathcal{P}X^{(l)}\|_{F}}=0$
with a linear rate of convergence $O(|\lambda_{2}/\lambda_{1}|^{l})$.
In particular, the theorem above implies that for a large spectral gap $|\lambda_{2}/\lambda_{1}|\ll 1$, $X^{(l)}$ is predominantly of rank one, namely
$X^{(l)}=\lambda_{1}^{l}\Big{(}uv^{T}+R(l)\Big{)}$
for some $v$, with $R(l)\sim O(|\lambda_{2}/\lambda_{1}|)^{l}$ and thus converging to 00 as $l\to\infty$. This results in weakly expressive feature representations, independently of the magnitude of the feature weights. This phenomenon can be effectively captured by measuring the rank of $X^{(l)}$, whereas Dirichlet-like energy measures may fail to detect it.
Another important limitation of these metrics is their dependence on a specific, known dominant eigenspace, which must either be explicitly known or computed in advance. Consequently, their applicability is strongly tied to the specific architecture of the network. In particular, the dominant eigenvector $u$ of $A^{(l)}$ must be known and remain the same for all $l$. This requirement excludes their use in most cases where $A^{(l)}$ varies with $l$, such as when $A^{(l)}$ is the standard weighted adjacency matrix of the graph.
5 The Rank as a Measure of Oversmoothing
Inspired by the behaviour observed in the linear case, we argue that measuring the rank of feature representations provides a more effective way to quantify oversmoothing, in alignment with recent work on oversmoothing (Guo et al., 2023). However, since the rank of a matrix is defined as the number of nonzero singular values, it is a discrete function and thus not suitable as a measure. A viable alternative is to use a continuous relaxation that closely approximates the rank itself.
Examples of possible continuous approximations of the rank include the numerical rank, the stable rank, and the effective rank (Roy & Vetterli, 2007; Rudelson & Vershynin, 2006; Arora et al., 2019). The stable rank is defined as $\text{StableRank}(X)=\|X\|_{*}^{2}/\|X\|_{F}^{2}$, where $\|X\|_{*}=\sum_{i}\sigma_{i}$ is the nuclear norm. The numerical rank is given by $\text{NumRank}(X)=\|X\|_{F}^{2}/\|X\|_{2}^{2}$. Finally, given the singular values $\sigma_{1}>\sigma_{2}>\dots>\sigma_{\min\{N,d\}}$ of $X$, the effective rank is defined as
$\text{Erank}(X)=\exp\left(-\textstyle{\sum_{k}p_{k}\log p_{k}}\right),$ (10)
where $p_{k}=\sigma_{k}/\sum_{i}\sigma_{i}$ is the $k$-th normalized singular value. These rank relaxation measures exhibit similar empirical behaviour as shown by our numerical evaluation in Section 6.
In practice, measuring oversmoothing in terms of a continuous approximation of the rank is a reasonable approach that helps address the limitations of Dirichlet-like measures. Specifically, it offers the following advantages: (a) it is scale-invariant, meaning it remains informative even when the feature matrix converges to zero or explodes to infinity; (b) it does not rely on a fixed, predetermined eigenspace but instead captures convergence of the feature matrix toward an arbitrary lower-dimensional subspace; (c) it allows for the detection of oversmoothing in shallow networks without requiring exact convergence to rank one—since a small value of, say, the effective rank, directly implies that the feature representations are low-rank, which in turn suggests a potentially suboptimal network architecture.
In Table 1, we present a toy example illustrating that classical oversmoothing metrics fail to correctly capture oversmoothing unless the features are perfectly aligned. This observation implies that these metrics can quantify oversmoothing only when the rank of the feature matrix converges exactly to one. In contrast, continuous rank functions provide a more reliable measure of approximate feature alignment. Later, in Figure 1, we demonstrate that the same phenomenon occurs in GNNs trained on real datasets, where exact feature alignment is rare. In such cases, classical metrics remain roughly constant, whereas the rank decreases, coinciding with a sharp drop in GNN accuracy.
Table 1: A set of toy scenarios depicting the behaviour of oversmoothing metrics. Each plot contains 50 nodes, each with two features plotted as the x and y axis. # 1: all features are of the same value; # 2: all features are assumed to align with the same vector exactly; # 3: shifts one (red) point in # 2 in y-direction; # 4: samples all points from a uniform distribution. MAD (see Section 6) and $E_{\mathrm{Dir}}$ give false negative signals in # 3 while the features are considered oversmoothing by definition. $E_{\mathrm{Proj}}$ can hardly differentiate between # 3 and # 4, and is thus not robust in quantifying oversmoothing. To compute $E_{\mathrm{Proj}}$ and $E_{\mathrm{Dir}}$ the first feature was considered in place of $u$ in (8) and (9).
# 1 # 2 # 3 # 4
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
$E_{\mathrm{Dir}}$ 0 0 13.25 77.78
$E_{\mathrm{Proj}}$ 0 0 0.83 0.97
MAD 0 0.81 0.81 0.57
NumRank 1 1 1.01 1.78
Erank 1 1 1.36 1.99
5.1 Theoretical Analysis of Rank Decay
In this section, we provide an analytical study proving the decrease of the numerical rank for a broad class of graph neural network architectures under the assumption that the weight matrices are entrywise nonnegative. While this is a somewhat restrictive setting, our result is the first theoretical proof that oversmoothing occurs independently of the weight (and thus feature) magnitude.
We begin with several useful observations. Let $u$ be the dominant eigenvector of $A$ corresponding to $\lambda_{1}$ and satisfying $\|u\|=1$. Consider the projection matrix $\mathcal{P}=uu^{\top}$. Given a matrix $X$, we can decompose it as $X=\mathcal{P}X+(I-\mathcal{P})X$. Since $u$ is a unit vector, it follows that $\|\mathcal{P}\|_{2}=1$, and therefore,
$\|X\|_{2}=\|\mathcal{P}\|_{2}\|X\|_{2}\geq\|\mathcal{P}X\|_{2}.$ (11)
Moreover, since $\mathcal{P}X$ and $(I-\mathcal{P})X$ are orthogonal with respect to the Frobenius inner product, we have $\|\mathcal{P}X+(I-\mathcal{P})X\|_{F}^{2}=\|\mathcal{P}X\|_{F}^{2}+\|(I-\mathcal{P})X\|_{F}^{2}$. Thus, we obtain the following bound for the numerical rank of $X$:
$\text{NumRank}(X)$ $=\frac{\|\mathcal{P}X+(I-\mathcal{P})X\|_{F}^{2}}{\|X\|_{2}^{2}}$ (12)
$=\frac{\|\mathcal{P}X\|_{F}^{2}+\|(I-\mathcal{P})X\|_{F}^{2}}{\|X\|_{2}^{2}}$
$\leq 1+\frac{\|(I-\mathcal{P})X\|_{F}^{2}}{\|X\|_{2}^{2}},$
where we used (11) and the fact that $\|\mathcal{P}X\|_{F}=\|\mathcal{P}X\|_{2}$ since $\mathcal{P}X$ is a rank-one matrix.
The above inequality, together with Theorem 4.1, allows us to establish the convergence of the numerical rank for linear networks.
The Linear Case
Consider a linear GCN of the form $X^{(l+1)}=AX^{(l)}W^{(l)}$, where $A$ has a simple dominant eigenvalue $\lambda_{1}$ satisfying $|\lambda_{1}|\geq|\lambda_{2}|$.
We have already noted that $\|X\|_{2}\geq\|\mathcal{P}X\|_{2}$, meaning that the numerical rank converges to one if $\|(I-\mathcal{P})X\|_{F}/\|X\|_{2}$ decays to zero. This occurs whenever the features grow faster in the direction of the dominant eigenvector than in any other direction. As established in Theorem 4.1, this is almost surely the case in linear GNNs. As a direct consequence, we obtain the following result:
Theorem 5.1.
Let $X^{(l+1)}=AX^{(l)}W^{(l)}$ be a linear GCN. Under the same assumptions as in Theorem 4.1, the following identity holds almost surely:
$\lim_{l\rightarrow\infty}\mathrm{NumRank}(X^{(l)})=1.$
Extending the result above to general GNNs with nonlinear activation functions is highly nontrivial. In the next section, we present our main theoretical result, which generalizes this analysis to a broader class of GNNs under the assumption of nonnegative weights.
5.2 The Nonnegative Setting
To study the case of networks with nonnegative weights, we make use of tools from the nonlinear Perron-Frobenius theory; we refer to (Lemmens & Nussbaum, 2012; Gautier et al., 2023) and the reference therein for further details.
We assume all the intermediate features to be in the positive open cone $\mathcal{K}:=\mathbb{R}_{+}^{N}=\{x\in\mathbb{R}^{N}\>|\>x_{i}>0\;\forall i=1,\dots,N\}$. On $\mathcal{K}$, it is possible to introduce the partial ordering
$x\leq_{\mathcal{K}}y\>(x\ll_{\mathcal{K}}y)\quad\text{iff}\quad y-x\in\bar{\mathcal{K}}\>(y-x\in\mathcal{K})$ (13)
where $\bar{\mathcal{K}}$ denotes the nonnegative closed cone. Given two points $x,y\in\mathcal{K}$ we write $M(x/y)=\max_{i}x_{i}/y_{i}$ and $m(x/y)=\min_{i}x_{i}/y_{i}$. Then, we can define the following Hilbert distance between any two points $x,y\in\mathcal{K}$,
$d_{H}(x,y)=\log(M(x/y)/m(x/y)).$
Note that $d_{H}$ is not a distance on $\mathcal{K}$, indeed $d_{H}(\alpha x,\beta y)=d_{H}(x,y)$ for any $x,y\in\mathcal{K}$ and $\alpha,\beta>0$. However it is a distance up to scaling; that is, it becomes a concrete distance whenever we restrict ourselves to a slice of the cone.
Now, it is well-known that any matrix $A$ such that $Ax\in\mathcal{K}$ for any $x\in\mathcal{K}$ is non-expansive with respect to the Hilbert distance, i.e.
$d_{H}(Ax,Ay)\leq\beta d_{H}(x,y)\qquad\forall x,y\in\mathcal{K}$ (14)
with $0\leq\beta\leq 1$. The last inequality is a consequence of the fact that $A$ preserves the ordering induced by the cone, i.e. if $x\geq_{\mathcal{K}}y$, then $Ax\geq_{\mathcal{K}}Ay$. In particular, we recall that (14) always hold for $\beta<1$ when $A$ is entry-wise strictly positive, as positive matrices map the whole $\bar{\mathcal{K}}$ in its interior, i.e. $x\geq_{\mathcal{K}}y$ implies $Ax>Ay$. In the next definition, we include all the nonnegative matrices that have the same behaviour as the positive matrices but only on a neighbourhood of a positive vector in $\mathcal{K}$.
Recall that any irreducible nonnegative matrix $A$ has a positive eigenvector $u$ corresponding to its dominant eigenvalue.
Definition 5.2.
A family of nonnegative irreducible matrices (possibly one single matrix) $\{A^{(l)}\}$ with the same dominant eigenvector $u>0$ is uniformly contractive with respect to $u$ if for any $C>0$ there exists some $\beta_{C}<1$ such that $d_{H}(A^{(l)}x,u)\leq\beta_{C}d_{H}(x,u)$ for all $l$ and $x\in\mathcal{K}$ such that $d_{H}(x,u)\leq C$.
The above definition is significantly weaker than asking for $A$ to be strictly positive. We show this with an illustrative example in Section A.3.
Consider now first the case of a linear function that can be represented in the form
$F(X)=AXW,$ (15)
where $X$ is a positive $N\times d$ matrix, $A$ and $W$ are nonnegative $N\times N$ and $d\times d$ matrices, respectively. Then, under mild assumptions, we can prove that the columns of $F(X)$ are closer to the dominant eigenvector $u$ of $A$ as compared to the columns of $X$.
Lemma 5.3.
Let $F(X)=AXW$ with $A$ nonnegative and irreducible with dominant eigenvector $u\in\mathcal{K}$. Assume also $X$ to be striclty positve and $W$ nonnegative with $\min_{j}\max_{i}W_{ij}>0$. If $A$ is contractive with respect to $u$ and $C=\max_{i}d_{H}(X_{i},u)$ then
$\max_{i}d_{H}(F(X)_{i},u)\leq\beta_{C}\max_{i}d_{H}(X_{i},u),$
where $Y_{i}$ denotes the $i$-th column of $Y$, and $\beta_{C}\leq 1$.
Nonlinear Perron-Frobenius theory extends some of the results about nonnegative matrices to particular classes of nonlinear functions. A function $\sigma\in C(\mathcal{K},\mathcal{K})$ is order preserving if given any $x,y\in\mathcal{K}$ with $x\geq_{\mathcal{K}}y$ then $\sigma(x)\geq_{\mathcal{K}}\sigma(y)$. Moreover $\sigma$ is subhomogenenous if $\sigma(\lambda x)\leq_{\mathcal{K}}\lambda\sigma(x)$ for all $x\in\mathcal{K}$ and any $\lambda>1$. In particular, it is strictly subhomogeneous if $\sigma(\lambda x)\ll_{\mathcal{K}}\lambda\sigma(x)$ for all $x\in\mathcal{K}$ and $\lambda>1$ and homogeneous if $\sigma(\lambda x)=\lambda\sigma(x)$ for all $x\in\mathcal{K}$ and $\lambda\geq 0$
Subohomogeneity is a useful property and of practical utility. In fact, as discussed in, e.g. (Sittoni & Tudisco, 2024; Piotrowski et al., 2024), it is not difficult to verify that a broad range of activation functions commonly used in deep learning is subhomogeneous on $\mathcal{K}$. For this family of activation functions, Lemma 5.3 combined with arguments from nonlinear Perron–Frobenius theory yield the following main result
Theorem 5.4.
Consider a positive GNN
$X^{(l+1)}=\sigma(A^{(l)}X^{(l)}W^{(l)})$ (16)
with $X^{(0)}_{i}\in\mathcal{K}$ for any $i=1,\dots,d$ and $A^{(l)}$ and $W^{(l)}$ nonnegative for any $l$. Assume also that $u\in\mathcal{K}$ is the dominant eigenvector of all of the matrices $A^{(i)}$ and that it is also an eigenvector of $\sigma$. Then, if the matrices $\{A^{(i)}\}$ are uniformly contractive with respect to $u$ and $\min_{j}\max_{i}{W^{(l)}}_{ij}>0$ for any $l$, it holds
$\lim_{l\rightarrow\infty}\mathrm{NumRank}(X^{(l)})=1.$
We conclude with a formal investigation of the eigenpairs of activation functions that are entrywise subhomogeneous. Let $\sigma=\otimes^{N}\psi$ with $\psi\in C(\mathbb{R},\mathbb{R})$ that is subhomogeneous on $\mathbb{R}_{+}$. Then one can easily show that $\sigma$ is itself subhomogeneous on +$\mathbb{R}^{N}_{+}$. We have,
Proposition 5.5.
Let $\sigma=\otimes^{N}\psi$ with $\psi\in C(\mathbb{R}_{+},\mathbb{R}_{+})$ be order preserving. Then: 1) If $\sigma$ is homogeneous, any positive vector is an eigenvector of $\sigma$. 2) If $\sigma$ is strictly subhomogeneous, the only eigenvector of $\sigma$ in $\mathcal{K}$ is the constant vector.
As a consequence of the above result, we find that the numerical rank of the features collapses both for GCNs with LeakyReLU activation function (that is homogeneous) as well as for GATs with any kind of subhomogeneous activation function. In fact, the constant vector is always an eigenvector of an entriwise nonlinear map.
Table 2: The table lists the correlation coefficients of logarithm of metric values and the classification accuracies for GNNs of depth 2 - 24 layers and separately trained on Cora dataset. For Erank and NumRank, we subtract 1 so that both metrics converge to zero. The classification accuracy ratio are computed on 2-layered and 24-layered GNNs.
Architecture $E_{\mathrm{Dir}}$ $E_{\mathrm{Proj}}$ MAD Erank NumRank Accuracy ratio
Standard Normalized Standard Normalized
LReLU + GCN -0.7871 0.6644 -0.8106 -0.8309 -0.2460 0.9724 0.5885 0.2689
Tanh + GCN 0.5243 0.9403 0.8610 0.9768 0.9734 0.9923 0.9784 0.1937
LReLU + GCN + Bias -0.9300 0.7505 -0.9414 -0.3642 -0.2828 0.9847 0.7833 0.2108
Tanh + GCN + Bias 0.4086 0.8950 0.7479 0.9026 0.8996 0.9926 0.9814 0.2133
LReLU + GCN + LayerNorm -0.9132 0.8445 -0.9424 0.5591 0.9753 0.9736 0.9736 0.4882
Tanh + GCN + LayerNorm -0.7119 0.2560 -0.2069 0.9716 0.8695 0.9576 0.9684 0.1885
LReLU + GCN + PairNorm -0.8165 0.3731 -0.8106 0.3885 0.4850 0.5597 0.6244 0.8252
Tanh + GCN + PairNorm -0.5963 0.0346 -0.5401 -0.3358 -0.0631 0.9838 0.9298 0.2902
LReLU + GAT -0.9189 0.6703 -0.9469 -0.6054 0.8251 0.9722 0.7612 0.2487
Tanh + GAT 0.8300 0.8501 0.8676 0.9066 0.8603 0.9600 0.9515 0.1767
Average correlation -0.3911 0.6279 -0.2722 0.2568 0.5296 0.9349 0.8541
6 Experiments
Empirical studies on the evolution of oversmoothing measures often use untrained, hundred-layer-deep GNNs (Rusch et al., 2022; Wang et al., 2022; Rusch et al., 2023b; Wu et al., 2023). We emphasize that this is an overly simplified setting. In more realistic settings, as the ones considered in this section, a trained GNN may suffer from significant performance degradation after only few-layers, at which stage the convergence patterns of most of the metrics are difficult to observe. The experiments that we present in this section validate the robustness of the effective rank and numerical rank in quantifying oversmoothing in GNNs against the other metrics.
In particular, we compare how different overmoothing metrics behave compared to the classification accuracy, varying the GNN architectures for node classification on real-world graph data. In our experiments, we consider the following metrics:
  • The Dirichlet Energy $E_{\mathrm{Dir}}$ (Cai & Wang, 2020; Rusch et al., 2023a) and its variant $E_{\mathrm{Proj}}$ (Wu et al., 2023). Both are discussed in Section 3.1, see in particular (8) and (9).
  • Normalized versions of the Dirichlet energy and its variant, namely $E_{\mathrm{Dir}}(X)/\|X\|_{F}^{2}$ and $E_{\mathrm{Proj}}(X)/\|X\|_{F}$. Indeed, from our previous discussion, a robust oversmoothing measure should be scale invariant with respect to the features. Metrics with global normalization like the ones we consider here have also been proposed in (Di Giovanni et al., 2023; Roth & Liebig, 2023; Maskey et al., 2023).
  • The Mean Average Distance (MAD) (Chen et al., 2020)
    $\text{MAD}(X)=\frac{1}{|\mathcal{E}|}\sum_{(i,j)\in\mathcal{E}}\left(1-\frac{X_{i}^{\top}X_{j}}{|X_{i}||X_{j}|}\right).$ (17)
    It measures the cosine similarity between the neighbouring nodes. Unlike previous baselines, this oversmoothing metric does not take into account the dominant eigenvector of the matrices $A^{(l)}$.
  • Relaxed rank metrics: We consider the Numerical Rank and Effective Rank. Both are discussed in Section 5. We point out that from our theoretical investigation, in particular from (12), the numerical rank decays to $1$ faster than the decay of the normalized $E_{\mathrm{Proj}}$ energy to zero. This further supports the use of the Numerical Rank as an improved measure of oversmoothing with respect to $E_{\mathrm{Proj}}$.
Figure 1: Four examples of GCNs corresponding to the first fours rows in table 2. For Erank and Numrank, we measure Erank$(X)-r^{*}_{\mathrm{ER}}$ and $\mathrm{NumRank}(X)-r^{*}_{\mathrm{NR}}$ for some $r^{*}>1$. In these particular cases, $r^{*}_{\mathrm{ER}}<1.85$, $r^{*}_{\mathrm{NR}}<1.3$. Note that the effective rank and numerical rank of the input features $X^{(0)}$ is about 1084 and 13.6, respectively.
In Table 2 and Figure 1, we train GNNs of a fixed hidden dimension equal to 32 on the Cora dataset. More results on Cora, as well as other datasets, are reported in Appendix B. We follow the standard setups of GCN and GAT as stated in equation (2) and (5), and use homogeneous LeakyReLU (LReLU) and subhomogenous Tanh as activation functions. For the GCN model, we also consider adding additional components, i.e. bias, LayerNorm (Ba et al., 2016) and PairNorm (Zhao & Akoglu, 2019). For each configuration, GNNs of eight different depths ranging from 2 to 24 are trained. The oversmoothing metric and accuracy results are averaged over 10 separately trained GNNs. All GNNs are trained with NAdam Optimizer and a constant learning rate of 0.01. The oversmoothing metrics are computed at the last hidden layer before the output layer. In Figure 1 and in Appendix B, we plot the behaviour of the different oversmoothing measures as well as the norm of the features and the accuracy of the trained GNNs when the depth of the network is increased. These figures clearly show that the network suffers a significant drop in accuracy, which is not matched by any visible change in standard oversmoothing metrics. By contrast, the rank of the feature representations decreases drastically, following quite closely the behaviour of the network’s accuracy. These findings are further sustained by the results shown in Table 2, where we compute the Pearson correlation coefficient between the logarithm of every measure and the classification accuracy of every GNN model. The use of a logarithmic transformation is based on the understanding that oversmoothing grows exponentially with the length of the network.
7 Conclusion
In this paper, we have discussed the problem of quantifying oversmoothing in GNNs. After discussing the limitations of the leading oversmoothing measures, we have proposed the use of the rank of the features as a better measure of oversmoothing. The experiments that we provided validate the robustness of the effective rank against the classical measures. In addition, we have proved theoretically the decay of the rank of the features for linear and nonnegative GNNs.
Appendix A Proofs of the main results
A.1 proof of Lemma 3.2
Let $\pi:=\mathrm{span}\{uv^{T}\,|\;v\in\mathbb{R}^{d}\}$ be the $1$-dimensional matrix subspace of the rank-$1$ matrices having columns aligned to $u$. Then it is easy to note that given some matrix $X$, $(I-\mathcal{P})X$ provides the projection of the matrix $X$ on the subspace $\pi$, i.e.
$(I-\mathcal{P})X=\mathrm{proj}_{\pi}(X).$ (18)
Indeed $\langle(I-\mathcal{P})X,uv^{T}\rangle_{F}=\mathrm{Tr}(vu^{T}(I-uu^{T})X)=0$. In particular, since the projection realizes the minimal distance, we have that
$\|X-\mathcal{P}X\|_{F}\leq\|X-uv^{T}\|_{F}\qquad\forall v\in\mathbb{R}^{d}.$ (19)
Now observe that $\sigma(uu^{T}A^{(l-1)}X^{(l-1)}W^{(l-1)})=u\bar{v}^{T}$ for some $\bar{v}$. Indeed, writing $v^{T}=u^{T}A^{(l-1)}X^{(l-1)}W^{(l-1)}$, we have that the $i$-th column of $\sigma(uu^{T}A^{(l-1)}X^{(l-1)}W^{(l-1)})$ is equal to $\sigma(v_{i}u)=\bar{v}_{i}u$ for some $\bar{v}_{i}$, because $u$ is an eigenvector of $\sigma$. As a consequence we have
$\|(I-\mathcal{P})X^{(l)}\|_{F}$ $\leq\|X^{(l)}-\sigma(uu^{T}A^{(l-1)}X^{(l-1)}W^{(l-1)})\|_{F}$ (20)
$=\|\sigma(A^{(l-1)}X^{(l-1)}W^{(l-1)})-\sigma(uu^{T}A^{(l-1)}X^{(l-1)}W^{(l-1)})\|_{F}$
$\leq\|(I-\mathcal{P})A^{(l-1)}X^{(l-1)}W^{(l-1)}\|_{F}$
where we have used the $1$-Lipschitz property of $\sigma$.
A.2 Proof for Theorem 4.1
Start by studying the norm of $(I-\mathcal{P})X^{(l)}$. Then looking at the shape of the powers of the Jordan blocks matrix it is not difficult to note that $\widetilde{T}^{l}=O({l\choose N}\lambda_{2}^{l-N})$ for $l$ larger than $N$. In particular if we look at the explicit expression of $(I-\mathcal{P})X^{(l)}$
$(I-P)X^{(l)}=\begin{pmatrix}0&(I-\mathcal{P})\tilde{M}\end{pmatrix}\begin{pmatrix}0&0\\ 0&\tilde{T}^{l}\end{pmatrix}M^{-1}X^{(0)}W^{(0)}\dots W^{(l-1)},$ (21)
we derive the upper bound
$\|(I-P)X^{(l)}\|_{F}\leq C{l\choose N}|\lambda_{2}|^{l-N}\|X^{(0)}W^{(0)}\dots W^{(l-1)}\|_{F},$ (22)
for some positive constant $C$ that is independent on $l$.
Similarly we can observe that
$\|\mathcal{P}X^{(l)}\|_{F}$ $\geq\|u^{T}A^{l}X^{(0)}W^{(0)}\dots W^{(l-1)}\|_{F}=$ (23)
$\|\big{(}\lambda_{1}^{l}v_{1}^{T}+u^{T}\widetilde{M}O\Big{(}{l\choose N}\lambda_{2}^{l-N}\Big{)}\widetilde{M^{\prime}}\big{)}X^{(0)}W^{(0)}\dots W^{(l-1)}\|_{F}\geq$
$|\lambda_{1}|^{l}\Big{(}\|v_{1}^{T}X^{(0)}W^{(0)}\dots W^{(l-1)}\|_{F}-\Big{\|}u^{T}\widetilde{M}O\Big{(}{l\choose N}\Big{(}\frac{\lambda_{2}}{\lambda_{1}}\Big{)}^{l}\widetilde{M^{\prime}}X^{(0)}W^{(0)}\dots W^{(l-1)}\Big{\|}_{F}\Big{)}\geq$
$|\lambda_{1}|^{l}\|v_{1}^{T}X^{(0)}W^{(0)}\dots W^{(l-1)}\|_{F}\Big{(}1-O\Big{(}{l\choose N}\Big{|}\frac{\lambda_{2}}{\lambda_{1}}\Big{|}^{l}\frac{\|X^{(0)}W^{(0)}\dots W^{(l-1)}\|_{F}}{\|v_{1}^{T}X^{(0)}W^{(0)}\dots W^{(l-1)}\|_{F}}\Big{)}.$
Now observe that under the randomness hypothesis from (Furstenberg & Kifer, 1983) and more generally from the Oseledets ergodic multiplicative theorem, we have that for almost any $w\in\mathbb{R}^{d}$ the limit exists and is equal to the maximal Lyapunov exponent of the system, i.e. a constant $c(\nu)\geq 0$ depending only on the distribution $\nu$, $\lim_{l\rightarrow\infty}\frac{1}{l}\log\|w^{T}W^{(0)}\dots W^{(l-1)}\|=c(\nu)$. In particular for any $w$ and $\epsilon>0$ there exists $l_{w,\epsilon}$ sufficiently large such that for any $l>l_{w,\epsilon}$
$c(\nu)-\epsilon\leq\frac{1}{l}\log\|w^{T}W^{(0)}\dots W^{(l-1)}\|<c(\nu)+\epsilon.$ (24)
i.e.
$e^{l(c(\nu)-\epsilon)}\leq\|w^{T}W^{(0)}\dots W^{(l-1)}\|<e^{l(c(\nu)+\epsilon)}\quad\forall l\geq l_{w,\epsilon}$ (25)
Now take as vector $w$ first the rows of $X^{(0)}$ and then the vector $v_{1}^{T}X^{(0)}$, than almost surely for any $\epsilon$ there exists $l_{\epsilon}$ such that for any $l>l_{\epsilon}$
$e^{l(c(\nu)-\epsilon)}\leq\|w^{T}W^{(0)}\dots W^{(l-1)}\|<e^{l(c(\nu)+\epsilon)},$ (26)
holding for any $l\geq l_{\epsilon}$ and any $w\in\{v_{1}\}\cup\{X_{0}^{T}e_{i}\}_{i=1}^{N}$.
Next recall that $\|X^{(0)}W^{(0)}\dots W^{(l-1)}\|_{F}=\sqrt{\sum_{i}\|e_{i}^{T}X^{(0)}W^{(0)}\|^{2}}$, meaning that almost surely, for $l\geq l_{\epsilon}$:
$Ne^{l(c(\nu)-\epsilon)}\leq\|X^{(0)}W^{(0)}\dots W^{(l-1)}\|_{F}\leq Ne^{l(c(\nu)+\epsilon)}.$ (27)
In particular for any $\epsilon$, there exists $l$ sufficiently large such that
($\Big{(}{l\choose N}\Big{|}\frac{\lambda_{2}}{\lambda_{1}}\Big{|}^{l}\frac{\|X^{(0)}W^{(0)}\dots W^{(l-1)}\|_{F}}{\|v_{1}^{T}X^{(0)}W^{(0)}\dots W^{(l-1)}\|_{F}}\Big{)}\leq\Big{(}{l\choose N}\Big{|}\frac{\lambda_{2}}{\lambda_{1}}\Big{|}^{l}e^{2l\epsilon}\Big{)}$ (28)
and thus, since $|\lambda_{2}|<|\lambda_{1}|$ and we can choose $\epsilon$ arbitrarily small, almost surely it has limit equal to zero. In particular we can write
$\lim_{l}\frac{\|(I-\mathcal{P}X^{(l)})\|_{F}}{\|\mathcal{P}X^{(l)}\|_{F}}\sim\lim_{l}\frac{{l\choose N}|\lambda_{2}|^{l-N}\|X^{(0)}W^{(0)}\dots W^{(l)}\|_{F}}{|\lambda_{1}|^{l}\|v_{1}^{T}X^{(0)}W^{(0)}\dots W^{(l)}\|_{F}}=0$ (29)
where we have used the same argument as before to state that the limit is zero.
A.3 An illustrative example
Consider the stochastic nonnegative primitive matrix
$A:=\begin{pmatrix}0&1\\[4.30554pt] 1/2&1/2\end{pmatrix}.$ (30)
Since $A$ is row stochastic, the dominant eigenvector is given by the constant vector $u=(1,1)$. In particular given a vector $x=(x_{1},x_{2})\in\mathcal{K}$ we have that
$d_{H}(x,u)=\log\Big{(}\frac{\max_{i=1,2}\{x_{i}\}}{\min_{i=1,2}\{x_{i}\}}\Big{)}=\begin{cases}\log\big{(}x_{1}/x_{2}\big{)}\quad\text{if}\;x_{1}\geq x_{2}\\ \log\big{(}x_{2}/x_{1}\big{)}\quad\text{if}\;x_{2}\geq x_{1}\end{cases}.$ (31)
On the other hand $Ax=\big{(}x_{2}\,,\,(x_{1}+x_{2})/2\big{)}$, thus
$d_{H}(Ax,u)=\begin{cases}\log\big{(}(x_{1}+x_{2})/2x_{2}\big{)}\quad\text{if}\;x_{1}\geq x_{2}\\ \log\big{(}2x_{2}/(x_{1}+x_{2})\big{)}\quad\text{if}\;x_{2}\geq x_{1}\end{cases}.$ (32)
Now observe that the set of vectors $x$ such that $d_{H}(x,u)\leq C$ is given by the points $\{x\,|\;x_{1}\leq x_{2}\leq C^{*}x_{1}\}\cup\{x\,|\;x_{2}\leq x_{1}\leq C^{*}x_{2}\}$ with $C^{*}=e^{C}>1$. We want to prove the existence of some $\beta_{C}=1-\epsilon_{C}\in(0,1)$ such that $d_{H}(Ax,u)\leq\beta_{C}d_{H}(x,u)$ for any $x$ with $d_{H}(x,u)\leq C$.
To this end consider first the case of $x\in\Omega_{1}=\{x\,|\;x_{1}\leq x_{2}\leq C^{*}x_{1}\}$. In this case
$d_{H}(Ax,u)=\log\Big{(}\frac{2x_{2}}{x_{1}+x_{2}}\Big{)}\leq(1-\epsilon_{C})\log\Big{(}\frac{x_{2}}{x_{1}}\Big{)}=(1-\epsilon_{C})d_{H}(x,u)\qquad\forall x\in\Omega_{1}$ (33)
$\Longleftrightarrow\quad\epsilon_{C}\log\Big{(}\frac{x_{2}}{x_{1}}\Big{)}\leq\log\Big{(}\frac{x_{2}(x_{1}+x_{2})}{2x_{1}x_{2}}\Big{)}\qquad\forall x\in\Omega_{1}$
$\Longleftrightarrow\quad\epsilon_{C}\log(t)\leq\log\Big{(}\frac{1+t}{2}\Big{)}\qquad\forall t\in[1,C^{*}]$
$\Longleftrightarrow t^{\epsilon_{C}}\leq\frac{1+t}{2}\qquad\forall t\in[1,C^{*}]$
where we have written $t=x_{2}/x_{1}$. Second we consider the case of $x\in\Omega_{2}=\{x\,|\;x_{2}\leq x_{1}\leq C^{*}x_{2}\}$. In this case
$d_{H}(Ax,u)=\log\Big{(}\frac{x_{1}+x_{2}}{2x_{2}}\Big{)}\leq(1-\epsilon_{C})\log\Big{(}\frac{x_{1}}{x_{2}}\Big{)}=(1-\epsilon_{C})d_{H}(x,u)\qquad\forall x\in\Omega_{2}$ (34)
$\Longleftrightarrow\quad\epsilon_{C}\log\Big{(}\frac{x_{1}}{x_{2}}\Big{)}\leq\log\Big{(}\frac{2x_{1}x_{2}}{x_{2}(x_{1}+x_{2})}\Big{)}\qquad\forall x\in\Omega_{2}$
$\Longleftrightarrow\quad\epsilon_{C}\log(t)\leq\log\Big{(}\frac{2t}{t+1}\Big{)}\qquad\forall t\in[1,C^{*}]$
$\Longleftrightarrow t^{\epsilon_{C}}\leq\frac{2t}{t+1}\qquad\forall t\in[1,C^{*}]$
where we have written $t=x_{1}/x_{2}$. Both the functions $t\rightarrow(1+t)/2$ and $t\rightarrow 2t/(t+1)$ are monotonically increasing for $t\in[1,\infty)$ with derivative $1/2$ in $t=1$, thus for any $C<\infty$ there exists some $10<\epsilon_{C}<$ 1 such that both (33) and (33) are satisfied in the interval $t\in[1,C^{*}]$. This shows that $A$ is contractive with respect to $u$ even if $A$ is not strictly positive.
A.4 proof of Lemma 5.3
By the contraction properties of the matrix $A$ we know that if $\max_{i}d_{H}\big{(}(AX)_{i},u\big{)}\leq C$, then
$d_{H}\Big{(}(AX)_{i},u\Big{)}=d_{H}\Big{(}(AX)_{i},\lambda_{1}(A)u\Big{)}=d_{H}\Big{(}(AX)_{i},Au\Big{)}\leq\beta d_{H}\big{(}X_{i},u\big{)}\qquad\forall i.$ (35)
for some $\beta<1$, where we have used $\lambda_{1}(A)>0$ and the scaling invariant property of the Hilbert distance.
Then note that, for any $i$, we can write $F(X)_{i}$ as follows
$\big{(}F(X)\big{)}_{j}=\sum_{j}W_{ij}(AX)_{j}\,.$ (36)
Thus we CLAIM that given $x_{1},x_{2},y\in\mathcal{K}$ then
$d_{H}(x_{1}+x_{2},y)\leq\max\{d_{H}(x_{1},y),d_{H}(x_{2},y)\}.$ (37)
Observe that if the claim holds we have conlcuded the proof, indeed by induction it can trivially be extended from $2$ to $d$ points yielding
$d_{H}\Big{(}F(X)_{j},u\Big{)}\leq\max_{i}d_{H}\Big{(}W_{ij}(AX)_{i},u\Big{)}\leq\max_{j}d_{H}\Big{(}(AX)_{j},u\Big{)}\leq\beta\max_{j}d_{H}\Big{(}X_{j},u\Big{)},$ (38)
where we have used the scale-invariance property of the Hilbert distance and the fact that $\max_{i}W_{ij}>0$ for all $j$.
It remains to prove the claim. To this end, exploiting the expression of the Hilbert distance we write
$d_{H}(x_{1}+x_{2},y)$ $=\log\bigg{(}\sup_{j}\sup_{i}\frac{(x_{1})_{i}+(x_{2})_{i}}{(y)_{i}}\frac{(y)_{j}}{(x_{1})_{j}+(x_{2})_{j}}\bigg{)}$ (39)
$\leq\log\bigg{(}\sup_{i}\sup_{j}\max_{x_{1},x_{2}}\bigg{\{}\frac{(x_{1})_{i}}{(x_{1})_{j}},\frac{(x_{2})_{i}}{(x_{2})_{j}}\bigg{\}}\frac{(y)_{j}}{(y)_{i}}\bigg{)}$
$=\max_{x_{1},x_{2}}\big{\{}d_{H}(x_{1},y),d_{H}(x_{2},y)\big{\}}$
concluding the proof.
A.5 proof of Theorem 5.4
To prove the theorem we start proving that, a continuous subhomogeneous and order-preserving function $\sigma$ with an eigenvector $u$ in the cone, is not nonexpansive in Hilbert distance with respect to $u$. Formally we claim that
$d_{H}\big{(}\sigma(y),u\big{)}\leq d_{H}\big{(}y,u\big{)}\qquad\forall y\in\mathcal{K}.$ (40)
To prove it let $y\in\mathcal{K}$ and assume w.l.o.g. that $\|y\|_{1}=t>0$ and $\|u\|_{1}=1$, then
$M(y/tu)=\max_{i=1,\dots,N}\frac{y_{i}}{t(u)_{i})}\geq\frac{\|y\|_{1}}{t\|u\|}=1\qquad m(y/tu)=\min_{i=1,\dots,N}\frac{y_{i}}{t(u)_{i}}\leq\frac{\|y\|_{1}}{t\|u\|_{1}}=1.$ (41)
By definition given $x,y\in\mathcal{K}$, $m(y/x)x\leq_{\mathcal{K}}y\leq_{\mathcal{K}}M(y/x)x$. Moreover we recall that since $u$ is an eigenvector for any $t>0$ there exists $\lambda_{t}>0$ such that $\sigma(tu)=\lambda_{t}u$. Thus we use the subhomogeneity of $\sigma$ and the fact that $u$ is an eigenvector of $\sigma$ to get the following inequalities:
$m(y/tu)\lambda_{t}tu\leq_{\mathcal{K}}\sigma\big{(}m(y/tu)tu\big{)}\leq_{\mathcal{K}}f(y)\leq_{\mathcal{K}}f\big{(}M(y/tu)tu\big{)}\leq_{\mathcal{K}}M(y/x_{c})\lambda_{t}tu.$ (42)
In particular we have $m(f(y)/tu)\geq\lambda_{t}m(y/tu)$ and $M(f(y)/tu)\leq\lambda_{t}M(y/tu)$. Finally the last inequalities and the scale invariance property of $d_{H}$ yield the thesis:
$d_{H}(f(y),u)=d_{H}(f(y),tu)=\log\Big{(}\frac{M(f(y)/tu)}{m(f(y)/tu)}\Big{)}\leq\log\Big{(}\frac{M(y/tu)}{m(y/tu)}\Big{)}=d_{H}(y,tu)=d_{H}(y,u)\,.$ (43)
Now using the claim above and Lemma 5.3, we can easily conclude that
$\lim_{l\rightarrow\infty}\max_{i}d_{H}(X_{i}^{(l)},u)=0.$ (44)
Indeed by the uniform contractivity of $\{A^{(l)}\}$, if $C=\max_{i}d_{H}(X^{(0)}_{i},u)$, then there exists some $\beta_{C}<1$ such that
$\max_{i}d_{H}(X^{(l)}_{i},u)\leq\beta_{C}^{l}\max_{i}d_{H}(X^{(0)}_{i},u).$ (45)
Finally to conclude we prove that, as a consequence of (44),
$\lim_{l\rightarrow\infty}\frac{\|(I-\mathcal{P})X^{(l)}\|_{F}}{\|\mathcal{P}X^{(l)}\|_{F}}=0.$ (46)
As a consequence of the inequality:
$1\leq\mathrm{NumRank}(X^{(l)})\leq 1+\frac{\|(I-\mathcal{P})X^{(l)}\|_{F}^{2}}{\|X^{(l)}\|_{2}^{2}}\leq 1+\frac{\|(I-\mathcal{P})X^{(l)}\|_{F}^{2}}{\|\mathcal{P}X^{(l)}\|_{F}^{2}},$ (47)
we have that (46) finally yields the thesis. To prove (46), we recall from Lemma 2.5.1 in (Lemmens & Nussbaum, 2012) that for any $w$ such that $u^{T}w=c$
$\|w-\mathcal{P}w\|_{u}\leq\|\mathcal{P}w\|_{u}(e^{d_{T}(w,\mathcal{P}w)}-1),$ (48)
where $d_{T}(x,y)=\log(\max\{M(x/y),m(x/y)^{-1}\})$ and where since the dual cone of +$\mathbb{R}^{n}_{+}$ is +$\mathbb{R}^{n}_{+}$ itself, we are considering the norm induced by $u$ on the cone, i.e. $\|x\|_{u}=u^{T}x$ for any $x$ in the cone. In practice the norm induced by $u$ $\|\cdot\|_{u}$ can be defined by the Minkowki functional of the set $\Omega=\text{ConvexHull}\{\{\Omega_{1}\}\cup\{-\Omega_{1}\}\}$ where $\Omega_{1}=\{x\in\mathcal{K}\,\;u^{T}x\leq 1\}$
Then since $\|\mathcal{P}w\|_{u}=\|w\|_{u}=u^{T}w$, we have that $M(w/\mathcal{P}w)\geq 1$ and $m(w/\mathcal{P}w)\leq 1$. Thus $d_{T}(w,\mathcal{P}w)\leq d_{H}(w,\mathcal{P}w)$, yielding
$\|w-\mathcal{P}w\|_{u}\leq\|\mathcal{P}w\|_{u}(e^{d_{H}(w,\mathcal{P}w)}-1).$ (49)
From the equivalence of the norms there exists some constant $c>0$ such that we can equivalently write
$\|w-\mathcal{P}w\|_{2}\leq C\|\mathcal{P}w\|_{2}(e^{d_{H}(w,\mathcal{P}w)}-1).$ (50)
Now recall that the squared Frobenius norm of a matrix is the sum of the squared 2-norms of the its columns, so we can apply the last inequality to the matrix $X^{(l)}$ columnwise obtaining:
$\|(I-\mathcal{P})X^{(l)}\|_{F}^{2}\leq C\|\mathcal{P}X^{(l)}\|_{F}^{2}\big{(}e^{\max_{i}\{d_{H}(X^{(l)}_{i},\mathcal{P}X^{(l)}_{i})\}}-1\big{)}^{2}.$ (51)
The proof is concluded observing that by the scale invariance property of the Hilbert distance $d_{H}(X^{(l)}_{i},\mathcal{P}X^{(l)}_{i})=d_{H}(X^{(l)}_{i},u)$, yielding:
$1\leq\big{(}\mathrm{NumRank}(X^{(l)})\big{)}\leq 1+\frac{\|(I-\mathcal{P})X^{(l)}\|_{F}^{2}}{\|\mathcal{P}X^{(l)}\|_{F}^{2}}\leq 1+C\big{(}e^{\max_{i}\{d_{H}(X^{(l)}_{i},u)\}}-1\big{)}^{2}.$ (52)
A.6 proof of Proposition 5.5
We start from the homogenous case. Note that since $\psi$ is homogeneous we have that necessarily $f(t)=ct$ for all $t,c\geq 0$, this in particular means that every $u\in\mathbb{R}^{N}_{+}$ is an eigenvector of $\sigma$ with corresponding eigenvalue $\lambda_{1}=c$.
Then we can consider the subhomogeneous case. Assume that we have $u\in\mathbb{R}_{+}^{N}$ that is an eigenvector of $\sigma$ with eigenvalue $\lambda$ and $u_{i}>0$ for all $i$, then
$\psi(u_{i})=\lambda u_{i}\qquad\forall i=1,\dots,N.$ (53)
By strict subhomogeneity this means that necessarily $u$ is constant, indeed if $u_{i}>u_{j}>0$ then
$\lambda u_{j}=\psi(u_{j})=\psi\Big{(}u_{j}\frac{u_{i}}{u_{i}}\Big{)}>\frac{u_{j}}{u_{i}}\psi(u_{i})=\lambda u_{j},$ (54)
yielding a contradiction. In particular any constant vector $u$ in $\mathbb{R}_{+}^{N}$ is easily proved to be an eigenvector of $\sigma$ relative to the eigenvalue $\lambda=\|\sigma(u)\|_{1}/\|u\|_{1}=\psi(u_{i})/u_{i}$ where $u_{i}$ is any entry of $u$.
Appendix B Additional Empirical Results on Real Dataset
LReLU + GCN $r^{*}_{\mathrm{ER}}=1.450100,\quad r^{*}_{\mathrm{NR}}=1.207918$ Tanh + GCN $r^{*}_{\mathrm{ER}}=1.690520,\quad r^{*}_{\mathrm{NR}}=1.007135$
         [Uncaptioned image]          [Uncaptioned image]
LReLU + GCN + Bias $r^{*}_{\mathrm{ER}}=1.454981,\quad r^{*}_{\mathrm{NR}}=1.020319$ Tanh + GCN + Bias $r^{*}_{\mathrm{ER}}=1.858602,\quad r^{*}_{\mathrm{NR}}=1.021634$
         [Uncaptioned image]          [Uncaptioned image]
LReLU + GCN + LayerNorm $r^{*}_{\mathrm{ER}}=1.017424,\quad r^{*}_{\mathrm{NR}}=1.000101$ Tanh + GCN + LayerNorm $r^{*}_{\mathrm{ER}}=1.528688,\quad r^{*}_{\mathrm{NR}}=1.003682$
         [Uncaptioned image]          [Uncaptioned image]
LReLU + GAT $r^{*}_{\mathrm{ER}}=1.295623,\quad r^{*}_{\mathrm{NR}}=1.079250$ Tanh + GAT $r^{*}_{\mathrm{ER}}=1.002191,\quad r^{*}_{\mathrm{NR}}=1.000006$
         [Uncaptioned image]          [Uncaptioned image]
Table 3: The table showcases the behaviour of different metrics and the classification accuracies for 8 GNNs separately trained on Cora Dataset. This table is an extension of figure 1
LReLU + GCN $r^{*}_{\mathrm{ER}}=1.750563,\quad r^{*}_{\mathrm{NR}}=1.142120$ Tanh + GCN $r^{*}_{\mathrm{ER}}=1.436012,\quad r^{*}_{\mathrm{NR}}=1.003035$
         [Uncaptioned image]          [Uncaptioned image]
LReLU + GCN + Bias $r^{*}_{\mathrm{ER}}=1.526297,\quad r^{*}_{\mathrm{NR}}=1.045451$ Tanh + GCN + Bias $r^{*}_{\mathrm{ER}}=1.752850,\quad r^{*}_{\mathrm{NR}}=1.007417$
         [Uncaptioned image]          [Uncaptioned image]
LReLU + GCN + LayerNorm $r^{*}_{\mathrm{ER}}=1.006405,\quad r^{*}_{\mathrm{NR}}=1.000009$ Tanh + GCN + LayerNorm $r^{*}_{\mathrm{ER}}=1.473417,\quad r^{*}_{\mathrm{NR}}=1.002555$
         [Uncaptioned image]          [Uncaptioned image]
LReLU + GAT $r^{*}_{\mathrm{ER}}=1.541711,\quad r^{*}_{\mathrm{NR}}=1.148837$ Tanh + GAT $r^{*}_{\mathrm{ER}}=1.003797,\quad r^{*}_{\mathrm{NR}}=1.000005$
         [Uncaptioned image]          [Uncaptioned image]
Table 4: The table showcases the behaviour of different metrics and the classification accuracies for 8 GNNs separately trained on Citeseer Dataset.
LReLU + GCN $r^{*}_{\mathrm{ER}}=1.420913,\quad r^{*}_{\mathrm{NR}}=1.085279$ Tanh + GCN $r^{*}_{\mathrm{ER}}=1.838982,\quad r^{*}_{\mathrm{NR}}=1.014851$
         [Uncaptioned image]          [Uncaptioned image]
LReLU + GCN + Bias $r^{*}_{\mathrm{ER}}=1.450100,\quad r^{*}_{\mathrm{NR}}=1.207918$ Tanh + GCN + Bias $r^{*}_{\mathrm{ER}}=1.443576,\quad r^{*}_{\mathrm{NR}}=1.018361$
         [Uncaptioned image]          [Uncaptioned image]
LReLU + GCN + LayerNorm $r^{*}_{\mathrm{ER}}=1.002088,\quad r^{*}_{\mathrm{NR}}=1.000074$ Tanh + GCN + LayerNorm $r^{*}_{\mathrm{ER}}=1.668289,\quad r^{*}_{\mathrm{NR}}=1.009275$
         [Uncaptioned image]          [Uncaptioned image]
LReLU + GAT $r^{*}_{\mathrm{ER}}=1.577163\quad r^{*}_{\mathrm{NR}}=1.119213$ Tanh + GAT $r^{*}_{\mathrm{ER}}=1.016096,\quad r^{*}_{\mathrm{NR}}=1.000019$
         [Uncaptioned image]          [Uncaptioned image]
Table 5: The table showcases the behaviour of different metrics and the classification accuracies for 8 GNNs separately trained on Pubmed Dataset.

### v3

@@
 We provide theoretical support for this approach, demonstrating that the numerical rank of feature representations converges to one for a broad family of nonlinear activation functions under the assumption of nonnegative trained weights.
-To the best of our knowledge, this is the first result that proves the occurrence of oversmoothing without assumptions on the boundedness of the weight matrices.
+To the best of our knowledge, this is the first result that proves the occurrence of oversmoothing in the nonlinear setting without assumptions on the boundedness of the weight matrices.
 Along with the theoretical findings, we provide extensive numerical evaluation across diverse graph architectures.
@@
 Notably, we reveal that a significant drop in the rank aligns closely with performance degradation, even in scenarios where energy metrics remain unchanged.
-Machine Learning, ICML
 1 Introduction
-Graph neural networks (GNNs) have emerged as a powerful framework for learning representations from graph-structured data, with applications spanning knowledge retrieval and reasoning (Peng et al., 2023; Tian et al., 2022), personalised recommendation systems (Damianou et al., 2024; Peng et al., 2022), social network analysis (Fan et al., 2019), and 3D mesh classification (Shi & Rajkumar, 2020).
+Graph neural networks (GNNs) have emerged as a powerful framework for learning representations from graph-structured data, with applications spanning knowledge retrieval and reasoning [25, 39], personalised recommendation systems [7, 26], social network analysis [10], and 3D mesh classification [37].
 Central to most GNN architectures is the message-passing paradigm, where node features are iteratively aggregated from their neighbours and transformed using learned functions, such as multi-layer perceptrons or graph-attention mechanisms.
-However, the performance of message-passing-based GNNs is known to deteriorate after only a few layers, essentially placing a soft limit on the depth of GNNs.
-This is often linked to the observation of increasing similarity between learned features as GNNs deepen and is named oversmoothing (Li et al., 2018; Nt & Maehara, 2019).
-Over the years, oversmoothing in GNNs, as well as methods to alleviate it, have been studied based on the decay of some node feature similarity metrics, such as the Dirichlet energy and its variants (Oono & Suzuki, 2019; Cai & Wang, 2020; Bodnar et al., 2022; Nguyen et al., 2022; Di Giovanni et al., 2023; Wu et al., 2023; Roth & Liebig, 2023).
+However, the performance of message-passing-based GNNs is known to deteriorate after only a few layers, essentially placing a limit on their depth.
+This issue, often linked to the increasingly similar learned features as GNNs deepen, is known as oversmoothing [19, 23].
+Over the years, oversmoothing in GNNs, as well as methods to alleviate it, have been studied based on the decay of some node feature similarity metrics, such as the Dirichlet energy and its variants [24, 4, 2, 21, 8, 45, 30].
 At a high level, most of these metrics directly measure the norm of the absolute deviation from the dominant eigenspace of the message-passing matrix.
@@
 However, when nonlinear activation functions or biases are used, the dominant eigenspace may change, causing these oversmoothing metrics to fail and give false negative signals about the oversmoothing state of the learned features.
-Therefore, these metrics are often considered as providing sufficient but not necessary evidence for oversmoothing (Rusch et al., 2023a).
-Despite this, there is a considerable body of literature using these somewhat unreliable metrics as part of their evidence for non-occurrence of oversmoothing in GNNs (Zhou et al., 2021; Chen et al., 2022; Rusch et al., 2022; Wang et al., 2022; Maskey et al., 2023; Nguyen et al., 2023; Rusch et al., 2023b; Epping et al., 2024; Scholkemper et al., 2024; Wang & Cho, 2024).
+Therefore, these metrics are often considered as sufficient but not necessary evidences for oversmoothing [33].
+Despite this, there is a considerable body of literature using these unreliable metrics as their evidences for non-occurrence of oversmoothing in GNNs [47, 6, 35, 44, 20, 22, 34, 9, 36, 43, 29, 42].
 As we show in Section 6, the performance degradation of GNNs trained on real datasets often happens well before any noticeable decay in these oversmoothing metrics can be observed.
 Most empirical studies in the literature that observe the decay of the Dirichlet-energy-like metrics are conducted over the layers of the same very deep untrained (with randomly sampled weights) or effectively untrained [^1] Deep networks (with, say, over 100 layers) that are trained but whose loss and accuracy remain far from acceptable.
-GNNs (Rusch et al., 2022; Wang et al., 2022; Rusch et al., 2023b; Wu et al., 2023), where the decay of the metrics is likely driven by the small weight initializations reducing all features to zero.
+GNNs [44, 35, 34, 45, 29, 42], where the decay of the metrics is likely driven by the small weight initializations reducing all features to zero.
 Instead, we observe that when GNNs of different depths are separately trained, these metrics do not correlate well with their performance degradation.
@@
 Overall, the main contributions of this paper are as follows:
-  • We review popular oversmoothing metrics in the current literature and provide a novel perspective on their theoretical analysis based on nonlinear activation eigenvectors.
-  • Based on this analysis and observations on the linear case, we propose the rank as a better metric for quantifying oversmoothing, and thereby re-defining oversmoothing in GNNs as the convergence towards a low-rank matrix rather than to a matrix of exactly rank one.
+  • We review popular oversmoothing metrics in the current literature and simplify their theoretical analysis from a novel perspective of nonlinear activation eigenvectors.
+  • We propose the rank as a better metric for quantifying oversmoothing, and thereby re-defining oversmoothing in GNNs as the convergence towards a low-rank matrix rather than to a matrix of exactly rank one.
   • We prove the convergence of the numerical rank towards one for linear GNNs and nonlinear GNNs where the eigenvector of the message-passing matrix is also the eigenvector of the nonlinear activation function under the assumption of non-negative weights.
-To our knowledge, this is the first theoretical result proving that oversmoothing may occur independently of the weights’ magnitude with nonlinear activation functions.
+To the best of our knowledge, this is the first theoretical result proving that oversmoothing may occur independently of the weights’ magnitude with nonlinear activation functions.
   • We provide extensive numerical evidence that continuous rank relaxation functions better capture oversmoothing than commonly used Dirichlet-like metrics.
@@
 Let $\mathcal{G}=(\mathcal{V},\mathcal{E})$ be an undirected graph with $\mathcal{V}$ denoting its set of vertices and $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$ its set of edges.
-Let $\widetilde{A}\in\mathbb{R}^{N\times N}$ be the unweighted adjacency matrix with $N=|\mathcal{V}|$ being the total number of nodes, $|\mathcal{E}|$ being the total number of edges of $\mathcal{G}$ and $A$ the corresponding symmetric adjacency matrix normalized by the node degrees:
-$A=\widetilde{D}^{-1/2}\widetilde{A}\widetilde{D}^{-1/2},$ (1)
-where $\widetilde{D}=D+I$, $D$ is the diagonal degree matrix of the graph $\mathcal{G}$, and $I$ is the identity matrix.
+Let $\widetilde{A}\in\mathbb{R}^{N\times N}$ be the unweighted adjacency matrix with $N=|\mathcal{V}|$ being the total number of nodes, $|\mathcal{E}|$ being the total number of edges of $\mathcal{G}$ and $A$ the corresponding symmetric adjacency matrix normalized by the node degrees: $A=\widetilde{D}^{-1/2}\widetilde{A}\widetilde{D}^{-1/2}$, where $\widetilde{D}=D+I$, $D$ is the diagonal degree matrix of the graph $\mathcal{G}$, and $I$ is the identity matrix.
 The rows of the feature matrix $X\in\mathbb{R}^{N\times d}$ are the concatenation of the $d$-dimensional feature vectors of all nodes in the graph.
-At each layer $l$, the GCN updates the node features as follows
-$X^{(l+1)}=\sigma(AX^{(l)}W^{(l)})$ (2)
-where $\sigma$ is a nonlinear activation function, applied componentwise, and $W^{(l)}$ is a trainable weight matrix.
+At each layer $l$, the node feature update of Graph Convolutional Network (GCN) [17] follows $X^{(l+1)}=\sigma(AX^{(l)}W^{(l)})$ where $\sigma$ is a nonlinear activation function, applied component-wise, and $W^{(l)}$ is a trainable weight matrix.
 2.2 Graph Attention Network
-While GCNs use a fixed normalized adjacency matrix to perform graph convolutions at each layer, Graph Attention Networks (GATs) (Veličković et al., 2017; Brody et al., 2021) perform graph convolution through a layer-dependent message-passing matrix $A^{(l)}$ learned through an attention mechanism as follows
-$A^{(l)}_{ij}=\text{softmax}_{j}(\sigma_{a}(p_{1}^{(l)\top}W^{(l)\top}X_{i}+p_{2}^{(l)\top}W^{(l)\top}X_{j}))$ (3)
-where $p_{i}^{(l)}$ are learnable parameter vectors, $X_{i},X_{j}$ denote the feature of the $i$th and $j$th nodes respectively, the activation $\sigma_{a}$ is typically chosen to be LeakyReLU, and $\text{softmax}_{j}$ corresponds to the row-wise normalization
-$\text{softmax}_{j}(A_{ij})=\frac{\exp(A_{ij})}{\sum_{j^{\prime}}\exp(A_{ij^{\prime}})}.$ (4)
-The corresponding feature update is
-$X^{(l+1)}=\sigma(A^{(l)}X^{(l)}W^{(l)})\,.$ (5)
+Graph Attention Networks (GATs) [41, 3] perform graph convolution via a layer-dependent message-passing matrix $A^{(l)}$ learned through an attention mechanism $A^{(l)}_{ij}=\text{softmax}_{j}(\sigma_{a}(p_{1}^{(l)\top}W^{(l)\top}X_{i,:}+p_{2}^{(l)\top}W^{(l)\top}X_{j,:}))$ where $p_{i}^{(l)}$ are learnable parameter vectors, $X_{i,:},X_{j,:}$ denote the feature of the $i$-th and $j$th nodes respectively, the activation $\sigma_{a}$ is typically chosen to be LeakyReLU, and $\text{softmax}_{j}$ corresponds to the row-wise normalization $\text{softmax}_{j}(A_{ij})=\exp(A_{ij})/\sum_{j^{\prime}}\exp(A_{ij^{\prime}})$.
+The corresponding feature update is $X^{(l+1)}=\sigma(A^{(l)}X^{(l)}W^{(l)})\,.$
 3 Oversmoothing
@@
 Some works define oversmoothing more rigorously as the alignment of all feature vectors with each other.
-This definition is motivated by the behaviour of a linear GCN:
-$X^{(l+1)}=A\cdots AX^{(0)}W^{(0)}\dots W^{(l)}.$ (6)
-Indeed if $\widetilde{A}$ is the adjacency matrix of a fully connected graph $\mathcal{G}$, $A$ will have a spectral radius equal to $1$ with multiplicity $1$, and $A^{l}$ will thus converge towards the eigenspace spanned by the dominant eigenvector.
-Precisely, we have
-$A^{l}\to uv^{\top}\quad\text{as}\quad l\rightarrow\infty,$ (7)
-where $Au=u$, and $A^{\top}v=v$, see e.g. (Tudisco et al., 2015).
+This definition is motivated by the behaviour of a linear GCN: $X^{(l+1)}=A\cdots AX^{(0)}W^{(0)}\dots W^{(l)}$.
+Indeed, if $\widetilde{A}$ is the adjacency matrix of a fully connected graph, $A$ will have spectral radius equal to $1$ with multiplicity $1$, and $A^{l}$ will converge toward the eigenspace spanned by the dominant eigenvector.
+Precisely, $A^{l}\to uv^{\top}$ as $l\rightarrow\infty$, where $Au=u$ and $A^{\top}v=v$ [40].
 As a consequence, if the product of the weight matrices $W^{(0)}\cdots W^{(l)}$ converges in the limit $l\rightarrow\infty$, then the features degenerate to a matrix having rank at most one, where all the features are aligned with the dominant eigenvector $u$.
-Mathematically, if we assume $u$ to be such that $\|u\|=1$, this alignment can be expressed by stating that the difference between the features and their projection onto $u$, given by $\|X^{(l)}-uu^{T}X^{(l)}\|$, converges to zero.
+Mathematically, if we assume $u$ to be such that $\|u\|=1$, this alignment can be expressed by stating that the difference between the features and their projection onto $u$, given by $\|X^{(l)}-uu^{\top}X^{(l)}\|$, converges to zero.
 3.1 Existing Oversmoothing Metrics
-Motivated by the discussion about the linear case, oversmoothing is thus quantified and analysed in terms of the convergence of some node similarity metrics towards zero.
+Motivated by the discussion about the linear case, oversmoothing is thus typically quantified and analysed in terms of the convergence of some node similarity metrics towards zero.
 In particular, in most cases, it is measured exactly by the alignment of the features with the dominant eigenvector of the matrix $A$.
-The most prominent metric that has been used to quantify oversmoothing is the Dirichlet energy, which measures the norm of the difference between the degree-normalized neighbouring node features (Cai & Wang, 2020; Rusch et al., 2023a)
-$E_{\mathrm{Dir}}(X)=\sum_{(i,j)\in\mathcal{E}}\left\|\frac{X_{i}}{u_{i}}-\frac{X_{j}}{u_{j}}\right\|^{2}_{2},$ (8)
-where $u_{i}$ is the $i$-th entry of the dominant eigenvector of the message-passing matrix in (1).
+The most prominent metric that has been used to quantify oversmoothing is the Dirichlet energy, which measures the norm of the difference between the degree-normalized neighbouring node features [4, 33]
+$E_{\mathrm{Dir}}(X)=\sum_{(i,j)\in\mathcal{E}}\left\|\frac{X_{i,:}}{u_{i}}-\frac{X_{j,:}}{u_{j}}\right\|^{2}_{2},$ (1)
+where $u_{i}$ is the $i$-th entry of the dominant eigenvector of the message-passing matrix.
 It thus immediately follows from our discussion on the linear setting that $E_{\mathrm{Dir}}(X^{(l)})$ converges to zero as $l\to\infty$ for a linear GCN with converging weights product $W^{(0)}\cdots W^{(l)}$.
 This intuition suggests that a similar behaviour may occur for “smooth-enough” nonlinearities.
-In particular, in the case of a GCN, the dominant eigenvector $u$ is defined by $u_{i}=\sqrt{1+d_{i}}$ and (Cai & Wang, 2020) have proved that, using LeakyReLU activation functions $E_{\mathrm{Dir}}(X^{(l+1)})\leq s_{l}\bar{\lambda}E_{\mathrm{Dir}}(X^{(l)})$, where $s_{l}=\|W^{(l)}\|_{2}$ is the largest singular value of the weight matrix $W^{(l)}$, and $\bar{\lambda}=(1-\min_{i}\lambda_{i})^{2}$, where $\lambda_{i}\in(0,2]$ varies among the nonzero eigenvalues of the normalized graph Laplacian $\widetilde{\Delta}=I-A=I-\widetilde{D}^{-\frac{1}{2}}\widetilde{A}\widetilde{D}^{-\frac{1}{2}}$.
+In particular, in the case of a GCN, the dominant eigenvector $u$ is defined by $u_{i}=\sqrt{1+d_{i}}$ and [4] have proved that, using LeakyReLU activation functions, it holds $E_{\mathrm{Dir}}(X^{(l+1)})\leq s_{l}\bar{\lambda}E_{\mathrm{Dir}}(X^{(l)})$, where $s_{l}=\|W^{(l)}\|_{2}$ is the largest singular value of the weight matrix $W^{(l)}$, and $\bar{\lambda}=(1-\min_{i}\lambda_{i})^{2}$, where $\lambda_{i}\in(0,2]$ varies among the nonzero eigenvalues of the normalized graph Laplacian $\widetilde{\Delta}=I-A=I-\widetilde{D}^{-\frac{1}{2}}\widetilde{A}\widetilde{D}^{-\frac{1}{2}}$.
 Similarly, in the case of GATs, the matrices $A_{i}$ are all row stochastic, meaning that $u_{i}=1$ for all $i$.
-In this case, in (Wu et al., 2023), it has been proved that whenever the product of the entry-wise absolute value of the weights is bounded, that is $\|\Pi_{k=1}^{\infty}|W^{(k)}|\|<\infty$, then the following variant of the Dirichlet energy decays to zero
+In this case, it has been proved that whenever the product of the entry-wise absolute value of the weights is bounded, that is $\|\Pi_{k=1}^{\infty}|W^{(k)}|\|<\infty$, then the following variant of the Dirichlet energy decays to zero [45]
 $E_{\mathrm{Proj}}(X)=\|X-\mathcal{P}X\|^{2}_{F}$ (2)
@@
 Definition 3.1.
-We say that a vector $u\in\mathbb{R}^{N}\setminus\{0\}$ is an eigenvector of the (nonlinear) activation function $\sigma:\mathbb{R}^{N}\to\mathbb{R}^{n}$ if for any $t\in\mathbb{R}\setminus\{0\}$, there exists $\mu_{t}$ such that $\sigma(tu)=\mu_{t}u$.
+We say that a vector $u\in\mathbb{R}^{N}\setminus\{0\}$ is an eigenvector of the (nonlinear) activation function $\sigma:\mathbb{R}^{N}\to\mathbb{R}^{n}$ if for any $t\in\mathbb{R}\setminus\{0\}$, there exists $\mu_{t}\in\mathbb{R}$ such that $\sigma(tu)=\mu_{t}u$.
 With this definition, we can now provide a unifying characterization of message-passing operators $A^{(l)}$ and activation functions $\sigma$ that guarantee the convergence of the Dirichlet-like energy metric $E_{\mathrm{Proj}}$ to zero for the feature representation sequence defined by $X^{(l+1)}=\sigma(A^{(l)}X^{(l)}W^{(l)})$.
@@
 This assumption recurs throughout our theoretical analysis, aligning with existing results in the literature.
-For example, (a) the proof by Cai & Wang (2020) applies to GCNs with LeakyReLU, where the dominant eigenvector of $A$ is nonnegative by the Perron-Frobenius theorem and, therefore, an eigenvector of LeakyReLU; and (b) the proof by Wu et al.
-(2023) holds for stochastic message-passing matrices $A^{(l)}$, which inherently share a common dominant eigenvector with constant entries, that is an eigenvector for any pointwise nonlinear activation function.
-Lemma 3.2.
-Let $X^{(l+1)}=\sigma(A^{(l)}X^{(l)}W^{(l)})$ be the $l$-th layer of a GNN.
-Assume that $A^{(l)}$ has a simple dominant eigenvalue with corresponding eigenvector $u$ that is also an eigenvector of $\sigma$.
-If $\sigma$ is $1$-Lipschitz, namely $\|\sigma(x)-\sigma(y)\|\leq\|x-y\|$ for any $x,y$, then
-$\|(I-\mathcal{P})X^{(l+1)}\|_{F}\leq\|(I-\mathcal{P})A^{(l)}X^{(l)}W^{(l)}\|_{F},$
-where $\mathcal{P}=uu^{T}/\|u\|^{2}$ is the projection matrix on the linear space spanned by $u$.
-The following theorem follows as a consequence.
+For example, (a) the proof by [4] applies to GCNs with LeakyReLU, where the dominant eigenvector of $A$ is nonnegative by the Perron-Frobenius theorem and, therefore, an eigenvector of LeakyReLU; and (b) the proof by [45] holds for stochastic message-passing matrices $A^{(l)}$, which inherently share a common dominant eigenvector with constant entries, that is an eigenvector for any pointwise nonlinear activation function.
 Theorem 3.2.
-With the same notation of Lemma 3.2, if
-$\lim_{L\to\infty}\prod_{l=0}^{L}\|(I-\mathcal{P})A^{(l)}\|_{2}\|W^{(l)}\|_{2}=0,$
-then $E_{\mathrm{Proj}}(X^{(l)})\to 0$ as $l\to\infty$.
-Proof.
-The proof follows observing that in the decomposition $(I-\mathcal{P})A^{(l)}=(I-\mathcal{P})A^{(l)}\mathcal{P}+(I-\mathcal{P})A^{(l)}(I-\mathcal{P})$ the matrix $(I-\mathcal{P})A^{(l)}\mathcal{P}$ is zero because $A^{(l)}u=\lambda_{1}^{l}u$ for any $l$.
-Thus
-$(I-\mathcal{P})A^{(l)}=(I-\mathcal{P})A^{(l)}(I-\mathcal{P}),$
-and, from Lemma 3.2 and the inequality $\|AB\|_{F}\leq\|A\|_{2}\|B\|_{F}$, we have $\|(I-\mathcal{P})X^{(L)}\|_{F}\leq\Big{(}\Pi_{l=0}^{L-1}\|(I-\mathcal{P})A^{(l)}\|_{2}\|W^{(l)}\|_{2}\Big{)}\|X^{(0)}\|_{F}$.
-∎
+Let $X^{(l+1)}=\sigma(A^{(l)}X^{(l)}W^{(l)})$, $l=1,\dots,L$, be a GNN such that $u$ is the dominant eigenvector of $A_{l}$ for any $l$ and is also an eigenvector of the activation function $\sigma$.
+If $\sigma$ is $1$-Lipschitz, namely $\|\sigma(x)-\sigma(y)\|\leq\|x-y\|$ for any $x,y$, and $\lim_{L\to\infty}\prod_{l=0}^{L}\|(I-\mathcal{P})A^{(l)}\|_{2}\|W^{(l)}\|_{2}=0$, then
+$E_{\mathrm{Proj}}(X^{(L)})\to 0\quad\text{ as }\quad L\to\infty.$
 Note, in particular, that in the case of GCNs, the matrix $A^{(l)}=A$ is symmetric, and thus $\|I-\mathcal{P}A^{(l)}\|_{2}=\lambda_{2}$.
-Therefore, when $\sigma=\text{LeakyReLU}$, we obtain the result by Cai & Wang (2020) as convergence to zero is guaranteed if $\|W^{(l)}\|_{2}\leq\lambda_{2}$.
+Therefore, when $\sigma=\text{LeakyReLU}$, we obtain the result by [4] as convergence to zero is guaranteed if $\|W^{(l)}\|_{2}\leq\lambda_{2}$.
 Recall that, as discussed above, the choice $\sigma=\text{LeakyReLU}$ satisfies our eigenvector assumption since $u\geq 0$ by the Perron-Frobenius theorem, and thus LeakyReLU$(tu)=\alpha tu$ with $\alpha$ depending only on the sign of $t$.
 Similarly, in the case of GATs, the matrices $A^{(l)}$ are stochastic for all $l$, implying that $u=\mathbbm{1}$ is the constant vector with $(u)_{i}=1$ for all $i$.
-If $\sigma=\otimes\psi$ is a nonlinear activation function acting entrywise through $\psi$, then $\sigma(t\mathbbm{1})=\psi(t)\mathbbm{1}$.
+If $\sigma=\otimes\psi$ is a nonlinear activation function acting entry-wise through $\psi$, then $\sigma(t\mathbbm{1})=\psi(t)\mathbbm{1}$.
 Therefore, Theorem 3.2 implies that if the weights are sufficiently small, the features align independently of the activation function used.
-This is consistent with the results in (Wu et al., 2023).
-However, we note that the bounds on the weights required by Theorem 3.3 and those in (Wu et al., 2023) on the weights $W^{(l)}$ are not identical, and it is unclear which of the two is more significant.
+This is consistent with the results in [45].
+However, we note that the bounds on the weights required by Theorem 3.2 and those in [45] on the weights $W^{(l)}$ are not identical, and it is unclear which of the two is more significant.
 Nonetheless, in both cases, having bounded weights along with any $1$-Lipschitz pointwise activation function is a sufficient condition for observing oversmoothing via $E_{\mathrm{Proj}}$ in GATs.
-In addition to offering a different and unifying theoretical perspective on the results in (Cai & Wang, 2020; Wu et al., 2023), we highlight the simplicity of our eigenvector-based proof, which we hope provides added clarity on this phenomenon.
+In addition to offering a different and unifying theoretical perspective on the results in [4, 45], we highlight the simplicity of our eigenvector-based proof, which we hope provides added clarity on this phenomenon.
 4 Energy-like Metrics: What Can Go Wrong
@@
 This aligns with our intuition and what occurs in the linear case.
-Indeed, for a linear GCN, even when the features $X^{(l)}$ grow to infinity as $l\to\infty$, one observes that $X^{(l)}$ becomes dominated by the dominant eigenspace of $A$, even for finite and possibly small values of $l$, depending on the spectral gap of the graph.
+Indeed, for a linear GCN, even when the features $X^{(l)}$ grow to infinity as $l\to\infty$, one observes that $X^{(l)}$ is dominated by the dominant eigenspace of $A$, even for finite and possibly small values of $l$, depending on the spectral gap of the graph.
 More precisely, the following theorem holds:
 Theorem 4.1.
-Let $X^{(l+1)}=AX^{(l)}W^{(l)}$ be a linear GCN.
-Let $\lambda_{1},\lambda_{2}$ be the largest and second-largest eigenvalues (in modulus) of $A$, respectively.
-Assume the weights $\{W^{(l)}\}_{l=1}^{\infty}$ are randomly sampled from i.i.d.
-random variables with distribution $\nu$ such that
-[…]
 […]
 $\lim_{l\to\infty}\frac{\|(I-\mathcal{P})X^{(l)}\|_{F}}{\|\mathcal{P}X^{(l)}\|_{F}}=0$
-with a linear rate of convergence $O(|\lambda_{2}/\lambda_{1}|^{l})$.
-In particular, the theorem above implies that for a large spectral gap $|\lambda_{2}/\lambda_{1}|\ll 1$, $X^{(l)}$ is predominantly of rank one, namely
-$X^{(l)}=\lambda_{1}^{l}\Big{(}uv^{T}+R(l)\Big{)}$
-for some $v$, with $R(l)\sim O(|\lambda_{2}/\lambda_{1}|)^{l}$ and thus converging to 00 as $l\to\infty$.
+with a linear rate of convergence $|\lambda_{2}/\lambda_{1}|$.
+In particular, the theorem above implies that for a large spectral gap $|\lambda_{2}/\lambda_{1}|\ll 1$, $X^{(l)}$ is predominantly of rank one, namely $X^{(l)}=\lambda_{1}^{l}\Big{(}uv^{\top}+R(l)\Big{)}$ for some $v$, with $R(l)\sim O(|\lambda_{2}/\lambda_{1}|)^{l}$ and thus converging to 00 as $l\to\infty$.
 This results in weakly expressive feature representations, independently of the magnitude of the feature weights.
-This phenomenon can be effectively captured by measuring the rank of $X^{(l)}$, whereas Dirichlet-like energy measures may fail to detect it.
+This phenomenon can be effectively captured by measuring the rank of $X^{(l)}$, whereas Dirichlet-like energy measures may fail to detect it, this is e.g.
+the case when $\lambda_{1}>1$ having $X^{(l)}\rightarrow\infty$.
 Another important limitation of these metrics is their dependence on a specific, known dominant eigenspace, which must either be explicitly known or computed in advance.
@@
 5 The Rank as a Measure of Oversmoothing
-Inspired by the behaviour observed in the linear case, we argue that measuring the rank of feature representations provides a more effective way to quantify oversmoothing, in alignment with recent work on oversmoothing (Guo et al., 2023).
+Inspired by the behaviour observed in the linear case, we argue that measuring the rank of feature representations provides a more effective way to quantify oversmoothing, in alignment with recent work on oversmoothing [14].
 However, since the rank of a matrix is defined as the number of nonzero singular values, it is a discrete function and thus not suitable as a measure.
 A viable alternative is to use a continuous relaxation that closely approximates the rank itself.
-Examples of possible continuous approximations of the rank include the numerical rank, the stable rank, and the effective rank (Roy & Vetterli, 2007; Rudelson & Vershynin, 2006; Arora et al., 2019).
+Examples of possible continuous approximations of the rank include the numerical rank, the stable rank, and the effective rank [31, 32, 1].
 The stable rank is defined as $\text{StableRank}(X)=\|X\|_{*}^{2}/\|X\|_{F}^{2}$, where $\|X\|_{*}=\sum_{i}\sigma_{i}$ is the nuclear norm.
@@
 Finally, given the singular values $\sigma_{1}>\sigma_{2}>\dots>\sigma_{\min\{N,d\}}$ of $X$, the effective rank is defined as
-$\text{Erank}(X)=\exp\left(-\textstyle{\sum_{k}p_{k}\log p_{k}}\right),$ (10)
-where $p_{k}=\sigma_{k}/\sum_{i}\sigma_{i}$ is the $k$-th normalized singular value.
-These rank relaxation measures exhibit similar empirical behaviour as shown by our numerical evaluation in Section 6.
-In practice, measuring oversmoothing in terms of a continuous approximation of the rank is a reasonable approach that helps address the limitations of Dirichlet-like measures.
-Specifically, it offers the following advantages: (a) it is scale-invariant, meaning it remains informative even when the feature matrix converges to zero or explodes to infinity; (b) it does not rely on a fixed, predetermined eigenspace but instead captures convergence of the feature matrix toward an arbitrary lower-dimensional subspace; (c) it allows for the detection of oversmoothing in shallow networks without requiring exact convergence to rank one—since a small value of, say, the effective rank, directly implies that the feature representations are low-rank, which in turn suggests a potentially suboptimal network architecture.
-In Table 1, we present a toy example illustrating that classical oversmoothing metrics fail to correctly capture oversmoothing unless the features are perfectly aligned.
+$\text{Erank}(X)=\exp\left(-\textstyle{\sum_{k}p_{k}\log p_{k}}\right),$
+where $p_{k}=\sigma_{k}/\sum_{i}\sigma_{i}$ is the $k$-th normalized singular value. These rank relaxation measures exhibit similar empirical behaviour as shown in Section 6.
+In practice, measuring oversmoothing in terms of a continuous approximation of the rank is a reasonable approach that helps addressing the limitations of Dirichlet-like measures.
+Specifically, it offers the following advantages: (a) it is scale-invariant, meaning it remains informative even when the feature matrix converges to zero or explodes to infinity; (b) it does not rely on a fixed, predetermined eigenspace but instead captures convergence of the feature matrix toward an arbitrary lower-dimensional subspace; (c) it allows for the detection of oversmoothing in shallow networks without requiring exact convergence to rank one.
+A small value of the effective rank directly implies that the feature representations are low-rank, suggesting a potentially suboptimal network architecture.
+In Figure 1, we present a toy example illustrating that classical oversmoothing metrics fail to correctly capture oversmoothing unless the features are perfectly aligned.
 This observation implies that these metrics can quantify oversmoothing only when the rank of the feature matrix converges exactly to one.
@@
 In such cases, classical metrics remain roughly constant, whereas the rank decreases, coinciding with a sharp drop in GNN accuracy.
-Table 1: A set of toy scenarios depicting the behaviour of oversmoothing metrics.
-Each plot contains 50 nodes, each with two features plotted as the x and y axis.
-# 1: all features are of the same value; # 2: all features are assumed to align with the same vector exactly; # 3: shifts one (red) point in # 2 in y-direction; # 4: samples all points from a uniform distribution.
-MAD (see Section 6) and $E_{\mathrm{Dir}}$ give false negative signals in # 3 while the features are considered oversmoothing by definition.
-$E_{\mathrm{Proj}}$ can hardly differentiate between # 3 and # 4, and is thus not robust in quantifying oversmoothing.
-To compute $E_{\mathrm{Proj}}$ and $E_{\mathrm{Dir}}$ the first feature was considered in place of $u$ in (8) and (9).
 # 1 # 2 # 3 # 4
-[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
+Refer to caption Refer to caption Refer to caption Refer to caption
 $E_{\mathrm{Dir}}$ 0 0 13.25 77.78
@@
 Erank 1 1 1.36 1.99
+Figure 1: Toy scenarios depicting the behaviour of oversmoothing metrics.
+Each plot contains 50 nodes, each with two features plotted on the x-y axis.
+The features are: # 1 of the same value; # 2 perfectly aligned with the same vector; # 3 aligned to the same vector except for one (red) point; # 4 sampled from a uniform distribution.
+MAD (Sec.
+6) and $E_{\mathrm{Dir}}$ give false negative signals in # 3 although features are oversmoothing by definition.
+$E_{\mathrm{Proj}}$ can hardly differentiate between # 3 and # 4, and is thus not robust in quantifying oversmoothing.
+To compute $E_{\mathrm{Proj}}$ and $E_{\mathrm{Dir}}$, the first feature was considered in place of $u$ in (1) and (2).
 5.1 Theoretical Analysis of Rank Decay
-In this section, we provide an analytical study proving the decrease of the numerical rank for a broad class of graph neural network architectures under the assumption that the weight matrices are entrywise nonnegative.
+In this section, we provide an analytical study proving the decrease of the numerical rank for a broad class of graph neural network architectures under the assumption that the weight matrices are entry-wise nonnegative.
 While this is a somewhat restrictive setting, our result is the first theoretical proof that oversmoothing occurs independently of the weight (and thus feature) magnitude.
@@
 Let $u$ be the dominant eigenvector of $A$ corresponding to $\lambda_{1}$ and satisfying $\|u\|=1$.
-Consider the projection matrix $\mathcal{P}=uu^{\top}$.
+Consider the projection $\mathcal{P}=uu^{\top}$.
 Given a matrix $X$, we can decompose it as $X=\mathcal{P}X+(I-\mathcal{P})X$.
@@
 Moreover, since $\mathcal{P}X$ and $(I-\mathcal{P})X$ are orthogonal with respect to the Frobenius inner product, we have $\|\mathcal{P}X+(I-\mathcal{P})X\|_{F}^{2}=\|\mathcal{P}X\|_{F}^{2}+\|(I-\mathcal{P})X\|_{F}^{2}$.
-Thus, we obtain the following bound for the numerical rank of $X$:
-$\text{NumRank}(X)$ $=\frac{\|\mathcal{P}X+(I-\mathcal{P})X\|_{F}^{2}}{\|X\|_{2}^{2}}$ (12)
-$=\frac{\|\mathcal{P}X\|_{F}^{2}+\|(I-\mathcal{P})X\|_{F}^{2}}{\|X\|_{2}^{2}}$
-$\leq 1+\frac{\|(I-\mathcal{P})X\|_{F}^{2}}{\|X\|_{2}^{2}},$
-where we used (11) and the fact that $\|\mathcal{P}X\|_{F}=\|\mathcal{P}X\|_{2}$ since $\mathcal{P}X$ is a rank-one matrix.
+Thus, we obtain the following bound:
+$\text{NumRank}(X)=\frac{\|\mathcal{P}X+(I-\mathcal{P})X\|_{F}^{2}}{\|X\|_{2}^{2}}=\frac{\|\mathcal{P}X\|_{F}^{2}+\|(I-\mathcal{P})X\|_{F}^{2}}{\|X\|_{2}^{2}}\leq 1+\frac{\|(I-\mathcal{P})X\|_{F}^{2}}{\|X\|_{2}^{2}},$ (4)
 The above inequality, together with Theorem 4.1, allows us to establish the convergence of the numerical rank for linear networks.
@@
 Extending the result above to general GNNs with nonlinear activation functions is highly nontrivial.
-In the next section, we present our main theoretical result, which generalizes this analysis to a broader class of GNNs under the assumption of nonnegative weights.
-5.2 The Nonnegative Setting
-To study the case of networks with nonnegative weights, we make use of tools from the nonlinear Perron-Frobenius theory; we refer to (Lemmens & Nussbaum, 2012; Gautier et al., 2023) and the reference therein for further details.
-We assume all the intermediate features to be in the positive open cone $\mathcal{K}:=\mathbb{R}_{+}^{N}=\{x\in\mathbb{R}^{N}\>|\>x_{i}>0\;\forall i=1,\dots,N\}$.
+In the next subsection, we present our main theoretical result, which generalizes this analysis to a broader class of GNNs under the assumption of nonnegative weights.
+5.1.1 The Nonlinear Case
+To study the case of networks with nonlinear activations, we make use of tools from the nonlinear Perron-Frobenius theory; we refer to [18, 12] and the reference therein for further details.
+Assume all the intermediate features of the network to be in the positive open cone $\mathcal{K}:=\mathbb{R}_{+}^{N}=\{x\in\mathbb{R}^{N}\>|\>x_{i}>0\;\forall i=1,\dots,N\}$.
 On $\mathcal{K}$, it is possible to introduce the partial ordering
-$x\leq_{\mathcal{K}}y\>(x\ll_{\mathcal{K}}y)\quad\text{iff}\quad y-x\in\bar{\mathcal{K}}\>(y-x\in\mathcal{K})$ (13)
-where $\bar{\mathcal{K}}$ denotes the nonnegative closed cone.
+$x\leq_{\mathcal{K}}y\>(x\ll_{\mathcal{K}}y)\quad\text{iff}\quad y-x\in\bar{\mathcal{K}}\>(y-x\in\mathcal{K})$
+where $\bar{\mathcal{K}}$ denotes the closed cone of nonnegative vectors.
 Given two points $x,y\in\mathcal{K}$ we write $M(x/y)=\max_{i}x_{i}/y_{i}$ and $m(x/y)=\min_{i}x_{i}/y_{i}$.
-Then, we can define the following Hilbert distance between any two points $x,y\in\mathcal{K}$,
-$d_{H}(x,y)=\log(M(x/y)/m(x/y)).$
-Note that $d_{H}$ is not a distance on $\mathcal{K}$, indeed $d_{H}(\alpha x,\beta y)=d_{H}(x,y)$ for any $x,y\in\mathcal{K}$ and $\alpha,\beta>0$.
+Then, we can define the following Hilbert distance between any two points $x,y\in\mathcal{K}$, $d_{H}(x,y)=\log(M(x/y)/m(x/y)).$ Note that $d_{H}$ is not a distance on $\mathcal{K}$, indeed $d_{H}(\alpha x,\beta y)=d_{H}(x,y)$ for any $x,y\in\mathcal{K}$ and $\alpha,\beta>0$.
 However it is a distance up to scaling; that is, it becomes a concrete distance whenever we restrict ourselves to a slice of the cone.
-Now, it is well-known that any matrix $A$ such that $Ax\in\mathcal{K}$ for any $x\in\mathcal{K}$ is non-expansive with respect to the Hilbert distance, i.e.
+It is well-known that any matrix $A$ such that $Ax\in\mathcal{K}$ for any $x\in\mathcal{K}$ is non-expansive with respect to the Hilbert distance, i.e.
 $d_{H}(Ax,Ay)\leq\beta d_{H}(x,y)\qquad\forall x,y\in\mathcal{K}$ (5)
@@
 In the next definition, we include all the nonnegative matrices that have the same behaviour as the positive matrices but only on a neighbourhood of a positive vector in $\mathcal{K}$.
-Recall that any irreducible nonnegative matrix $A$ has a positive eigenvector $u$ corresponding to its dominant eigenvalue.
+To this end, we recall that any irreducible nonnegative matrix $A$ has a positive eigenvector $u$ corresponding to its dominant eigenvalue.
 Definition 5.2.
-A family of nonnegative irreducible matrices (possibly one single matrix) $\{A^{(l)}\}$ with the same dominant eigenvector $u>0$ is uniformly contractive with respect to $u$ if for any $C>0$ there exists some $\beta_{C}<1$ such that $d_{H}(A^{(l)}x,u)\leq\beta_{C}d_{H}(x,u)$ for all $l$ and $x\in\mathcal{K}$ such that $d_{H}(x,u)\leq C$.
-The above definition is significantly weaker than asking for $A$ to be strictly positive. We show this with an illustrative example in Section A.3.
-Consider now first the case of a linear function that can be represented in the form
-$F(X)=AXW,$ (15)
-where $X$ is a positive $N\times d$ matrix, $A$ and $W$ are nonnegative $N\times N$ and $d\times d$ matrices, respectively.
+A family of nonnegative irreducible matrices $\{A^{(l)}\}$ with the same dominant eigenvector $u>0$ is uniformly contractive with respect to $u$ if for any $C>0$ there exists some $\beta_{C}<1$ such that $d_{H}(A^{(l)}x,u)\leq\beta_{C}d_{H}(x,u)$ for all $l$ and $x\in\mathcal{K}$ such that $d_{H}(x,u)\leq C$
+Note that the above definition is significantly weaker than asking for $A$ to be strictly positive. We show this with an illustrative example in Section A.3.
+Consider now first the case of a linear function that can be represented in the form $F(X)=AXW$, where $X$ is a positive $N\times d$ matrix, $A$ and $W$ are nonnegative $N\times N$ and $d\times d$ matrices, respectively.
 Then, under mild assumptions, we can prove that the columns of $F(X)$ are closer to the dominant eigenvector $u$ of $A$ as compared to the columns of $X$.
@@
 Let $F(X)=AXW$ with $A$ nonnegative and irreducible with dominant eigenvector $u\in\mathcal{K}$.
-Assume also $X$ to be striclty positve and $W$ nonnegative with $\min_{j}\max_{i}W_{ij}>0$.
+Assume also $X$ to be strictly positive and $W$ nonnegative with $\min_{j}\max_{i}W_{ij}>0$.
 If $A$ is contractive with respect to $u$ and $C=\max_{i}d_{H}(X_{i},u)$ then
-$\max_{i}d_{H}(F(X)_{i},u)\leq\beta_{C}\max_{i}d_{H}(X_{i},u),$
-where $Y_{i}$ denotes the $i$-th column of $Y$, and $\beta_{C}\leq 1$.
+$\max_{i}d_{H}(F(X)_{:,i},u)\leq\beta_{C}\max_{i}d_{H}(X_{:,i},u),$
+where $Y_{:,i}$ denotes the $i$-th column of $Y$, and $\beta_{C}<1$.
 Nonlinear Perron-Frobenius theory extends some of the results about nonnegative matrices to particular classes of nonlinear functions.
@@
 In particular, it is strictly subhomogeneous if $\sigma(\lambda x)\ll_{\mathcal{K}}\lambda\sigma(x)$ for all $x\in\mathcal{K}$ and $\lambda>1$ and homogeneous if $\sigma(\lambda x)=\lambda\sigma(x)$ for all $x\in\mathcal{K}$ and $\lambda\geq 0$
-Subohomogeneity is a useful property and of practical utility.
-In fact, as discussed in, e.g.
-(Sittoni & Tudisco, 2024; Piotrowski et al., 2024), it is not difficult to verify that a broad range of activation functions commonly used in deep learning is subhomogeneous on $\mathcal{K}$.
+Subhomogeneity is a useful property and of practical utility.
+In fact, as discussed in  [38, 27], a broad range of activation functions commonly used in deep learning is subhomogeneous on $\mathcal{K}$.
 For this family of activation functions, Lemma 5.3 combined with arguments from nonlinear Perron–Frobenius theory yield the following main result
 Theorem 5.4.
-Consider a positive GNN
-$X^{(l+1)}=\sigma(A^{(l)}X^{(l)}W^{(l)})$ (16)
-with $X^{(0)}_{i}\in\mathcal{K}$ for any $i=1,\dots,d$ and $A^{(l)}$ and $W^{(l)}$ nonnegative for any $l$.
-Assume also that $u\in\mathcal{K}$ is the dominant eigenvector of all of the matrices $A^{(i)}$ and that it is also an eigenvector of $\sigma$.
-Then, if the matrices $\{A^{(i)}\}$ are uniformly contractive with respect to $u$ and $\min_{j}\max_{i}{W^{(l)}}_{ij}>0$ for any $l$, it holds
+Consider a GNN of the form $X^{(l+1)}=\sigma(A^{(l)}X^{(l)}W^{(l)})$ with $X^{(l)}_{:,i}\in\mathcal{K}$ for any $i=1,\dots,d$ and $A^{(l)}$ and $W^{(l)}$ nonnegative, for any $l$.
+Assume also that $u\in\mathcal{K}$ is the dominant eigenvector of all of the matrices $A^{(l)}$ and that it is also an eigenvector of $\sigma$.
+Then, if the matrices $\{A^{(l)}\}$ are uniformly contractive with respect to $u$ and $\min_{j}\max_{i}W^{(l)}_{ij}>0$ for any $l$, it holds
 $\lim_{l\rightarrow\infty}\mathrm{NumRank}(X^{(l)})=1.$
-We conclude with a formal investigation of the eigenpairs of activation functions that are entrywise subhomogeneous.
+We conclude with a formal investigation of the eigenpairs of activation functions that are entry-wise subhomogeneous.
 Let $\sigma=\otimes^{N}\psi$ with $\psi\in C(\mathbb{R},\mathbb{R})$ that is subhomogeneous on $\mathbb{R}_{+}$.
@@
 2) If $\sigma$ is strictly subhomogeneous, the only eigenvector of $\sigma$ in $\mathcal{K}$ is the constant vector.
-As a consequence of the above result, we find that the numerical rank of the features collapses both for GCNs with LeakyReLU activation function (that is homogeneous) as well as for GATs with any kind of subhomogeneous activation function.
-In fact, the constant vector is always an eigenvector of an entriwise nonlinear map.
-Table 2: The table lists the correlation coefficients of logarithm of metric values and the classification accuracies for GNNs of depth 2 - 24 layers and separately trained on Cora dataset.
-For Erank and NumRank, we subtract 1 so that both metrics converge to zero.
-The classification accuracy ratio are computed on 2-layered and 24-layered GNNs.
-Architecture $E_{\mathrm{Dir}}$ $E_{\mathrm{Proj}}$ MAD Erank NumRank Accuracy ratio
+As a consequence of the above result, we find that the numerical rank of the features collapses to one both for GCNs with LeakyReLU activation function (the setting of [4]) as well as for GATs with any kind of subhomogeneous activation function (the setting of [45]), as the constant vector is always an eigenvector of an entry-wise nonlinear map.
+For GCNs with any other subhomogeneous activation functions, the convergence of the numerical rank to 1 as stated in Theorem 5.1 may not hold as discussed in Appendix D.
+Dataset Model $E_{\mathrm{Dir}}$ $E_{\mathrm{Proj}}$ MAD Erank NumRank Accuracy ratio
 Standard Normalized Standard Normalized
-LReLU + GCN -0.7871 0.6644 -0.8106 -0.8309 -0.2460 0.9724 0.5885 0.2689
-Tanh + GCN 0.5243 0.9403 0.8610 0.9768 0.9734 0.9923 0.9784 0.1937
-LReLU + GCN + Bias -0.9300 0.7505 -0.9414 -0.3642 -0.2828 0.9847 0.7833 0.2108
-Tanh + GCN + Bias 0.4086 0.8950 0.7479 0.9026 0.8996 0.9926 0.9814 0.2133
-LReLU + GCN + LayerNorm -0.9132 0.8445 -0.9424 0.5591 0.9753 0.9736 0.9736 0.4882
-Tanh + GCN + LayerNorm -0.7119 0.2560 -0.2069 0.9716 0.8695 0.9576 0.9684 0.1885
-LReLU + GCN + PairNorm -0.8165 0.3731 -0.8106 0.3885 0.4850 0.5597 0.6244 0.8252
-Tanh + GCN + PairNorm -0.5963 0.0346 -0.5401 -0.3358 -0.0631 0.9838 0.9298 0.2902
-LReLU + GAT -0.9189 0.6703 -0.9469 -0.6054 0.8251 0.9722 0.7612 0.2487
-Tanh + GAT 0.8300 0.8501 0.8676 0.9066 0.8603 0.9600 0.9515 0.1767
-Average correlation -0.3911 0.6279 -0.2722 0.2568 0.5296 0.9349 0.8541
+Cora GCN -0.7871 0.6644 -0.8106 -0.8309 -0.2460 0.9724 0.5885 0.2693
+GAT -0.9189 0.6703 -0.9469 -0.6054 0.8251 0.9722 0.7612 0.2493
+Citeseer GCN -0.8442 0.4350 -0.8913 -0.8667 -0.7169 0.9700 0.6795 0.4380
+GAT -0.9576 0.0664 -0.9585 -0.9080 0.3722 0.9915 0.8047 0.4672
+Pubmed GCN -0.9068 0.7006 -0.8508 -0.1109 0.6205 0.9464 0.9268 0.5225
+GAT -0.8735 -0.3684 -0.8541 -0.4102 -0.3932 0.9270 0.9721 0.5564
+Squirrel GCN -0.7774 0.4171 -0.7602 -0.3258 -0.8247 0.6316 0.9582 0.8457
+GAT -0.6864 -0.5503 -0.7364 -0.7253 0.5002 0.8538 0.6840 0.7533
+Chameleon GCN -0.9223 0.1504 -0.9163 -0.8201 -0.8809 0.9387 0.9014 0.6195
+GAT -0.8721 0.1942 -0.9089 -0.8234 0.2803 0.9446 0.8799 0.6332
+Amazon Ratings GCN -0.9297 0.8809 -0.9079 -0.3423 0.9201 0.9301 0.8049 0.8562
+GAT -0.9388 0.5277 -0.9089 -0.1617 0.6545 0.9248 0.8764 0.8384
+OGB-Arxiv GCN 0.7738 0.9194 0.5740 -0.2738 0.2822 0.9682 0.9091 0.0939
+GAT -0.4097 0.9439 -0.7230 0.8985 0.8492 0.7740 0.9781 0.2310
+Average correlation -0.7179 0.4036 -0.7571 -0.4504 0.1601 0.9103 0.8374
+Table 1: Correlation between the classification accuracy and the logarithm of metric values on GNNs with LeakyReLU and of depths from 2 to 24 layers, separately trained on different homophilic (Cora, Citeseer, Pubmed), heterophilic (Squirrel, Chameleon, Amazon Ratings) and large-scale (OGB-Arxiv) datasets.
+For Erank and NumRank, we subtract 1 so that both metrics approach zero.
+The rightmost column reports the ratio of classification accuracy between GNNs with 2 and 24 layers.
+Some heterophilic datasets may be more resilient to the increasing network depth, in-line with observations from the literature, e.g.
+[13].
+Additional results on other datasets and activation functions are presented in Appendix E, and with additional network components in Appendix F.
 6 Experiments
-Empirical studies on the evolution of oversmoothing measures often use untrained, hundred-layer-deep GNNs (Rusch et al., 2022; Wang et al., 2022; Rusch et al., 2023b; Wu et al., 2023).
+Empirical studies on the evolution of oversmoothing measures often use untrained, hundred-layer-deep GNNs [35, 44, 34, 45].
 We emphasize that this is an overly simplified setting.
 In more realistic settings, as the ones considered in this section, a trained GNN may suffer from significant performance degradation after only few-layers, at which stage the convergence patterns of most of the metrics are difficult to observe.
-The experiments that we present in this section validate the robustness of the effective rank and numerical rank in quantifying oversmoothing in GNNs against the other metrics.
-In particular, we compare how different overmoothing metrics behave compared to the classification accuracy, varying the GNN architectures for node classification on real-world graph data.
+The experiments that we present in this section validate the robustness of the relaxed rank metrics in quantifying oversmoothing in GNNs against the other metrics.
+In particular, we compare how different oversmoothing metrics behave compared to the classification accuracy, varying the GNN architectures for node classification on real-world graph data.
 In our experiments, we consider the following metrics:
-  • The Dirichlet Energy $E_{\mathrm{Dir}}$ (Cai & Wang, 2020; Rusch et al., 2023a) and its variant $E_{\mathrm{Proj}}$ (Wu et al., 2023).
-Both are discussed in Section 3.1, see in particular (8) and (9).
-  • Normalized versions of the Dirichlet energy and its variant, namely $E_{\mathrm{Dir}}(X)/\|X\|_{F}^{2}$ and $E_{\mathrm{Proj}}(X)/\|X\|_{F}$.
+  • The Dirichlet Energy $E_{\mathrm{Dir}}$ [4, 33] and its variant $E_{\mathrm{Proj}}$ [45]. Both are discussed in Section 3.1, see in particular (1) and (2).
+  • Normalized versions of Dirichlet energy and its variant, $E_{\mathrm{Dir}}(X)/\|X\|_{F}^{2}$ and $E_{\mathrm{Proj}}(X)/\|X\|_{F}$.
 Indeed, from our previous discussion, a robust oversmoothing measure should be scale invariant with respect to the features.
-Metrics with global normalization like the ones we consider here have also been proposed in (Di Giovanni et al., 2023; Roth & Liebig, 2023; Maskey et al., 2023).
-  • The Mean Average Distance (MAD) (Chen et al., 2020)
-    $\text{MAD}(X)=\frac{1}{|\mathcal{E}|}\sum_{(i,j)\in\mathcal{E}}\left(1-\frac{X_{i}^{\top}X_{j}}{|X_{i}||X_{j}|}\right).$ (17)
+Metrics with global normalization like the ones we consider here have also been proposed in [8, 30, 20].
+  • The Mean Average Distance (MAD) [5]
+    $\textstyle\text{MAD}(X)=\frac{1}{|\mathcal{E}|}\sum_{(i,j)\in\mathcal{E}}\Big{(}1-\frac{X_{i,:}^{\top}X_{j,:}}{|X_{i,:}||X_{j,:}|}\Big{)}.$
     It measures the cosine similarity between the neighbouring nodes.
@@
 This further supports the use of the Numerical Rank as an improved measure of oversmoothing with respect to $E_{\mathrm{Proj}}$.
-Figure 1: Four examples of GCNs corresponding to the first fours rows in table 2.
+Figure 2: Four examples of the metric behaviours computed at the last hidden layer of separately trained GCNs of increasing depths.
 For Erank and Numrank, we measure Erank$(X)-r^{*}_{\mathrm{ER}}$ and $\mathrm{NumRank}(X)-r^{*}_{\mathrm{NR}}$ for some $r^{*}>1$.
 In these particular cases, $r^{*}_{\mathrm{ER}}<1.85$, $r^{*}_{\mathrm{NR}}<1.3$.
-Note that the effective rank and numerical rank of the input features $X^{(0)}$ is about 1084 and 13.6, respectively.
-In Table 2 and Figure 1, we train GNNs of a fixed hidden dimension equal to 32 on the Cora dataset.
-More results on Cora, as well as other datasets, are reported in Appendix B.
-We follow the standard setups of GCN and GAT as stated in equation (2) and (5), and use homogeneous LeakyReLU (LReLU) and subhomogenous Tanh as activation functions.
-For the GCN model, we also consider adding additional components, i.e.
-bias, LayerNorm (Ba et al., 2016) and PairNorm (Zhao & Akoglu, 2019).
+Note that the effective rank and numerical rank of the input features $X^{(0)}$ are about 1084 and 13.6, respectively.
+Additional results are attached in Appendix G.
+In Table 1 and Figure 2, we train GNNs of a fixed hidden dimension equal to 32 on homophilic, heterophilic and large-scale datasets in their default splits.
+We follow the standard setups of GCN and GAT as stated in Sections 2.1 and 2.2, and use homogeneous LeakyReLU (LReLU) as the activation function.
 For each configuration, GNNs of eight different depths ranging from 2 to 24 are trained.
@@
 The oversmoothing metrics are computed at the last hidden layer before the output layer.
-In Figure 1 and in Appendix B, we plot the behaviour of the different oversmoothing measures as well as the norm of the features and the accuracy of the trained GNNs when the depth of the network is increased.
+In Figure 2 and in Appendix G, we plot the behaviour of the different oversmoothing measures, the norm of the features, and the accuracy of the trained GNNs with increasing depth.
 These figures clearly show that the network suffers a significant drop in accuracy, which is not matched by any visible change in standard oversmoothing metrics.
-By contrast, the rank of the feature representations decreases drastically, following quite closely the behaviour of the network’s accuracy.
+By contrast, the rank of the feature representations decreases drastically, following closely the behaviour of the network’s accuracy.
 These findings are further sustained by the results shown in Table 1, where we compute the Pearson correlation coefficient between the logarithm of every measure and the classification accuracy of every GNN model.
 The use of a logarithmic transformation is based on the understanding that oversmoothing grows exponentially with the length of the network.
+Extensions of these results are provided in Appendices E and F.
+In Appendix D, we perform an asymptotic ablation study on very deep (300-layer) synthetic networks with randomly sampled, untrained weights.
+This study serves to validate our theoretical findings on the convergence of relaxed rank metrics and to demonstrate that such untrained settings offer little insight into the ability of existing metrics to quantify oversmoothing in realistic, trained networks.
 7 Conclusion
-In this paper, we have discussed the problem of quantifying oversmoothing in GNNs.
-After discussing the limitations of the leading oversmoothing measures, we have proposed the use of the rank of the features as a better measure of oversmoothing.
-The experiments that we provided validate the robustness of the effective rank against the classical measures.
-In addition, we have proved theoretically the decay of the rank of the features for linear and nonnegative GNNs.
+In this paper, we have discussed the problem of quantifying oversmoothing in message-passing GNNs.
+After simplifying the existing theoretical analysis using nonlinear activation eigenvectors and discussing the limitations of the leading oversmoothing measures, we propose the use of the rank of the features as a better measure of oversmoothing.
+We provide extensive experiments to validate the robustness of the effective rank against the classical measures.
+In addition, we have proved theoretically the decay of the rank of the features for message-passing GNNs.
+Acknowledgement
+KZ was supported by the EPSRC Centre for Doctoral Training in Mathematical Modelling, Analysis and Computation (MAC-MIGS) funded by the UK Engineering and Physical Sciences Research Council (grant EP/S023291/1), Heriot-Watt University and the University of Edinburgh.
+PD and FT are members of the Gruppo Nazionale Calcolo Scientifico - Istituto Nazionale di Alta Matematica (GNCS-INdAM).
+PD was supported by the MUR-PRO3 grant STANDS.
+DJH was supported by a Fellowship from the Leverhulme Trust.
+FT was partially funded by the PRIN-MUR project MOLE, code 2022ZK5ME7, and PRIN-PNRR project FIN4GEO within the European Union’s Next Generation EU framework, Mission 4, Component 2, CUP P2022BNB97.
 Appendix A Proofs of the main results
-A.1 proof of Lemma 3.2
-Let $\pi:=\mathrm{span}\{uv^{T}\,|\;v\in\mathbb{R}^{d}\}$ be the $1$-dimensional matrix subspace of the rank-$1$ matrices having columns aligned to $u$.
+A.1 Proof of Theorem 3.2
+We start proving that
+$\|(I-\mathcal{P})X^{(l+1)}\|_{F}\leq\|(I-\mathcal{P})A^{(l)}X^{(l)}W^{(l)}\|_{F},$ (6)
+where $\mathcal{P}=uu^{\top}/\|u\|^{2}$ is the projection matrix on the linear space spanned by $u$.
+To this end, let $\pi:=\mathrm{span}\{uv^{\top}\,|\;v\in\mathbb{R}^{d}\}$ be the $1$-dimensional matrix subspace of the rank-$1$ matrices having columns aligned to $u$.
 Then it is easy to note that given some matrix $X$, $(I-\mathcal{P})X$ provides the projection of the matrix $X$ on the subspace $\pi$, i.e.
 $(I-\mathcal{P})X=\mathrm{proj}_{\pi}(X).$ (7)
-Indeed $\langle(I-\mathcal{P})X,uv^{T}\rangle_{F}=\mathrm{Tr}(vu^{T}(I-uu^{T})X)=0$. In particular, since the projection realizes the minimal distance, we have that
-$\|X-\mathcal{P}X\|_{F}\leq\|X-uv^{T}\|_{F}\qquad\forall v\in\mathbb{R}^{d}.$ (19)
-Now observe that $\sigma(uu^{T}A^{(l-1)}X^{(l-1)}W^{(l-1)})=u\bar{v}^{T}$ for some $\bar{v}$.
-Indeed, writing $v^{T}=u^{T}A^{(l-1)}X^{(l-1)}W^{(l-1)}$, we have that the $i$-th column of $\sigma(uu^{T}A^{(l-1)}X^{(l-1)}W^{(l-1)})$ is equal to $\sigma(v_{i}u)=\bar{v}_{i}u$ for some $\bar{v}_{i}$, because $u$ is an eigenvector of $\sigma$.
+Indeed $\langle(I-\mathcal{P})X,uv^{\top}\rangle_{F}=\mathrm{Tr}(vu^{\top}(I-uu^{\top}/\|u\|^{2})X)=0$. In particular, since the projection realizes the minimal distance, we have that
+$\|X-\mathcal{P}X\|_{F}\leq\|X-uv^{\top}\|_{F}\qquad\forall v\in\mathbb{R}^{d}.$ (8)
+Now observe that $\sigma(uu^{\top}A^{(l-1)}X^{(l-1)}W^{(l-1)})=u\bar{v}^{\top}$ for some $\bar{v}$.
+Indeed, writing $v^{\top}=u^{\top}A^{(l-1)}X^{(l-1)}W^{(l-1)}$, we have that the $i$-th column of $\sigma(uu^{\top}A^{(l-1)}X^{(l-1)}W^{(l-1)})$ is equal to $\sigma(v_{i}u)=\bar{v}_{i}u$ for some $\bar{v}_{i}$, because $u$ is an eigenvector of $\sigma$.
 As a consequence we have
-$\|(I-\mathcal{P})X^{(l)}\|_{F}$ $\leq\|X^{(l)}-\sigma(uu^{T}A^{(l-1)}X^{(l-1)}W^{(l-1)})\|_{F}$ (20)
-$=\|\sigma(A^{(l-1)}X^{(l-1)}W^{(l-1)})-\sigma(uu^{T}A^{(l-1)}X^{(l-1)}W^{(l-1)})\|_{F}$
+$\|(I-\mathcal{P})X^{(l)}\|_{F}$ $\leq\|X^{(l)}-\sigma(uu^{\top}A^{(l-1)}X^{(l-1)}W^{(l-1)})\|_{F}$ (9)
+$=\|\sigma(A^{(l-1)}X^{(l-1)}W^{(l-1)})-\sigma(uu^{\top}A^{(l-1)}X^{(l-1)}W^{(l-1)})\|_{F}$
 $\leq\|(I-\mathcal{P})A^{(l-1)}X^{(l-1)}W^{(l-1)}\|_{F}$
-where we have used the $1$-Lipschitz property of $\sigma$.
+where we have used the $1$-Lipschitz property of $\sigma$. This concludes the proof of (6)
+To conclude the proof of the theorem observe that, in the decomposition
+$(I-\mathcal{P})A^{(l)}=(I-\mathcal{P})A^{(l)}\mathcal{P}+(I-\mathcal{P})A^{(l)}(I-\mathcal{P}),$
+the matrix $(I-\mathcal{P})A^{(l)}\mathcal{P}$ is zero because $A^{(l)}u=\lambda_{1}^{l}u$ for any $l$. Thus
+$(I-\mathcal{P})A^{(l)}=(I-\mathcal{P})A^{(l)}(I-\mathcal{P}),$
+and, from (6) and the inequality $\|AB\|_{F}\leq\|A\|_{2}\|B\|_{F}$, we have
+$\|(I-\mathcal{P})X^{(L)}\|_{F}\leq\Big{(}\Pi_{l=0}^{L-1}\|(I-\mathcal{P})A^{(l)}\|_{2}\|W^{(l)}\|_{2}\Big{)}\|X^{(0)}\|_{F}.$
+So the thesis follows from the hypothesis about the product $\Pi_{l=0}^{L-1}\|(I-\mathcal{P})A^{(l)}\|_{2}\|W^{(l)}\|_{2}$.
 A.2 Proof for Theorem 4.1
@@
 Similarly we can observe that
-$\|\mathcal{P}X^{(l)}\|_{F}$ $\geq\|u^{T}A^{l}X^{(0)}W^{(0)}\dots W^{(l-1)}\|_{F}=$ (23)
-$\|\big{(}\lambda_{1}^{l}v_{1}^{T}+u^{T}\widetilde{M}O\Big{(}{l\choose N}\lambda_{2}^{l-N}\Big{)}\widetilde{M^{\prime}}\big{)}X^{(0)}W^{(0)}\dots W^{(l-1)}\|_{F}\geq$
-$|\lambda_{1}|^{l}\Big{(}\|v_{1}^{T}X^{(0)}W^{(0)}\dots W^{(l-1)}\|_{F}-\Big{\|}u^{T}\widetilde{M}O\Big{(}{l\choose N}\Big{(}\frac{\lambda_{2}}{\lambda_{1}}\Big{)}^{l}\widetilde{M^{\prime}}X^{(0)}W^{(0)}\dots W^{(l-1)}\Big{\|}_{F}\Big{)}\geq$
-$|\lambda_{1}|^{l}\|v_{1}^{T}X^{(0)}W^{(0)}\dots W^{(l-1)}\|_{F}\Big{(}1-O\Big{(}{l\choose N}\Big{|}\frac{\lambda_{2}}{\lambda_{1}}\Big{|}^{l}\frac{\|X^{(0)}W^{(0)}\dots W^{(l-1)}\|_{F}}{\|v_{1}^{T}X^{(0)}W^{(0)}\dots W^{(l-1)}\|_{F}}\Big{)}.$
-Now observe that under the randomness hypothesis from (Furstenberg & Kifer, 1983) and more generally from the Oseledets ergodic multiplicative theorem, we have that for almost any $w\in\mathbb{R}^{d}$ the limit exists and is equal to the maximal Lyapunov exponent of the system, i.e.
-a constant $c(\nu)\geq 0$ depending only on the distribution $\nu$, $\lim_{l\rightarrow\infty}\frac{1}{l}\log\|w^{T}W^{(0)}\dots W^{(l-1)}\|=c(\nu)$.
+$\|\mathcal{P}X^{(l)}\|_{F}\geq\|u^{\top}A^{l}X^{(0)}W^{(0)}\dots W^{(l-1)}\|_{F}=$ (12)
+$=\|\big{(}\lambda_{1}^{l}v_{1}^{\top}+u^{\top}\widetilde{M}O\Big{(}{l\choose N}\lambda_{2}^{l-N}\Big{)}\widetilde{M^{\prime}}\big{)}X^{(0)}W^{(0)}\dots W^{(l-1)}\|_{F}\geq$
+$\geq|\lambda_{1}|^{l}\Big{(}\|v_{1}^{\top}X^{(0)}W^{(0)}\dots W^{(l-1)}\|_{F}-\Big{\|}u^{\top}\widetilde{M}O\Big{(}{l\choose N}\Big{(}\frac{\lambda_{2}}{\lambda_{1}}\Big{)}^{l}\widetilde{M^{\prime}}X^{(0)}W^{(0)}\dots W^{(l-1)}\Big{\|}_{F}\Big{)}\geq$
+$\geq|\lambda_{1}|^{l}\|v_{1}^{\top}X^{(0)}W^{(0)}\dots W^{(l-1)}\|_{F}\Big{(}1-O\Big{(}{l\choose N}\Big{|}\frac{\lambda_{2}}{\lambda_{1}}\Big{|}^{l}\frac{\|X^{(0)}W^{(0)}\dots W^{(l-1)}\|_{F}}{\|v_{1}^{\top}X^{(0)}W^{(0)}\dots W^{(l-1)}\|_{F}}\Big{)}$
+Now observe that under the randomness hypothesis from [11] and more generally from the Oseledets ergodic multiplicative theorem, we have that for almost any the limit $w\in\mathbb{R}^{d}$ $\lim_{l\rightarrow\infty}\frac{1}{l}\log\|w^{\top}W^{(0)}\dots W^{(l-1)}\|=c(\nu)$ exists and is equal to the maximal Lyapunov exponent of the system.
 In particular for any $w$ and $\epsilon>0$ there exists $l_{w,\epsilon}$ sufficiently large such that for any $l>l_{w,\epsilon}$
-$c(\nu)-\epsilon\leq\frac{1}{l}\log\|w^{T}W^{(0)}\dots W^{(l-1)}\|<c(\nu)+\epsilon.$ (24)
+$c(\nu)-\epsilon\leq\frac{1}{l}\log\|w^{\top}W^{(0)}\dots W^{(l-1)}\|<c(\nu)+\epsilon$ (13)
 i.e.
-$e^{l(c(\nu)-\epsilon)}\leq\|w^{T}W^{(0)}\dots W^{(l-1)}\|<e^{l(c(\nu)+\epsilon)}\quad\forall l\geq l_{w,\epsilon}$ (25)
-Now take as vector $w$ first the rows of $X^{(0)}$ and then the vector $v_{1}^{T}X^{(0)}$, than almost surely for any $\epsilon$ there exists $l_{\epsilon}$ such that for any $l>l_{\epsilon}$
-$e^{l(c(\nu)-\epsilon)}\leq\|w^{T}W^{(0)}\dots W^{(l-1)}\|<e^{l(c(\nu)+\epsilon)},$ (26)
-holding for any $l\geq l_{\epsilon}$ and any $w\in\{v_{1}\}\cup\{X_{0}^{T}e_{i}\}_{i=1}^{N}$.
-Next recall that $\|X^{(0)}W^{(0)}\dots W^{(l-1)}\|_{F}=\sqrt{\sum_{i}\|e_{i}^{T}X^{(0)}W^{(0)}\|^{2}}$, meaning that almost surely, for $l\geq l_{\epsilon}$:
+$e^{l(c(\nu)-\epsilon)}\leq\|w^{\top}W^{(0)}\dots W^{(l-1)}\|<e^{l(c(\nu)+\epsilon)}\quad\forall l\geq l_{w,\epsilon}$ (14)
+Now take as vector $w$ first the rows of $X^{(0)}$ and then the vector $v_{1}^{\top}X^{(0)}$, than almost surely for any $\epsilon$ there exists $l_{\epsilon}$ such that for any $l>l_{\epsilon}$
+$e^{l(c(\nu)-\epsilon)}\leq\|w^{\top}W^{(0)}\dots W^{(l-1)}\|<e^{l(c(\nu)+\epsilon)},$ (15)
+holding for any $l\geq l_{\epsilon}$ and any $w\in\{v_{1}\}\cup\{X_{0}^{\top}e_{i}\}_{i=1}^{N}$.
+Next recall that $\|X^{(0)}W^{(0)}\dots W^{(l-1)}\|_{F}=\sqrt{\sum_{i}\|e_{i}^{\top}X^{(0)}W^{(0)}\|^{2}}$, meaning that almost surely, for $l\geq l_{\epsilon}$:
 $Ne^{l(c(\nu)-\epsilon)}\leq\|X^{(0)}W^{(0)}\dots W^{(l-1)}\|_{F}\leq Ne^{l(c(\nu)+\epsilon)}.$ (16)
 In particular for any $\epsilon$, there exists $l$ sufficiently large such that
-($\Big{(}{l\choose N}\Big{|}\frac{\lambda_{2}}{\lambda_{1}}\Big{|}^{l}\frac{\|X^{(0)}W^{(0)}\dots W^{(l-1)}\|_{F}}{\|v_{1}^{T}X^{(0)}W^{(0)}\dots W^{(l-1)}\|_{F}}\Big{)}\leq\Big{(}{l\choose N}\Big{|}\frac{\lambda_{2}}{\lambda_{1}}\Big{|}^{l}e^{2l\epsilon}\Big{)}$ (28)
+($\left({l\choose N}\Big{|}\frac{\lambda_{2}}{\lambda_{1}}\Big{|}^{l}\frac{\|X^{(0)}W^{(0)}\dots W^{(l-1)}\|_{F}}{\|v_{1}^{\top}X^{(0)}W^{(0)}\dots W^{(l-1)}\|_{F}}\right)\leq\left({l\choose N}\Big{|}\frac{\lambda_{2}}{\lambda_{1}}\Big{|}^{l}e^{2l\epsilon}\right)$ (17)
 and thus, since $|\lambda_{2}|<|\lambda_{1}|$ and we can choose $\epsilon$ arbitrarily small, almost surely it has limit equal to zero. In particular we can write
-$\lim_{l}\frac{\|(I-\mathcal{P}X^{(l)})\|_{F}}{\|\mathcal{P}X^{(l)}\|_{F}}\sim\lim_{l}\frac{{l\choose N}|\lambda_{2}|^{l-N}\|X^{(0)}W^{(0)}\dots W^{(l)}\|_{F}}{|\lambda_{1}|^{l}\|v_{1}^{T}X^{(0)}W^{(0)}\dots W^{(l)}\|_{F}}=0$ (29)
+$\lim_{l}\frac{\|(I-\mathcal{P}X^{(l)})\|_{F}}{\|\mathcal{P}X^{(l)}\|_{F}}\sim\lim_{l}\frac{{l\choose N}|\lambda_{2}|^{l-N}\|X^{(0)}W^{(0)}\dots W^{(l)}\|_{F}}{|\lambda_{1}|^{l}\|v_{1}^{\top}X^{(0)}W^{(0)}\dots W^{(l)}\|_{F}}=0$ (18)
 where we have used the same argument as before to state that the limit is zero.
@@
 A.4 proof of Lemma 5.3
-By the contraction properties of the matrix $A$ we know that if $\max_{i}d_{H}\big{(}(AX)_{i},u\big{)}\leq C$, then
-$d_{H}\Big{(}(AX)_{i},u\Big{)}=d_{H}\Big{(}(AX)_{i},\lambda_{1}(A)u\Big{)}=d_{H}\Big{(}(AX)_{i},Au\Big{)}\leq\beta d_{H}\big{(}X_{i},u\big{)}\qquad\forall i.$ (35)
+By the contraction properties of the matrix $A$ we know that if $\max_{i}d_{H}\big{(}(AX)_{:,i},u\big{)}\leq C$, then
+$d_{H}\Big{(}(AX)_{:,i},u\Big{)}=d_{H}\Big{(}(AX)_{:,i},\lambda_{1}(A)u\Big{)}=d_{H}\Big{(}(AX)_{:,i},Au\Big{)}\leq\beta d_{H}\big{(}X_{:,i},u\big{)}\qquad\forall i.$ (24)
 for some $\beta<1$, where we have used $\lambda_{1}(A)>0$ and the scaling invariant property of the Hilbert distance.
-Then note that, for any $i$, we can write $F(X)_{i}$ as follows
-$\big{(}F(X)\big{)}_{j}=\sum_{j}W_{ij}(AX)_{j}\,.$ (36)
+Then note that, for any $i$, we can write $F(X)_{:,j}$ as follows
+$\big{(}F(X)\big{)}_{:,j}=\sum_{j}W_{ij}(AX)_{:,j}\,.$ (25)
 Thus we CLAIM that given $x_{1},x_{2},y\in\mathcal{K}$ then
@@
 Observe that if the claim holds we have conlcuded the proof, indeed by induction it can trivially be extended from $2$ to $d$ points yielding
-$d_{H}\Big{(}F(X)_{j},u\Big{)}\leq\max_{i}d_{H}\Big{(}W_{ij}(AX)_{i},u\Big{)}\leq\max_{j}d_{H}\Big{(}(AX)_{j},u\Big{)}\leq\beta\max_{j}d_{H}\Big{(}X_{j},u\Big{)},$ (38)
+$d_{H}\Big{(}F(X)_{:,j},u\Big{)}\leq\max_{i}d_{H}\Big{(}W_{ij}(AX)_{:,i},u\Big{)}\leq\max_{j}d_{H}\Big{(}(AX)_{:,j},u\Big{)}\leq\beta\max_{j}d_{H}\Big{(}X_{:,j},u\Big{)},$ (27)
 where we have used the scale-invariance property of the Hilbert distance and the fact that $\max_{i}W_{ij}>0$ for all $j$.
-It remains to prove the claim. To this end, exploiting the expression of the Hilbert distance we write
+We only miss to prove the claim. To this end, exploiting the expression of the Hilbert distance we write
 $d_{H}(x_{1}+x_{2},y)$ $=\log\bigg{(}\sup_{j}\sup_{i}\frac{(x_{1})_{i}+(x_{2})_{i}}{(y)_{i}}\frac{(y)_{j}}{(x_{1})_{j}+(x_{2})_{j}}\bigg{)}$ (28)
 $\leq\log\bigg{(}\sup_{i}\sup_{j}\max_{x_{1},x_{2}}\bigg{\{}\frac{(x_{1})_{i}}{(x_{1})_{j}},\frac{(x_{2})_{i}}{(x_{2})_{j}}\bigg{\}}\frac{(y)_{j}}{(y)_{i}}\bigg{)}$
-$=\max_{x_{1},x_{2}}\big{\{}d_{H}(x_{1},y),d_{H}(x_{2},y)\big{\}}$
+$=\max_{x_{1},x_{2}}\big{\{}d_{H}(x_{1},y),d_{H}(x_{2},y)\big{\}}.$
 concluding the proof.
@@
 Now using the claim above and Lemma 5.3, we can easily conclude that
-$\lim_{l\rightarrow\infty}\max_{i}d_{H}(X_{i}^{(l)},u)=0.$ (44)
-Indeed by the uniform contractivity of $\{A^{(l)}\}$, if $C=\max_{i}d_{H}(X^{(0)}_{i},u)$, then there exists some $\beta_{C}<1$ such that
-$\max_{i}d_{H}(X^{(l)}_{i},u)\leq\beta_{C}^{l}\max_{i}d_{H}(X^{(0)}_{i},u).$ (45)
+$\lim_{l\rightarrow\infty}\max_{i}d_{H}(X_{:,i}^{(l)},u)=0.$ (33)
+Indeed by the uniform contractivity of $\{A^{(l)}\}$, if $C=\max_{i}d_{H}(X^{(0)}_{:,i},u)$, then there exists some $\beta_{C}<1$ such that
+$\max_{i}d_{H}(X^{(l)}_{:,i},u)\leq\beta_{C}^{l}\max_{i}d_{H}(X^{(0)}_{:,i},u).$ (34)
 Finally to conclude we prove that, as a consequence of (33),
@@
 $1\leq\mathrm{NumRank}(X^{(l)})\leq 1+\frac{\|(I-\mathcal{P})X^{(l)}\|_{F}^{2}}{\|X^{(l)}\|_{2}^{2}}\leq 1+\frac{\|(I-\mathcal{P})X^{(l)}\|_{F}^{2}}{\|\mathcal{P}X^{(l)}\|_{F}^{2}},$ (36)
-we have that (46) finally yields the thesis. To prove (46), we recall from Lemma 2.5.1 in (Lemmens & Nussbaum, 2012) that for any $w$ such that $u^{T}w=c$
+we have that (35) finally yields the thesis. To prove (35), we recall from Lemma 2.5.1 in [18] that for any $w$ such that $u^{\top}w=c$
 $\|w-\mathcal{P}w\|_{u}\leq\|\mathcal{P}w\|_{u}(e^{d_{T}(w,\mathcal{P}w)}-1),$ (37)
 where $d_{T}(x,y)=\log(\max\{M(x/y),m(x/y)^{-1}\})$ and where since the dual cone of +$\mathbb{R}^{n}_{+}$ is +$\mathbb{R}^{n}_{+}$ itself, we are considering the norm induced by $u$ on the cone, i.e.
-$\|x\|_{u}=u^{T}x$ for any $x$ in the cone.
-In practice the norm induced by $u$ $\|\cdot\|_{u}$ can be defined by the Minkowki functional of the set $\Omega=\text{ConvexHull}\{\{\Omega_{1}\}\cup\{-\Omega_{1}\}\}$ where $\Omega_{1}=\{x\in\mathcal{K}\,\;u^{T}x\leq 1\}$
-Then since $\|\mathcal{P}w\|_{u}=\|w\|_{u}=u^{T}w$, we have that $M(w/\mathcal{P}w)\geq 1$ and $m(w/\mathcal{P}w)\leq 1$. Thus $d_{T}(w,\mathcal{P}w)\leq d_{H}(w,\mathcal{P}w)$, yielding
+$\|x\|_{u}=u^{\top}x$ for any $x$ in the cone.
+In practice the norm induce by $u$ $\|\cdot\|_{u}$ can be defines by the Minkowki functional of the set $\Omega=\text{ConvexHull}\{\{\Omega_{1}\}\cup\{-\Omega_{1}\}\}$ where $\Omega_{1}=\{x\in\mathcal{K}\,\;u^{\top}x\leq 1\}$
+Then since $\|\mathcal{P}w\|_{u}=\|w\|_{u}=u^{\top}w$, we have that $M(w/\mathcal{P}w)\geq 1$ and $m(w/\mathcal{P}w)\leq 1$. Thus $d_{T}(w,\mathcal{P}w)\leq d_{H}(w,\mathcal{P}w)$, yielding
 $\|w-\mathcal{P}w\|_{u}\leq\|\mathcal{P}w\|_{u}(e^{d_{H}(w,\mathcal{P}w)}-1).$ (38)
@@
 Now recall that the squared frobenius norm of a matrix is the sum of the squared 2-norms of the its columns, so we can apply the last inequality to the matrix $X^{(l)}$ columnwise obtaining:
-$\|(I-\mathcal{P})X^{(l)}\|_{F}^{2}\leq C\|\mathcal{P}X^{(l)}\|_{F}^{2}\big{(}e^{\max_{i}\{d_{H}(X^{(l)}_{i},\mathcal{P}X^{(l)}_{i})\}}-1\big{)}^{2}.$ (51)
-The proof is concluded observing that by the scale invariance property of the Hilbert distance $d_{H}(X^{(l)}_{i},\mathcal{P}X^{(l)}_{i})=d_{H}(X^{(l)}_{i},u)$, yielding:
-$1\leq\big{(}\mathrm{NumRank}(X^{(l)})\big{)}\leq 1+\frac{\|(I-\mathcal{P})X^{(l)}\|_{F}^{2}}{\|\mathcal{P}X^{(l)}\|_{F}^{2}}\leq 1+C\big{(}e^{\max_{i}\{d_{H}(X^{(l)}_{i},u)\}}-1\big{)}^{2}.$ (52)
+$\|(I-\mathcal{P})X^{(l)}\|_{F}^{2}\leq C\|\mathcal{P}X^{(l)}\|_{F}^{2}\big{(}e^{\max_{i}\{d_{H}(X^{(l)}_{:,i},\mathcal{P}X^{(l)}_{:,i})\}}-1\big{)}^{2}.$ (40)
+the proof is concluded using (33) and observing that by the scale invariance property of the Hilbert distance $d_{H}(X^{(l)}_{:,i},\mathcal{P}X^{(l)}_{:,i})=d_{H}(X^{(l)}_{:,i},u)$, yielding:
+$1\leq\big{(}\mathrm{NumRank}(X^{(l)})\big{)}\leq 1+\frac{\|(I-\mathcal{P})X^{(l)}\|_{F}^{2}}{\|\mathcal{P}X^{(l)}\|_{F}^{2}}\leq 1+C\big{(}e^{\max_{i}\{d_{H}(X^{(l)}_{:,i},u)\}}-1\big{)}^{2}$ (41)
+and concluding the proof.
 A.6 proof of Proposition 5.5
@@
 In particular any constant vector $u$ in $\mathbb{R}_{+}^{N}$ is easily proved to be an eigenvector of $\sigma$ relative to the eigenvalue $\lambda=\|\sigma(u)\|_{1}/\|u\|_{1}=\psi(u_{i})/u_{i}$ where $u_{i}$ is any entry of $u$.
-Appendix B Additional Empirical Results on Real Dataset
-LReLU + GCN $r^{*}_{\mathrm{ER}}=1.450100,\quad r^{*}_{\mathrm{NR}}=1.207918$ Tanh + GCN $r^{*}_{\mathrm{ER}}=1.690520,\quad r^{*}_{\mathrm{NR}}=1.007135$
+Appendix B Limitations
+Although a small effective rank or numerical rank indicates oversmoothing and can be subsequently linked to the underperformance of GNNs, a large effective rank or numerical rank does not necessarily correspond to a good network performance.
+Prior study has suggested some degree of smoothing can be beneficial [16], and as an extreme example, features sampled from a uniform distribution and with randomly assigned labels would almost surely have a large effective rank, but they cannot be classified accurately due to the lack of any underlying pattern.
+We note that this limitation is not specific to the two relaxed rank measures, as the same argument is directly applicable to all other oversmoothing metrics.
+Consequently, as shown in Table 4, when additional components are used to (partially) alleviate oversmoothing, particularly when residual connections are used, the accuracy ratio may remain large over the layers, and all oversmoothing metrics correlate poorly with the accuracy of GNNs.
+This in-turn suggests these oversmoothing metrics become less informative as the oversmoothing problem is mitigated or alleviated.
+Appendix C Computational Complexity Analysis
+Let $N$, $D$, $E$ the number of nodes, features and edges of a graph $\mathcal{G}$, the computational cost of the Dirichlet energy is $O(E\times D)$, while the cost of the Projection Energy is $O(N\times D)$.
+In contrast, most of the computational cost for the numerical rank and the effective rank is given by the computation of the spectral radius of $X$, and of all the singular values of $X$ for the effective rank, respectively.
+Standard results show that it is always possible to compute the full singular value decomposition (SVD), and subsequently the spectral radius, in $O(N\times D\times\min\{N,D\})$.
+However, in typical cases, the latter cost can be drastically reduced using different strategies.
+Firstly, in the case of numerical rank, the computational cost of the 2-norm is generally much smaller than the cost of the full SVD.
+Indeed, using Lanczos or power methods to compute it, the cost scales as $O(N\times D)$, and the methods converge typically very fast.
+In addition, both the effective rank and the numerical rank can be efficiently controlled by computing only the $k$-largest singular values of $X$.
+In particular, a truncated SVD containing the $k$-largest singular values can be computed in $O(N\times D\times k)$ using either deterministic algorithms or randomized SVD methods.
+In general, the computational cost of metrics to quantify oversmoothing is marginal compared to the cost of training.
+In production, measuring the emergence of oversmoothing is done by training the model and checking the performance on the validation and test sets as the number of layers changes.
+The cost of additionally computing effective or numerical ranks is marginal.
+Moreover, we note that computing the metrics on a subsection of a graph can be sufficiently informative, and as a consequence, most oversmoothing metrics studied in this paper can be computed in less than $10~{}ms$ for a graph (or a subsection of it) with less than 2000 nodes.
+Appendix D Additional Experimental Results with Synthetic Weights
+In this section, we conduct an asymptotic ablation study using randomly sampled (synthetic) untrained weights. The aim of these experiments is twofold:
+  • to demonstrate that similar untrained asymptotic experiments are inherently unrealistic, despite being extensively used in the literature [44, 35, 34, 45, 29, 42], as they fail to reliably capture oversmoothing in shallower GNNs where the performance degradation occurs.
+  • to examine the convergence properties of different oversmoothing metrics with weight size control and to empirically validate Theorems 5.1 and 5.4.
+We construct a 10-node Barabasi-Albert graph with each node having 32 features.
+The weights are either an identity matrix or randomly sampled at each layer from a uniform distribution $\mathcal{U}(0,s)$, where $s$ depends on the settings: small weights ($s=0.05$) lead to an exponentially decaying $\|X^{(l)}\|_{F}$, and large weights ($s=0.1$) lead to an exploding $\|X^{(l)}\|_{F}$ for uncapped activation functions.
+For identity weights, $\|X^{(l)}\|_{F}$ is roughly constant (LReLU) or slowly decaying (Tanh).
+The feature initialization $X^{(0)}$ is sampled from $\mathcal{U}(0,1)$, and is iterated over 300 layers.
+In this asymptotic synthetic setting, as presented in Table 2, the normalized $E_{\mathrm{Dir}}$ and $E_{\mathrm{Proj}}$ exhibit decay patterns similar to those of the effective rank and numerical rank, suggesting these metrics are equally sensitive to asymptotic rank collapse.
+However, this behaviour stands in stark contrast to results on trained networks presented in Tables 1 and 3 and Appendix G, where the normalized $E_{\mathrm{Dir}}$ and $E_{\mathrm{Proj}}$ often fail to detect oversmoothing.
+Moreover, these asymptotic results validate Theorems 5.1 and 5.4, showing that the numerical rank converges to one for GCN + LReLU and GAT + any subhomogeneous activation functions.
+Without making any additional assumption on the normalization of the adjacency matrix, the effective rank and numerical rank do not generally decay to one when subhomogeneous activation functions, e.g.
+Tanh, are used in GCNs.
+Architecture $E_{\mathrm{Dir}}$ $E_{\mathrm{Proj}}$ MAD Erank NumRank
+standard normalized standard normalized
+GCN+LReLU+identity weights
+GCN+Tanh+identity weights
+GAT+LReLU+identity weights
+GAT+Tanh+identity weights
+GCN+LReLU+small weights
+GCN+Tanh+small weights
+GAT+LReLU+small weights
+GAT+Tanh+small weights
+GCN+LReLU+large weights
+GCN+Tanh+large weights
+GAT+LReLU+large weights
+GAT+Tanh+large weights
+Table 2: Additional results on very deep (300 layers) synthetic networks with randomly sampled weights.
+For Erank and NumRank, we subtract 1 so that both metrics converge to zero.
+✓ indicates a decay of the corresponding metric to zero, ✗ indicates otherwise.
+Note that GAT has similar asymptotic behaviour to GCN with adjacency normalization $D^{-1}\widetilde{A}$.
+Appendix E Additional Experimental Results on Activation Functions and Different Datasets
+We extend Table 1 to subhomogenous Tanh activation function and a few additional datasets. The experimental setting is consistent with that of Section 6.
+Dataset Architecture $E_{\mathrm{Dir}}$ $E_{\mathrm{Proj}}$ MAD Erank NumRank Accuracy ratio
+Standard Normalized Standard Normalized
+Cora GCN+LReLU -0.7871 0.6644 -0.8106 -0.8309 -0.2460 0.9724 0.5885 0.2693
+GCN+Tanh 0.5243 0.9403 0.8610 0.9768 0.9734 0.9923 0.9784 0.1937
+GAT+LReLU -0.9189 0.6703 -0.9469 -0.6054 0.8251 0.9722 0.7612 0.2493
+GAT+Tanh 0.8300 0.8501 0.8676 0.9066 0.8603 0.9600 0.9515 0.1900
+Citeseer GCN+LReLU -0.8442 0.4350 -0.8913 -0.8667 -0.7169 0.9700 0.6795 0.4380
+GCN+Tanh 0.3420 0.8957 0.4631 0.9045 0.9605 0.9906 0.9457 0.3509
+GAT+LReLU -0.9576 0.0664 -0.9585 -0.9080 0.3722 0.9915 0.8047 0.4672
+GAT+Tanh 0.9234 0.8949 0.8276 0.8997 0.9176 0.9287 0.9024 0.3045
+Pubmed GCN+LReLU -0.9068 0.7006 -0.8508 -0.1109 0.6205 0.9464 0.9268 0.5225
+GCN+Tanh -0.2330 0.2657 0.2029 0.9137 0.8745 0.9328 0.9940 0.3883
+GAT+LReLU -0.8735 -0.3684 -0.8541 -0.4102 -0.3932 0.9270 0.9721 0.5564
+GAT+Tanh 0.2977 0.8411 0.7160 0.9331 0.8546 0.9303 0.8551 0.4464
+Squirrel GCN+LReLU -0.7774 0.4171 -0.7602 -0.3258 -0.8247 0.6316 0.9582 0.8457
+GCN+Tanh 0.6026 0.7736 0.1689 0.9377 0.7727 0.9680 0.9837 0.8152
+GAT+LReLU -0.6864 -0.5503 -0.7364 -0.7253 0.5002 0.8538 0.6840 0.7533
+GAT+Tanh -0.3606 -0.6557 -0.0363 0.8714 -0.7033 0.8911 0.8861 0.9103
+Chameleon GCN+LReLU -0.9223 0.1504 -0.9163 -0.8201 -0.8809 0.9387 0.9014 0.6195
+GCN+Tanh 0.2742 0.8796 -0.2541 0.8492 0.9272 0.9869 0.9841 0.7093
+GAT+LReLU -0.8721 0.1942 -0.9089 -0.8234 0.2803 0.9446 0.8799 0.6332
+GAT+Tanh 0.4090 0.5699 0.0743 0.8230 0.6006 0.9613 0.9143 0.7052
+Amazon Ratings GCN+LReLU -0.9297 0.8809 -0.9079 -0.3423 0.9201 0.9301 0.8049 0.8562
+GCN+Tanh -0.6960 -0.6289 -0.6910 0.8576 0.9423 0.9871 0.9327 0.8574
+GAT+LReLU -0.9388 0.5277 -0.9089 -0.1617 0.6545 0.9248 0.8764 0.8384
+GAT+Tanh 0.8354 0.8507 -0.6595 0.9245 0.9418 0.8954 0.8883 0.8382
+Roman Empire GCN+LReLU -0.5635 0.7703 -0.6772 0.2575 0.6420 0.5833 0.5368 0.3891
+GCN+Tanh 0.8018 0.9225 -0.6808 0.8359 0.8124 0.8570 0.8090 0.4067
+GAT+LReLU -0.9174 0.5390 -0.9407 0.1868 0.7582 0.7221 0.8722 0.3705
+GAT+Tanh 0.5767 0.5844 -0.4716 0.8819 0.7263 0.7332 0.8589 0.3652
+OGB-Arxiv GCN+LReLU 0.7738 0.9194 0.5740 -0.2738 0.2822 0.9682 0.9091 0.0957
+GCN+Tanh -0.6409 0.7041 -0.6808 0.9494 0.9487 0.9900 0.9876 0.1204
+GAT+LReLU -0.4097 0.9439 -0.7230 0.8985 0.8492 0.7740 0.9781 0.2310
+GAT+Tanh 0.7834 0.7896 -0.9277 0.9393 0.8222 0.7671 0.7861 0.2376
+Average correlation -0.1956 0.5137 -0.3886 0.2669 0.4960 0.9007 0.8684
+Table 3: Additional correlation coefficient results on homophilic (Cora, Citeseer, Pubmed), heterophilic (Squirrel, Chameleon, Amazon Ratings, Roman Empire) and large-scale (OGB-Arxiv) dataset.
+Appendix F Additional Experimental Results with Different Network Components
+Prior literature indicates that when adding additional components, such as bias [33] or residual terms [36], the Dirichlet energy does not decay.
+Therefore, we compare the correlation coefficients in Table 4 between existing oversmoothing metrics when additional components are added, such as bias, LayerNorm, BatchNorm, PairNorm [46], DropEdge [28, 15] and residual connections [36].
+All experiments follow the setup described in Section 6. In addition, DropEdge has a probability of 0.5 in removing each edge at each layer. The residual connection is implemented as follows
+$X^{(l+1)}=\sigma(AX^{(l)}W_{1}^{(l)})+X^{(0)}W_{2}^{(l)}$
+Table 4 demonstrates that both the effective rank and numerical rank achieve a higher average correlation with the classification accuracy than alternative metrics.
+This finding confirms their superior consistency in detecting oversmoothing across a variety of architectural variants.
+Furthermore, we note that when oversmoothing is effectively alleviated, e.g.
+when residual connection is used, all metrics have a poor correlation with the classification accuracy.
+This generic limitation is discussed in Appendix B.
+Architecture $E_{\mathrm{Dir}}$ $E_{\mathrm{Proj}}$ MAD Erank NumRank Accuracy ratio
+Standard Normalized Standard Normalized
+GCN+LReLU+Bias -0.9300 0.7505 -0.9414 -0.3642 -0.2828 0.9847 0.7833 0.2110
+GCN+Tanh+Bias 0.4086 0.8950 0.7479 0.9026 0.8996 0.9926 0.9814 0.2133
+GCN+LReLU+LayerNorm -0.9132 0.8445 -0.9424 0.5591 0.9753 0.9736 0.9736 0.4882
+GCN+Tanh+LayerNorm -0.7119 0.2560 -0.2069 0.9716 0.8695 0.9576 0.9684 0.1886
+GCN+LReLU+BatchNorm 0.8789 0.7300 0.8872 0.7175 0.5094 0.6005 0.6577 0.8761
+GCN+Tanh+BatchNorm -0.6033 0.2883 -0.6587 0.7717 0.4984 0.7377 0.7038 0.8157
+GCN+LReLU+PairNorm -0.8165 0.3731 -0.8106 0.3885 0.4850 0.5597 0.6244 0.8556
+GCN+Tanh+PairNorm -0.5963 0.0346 -0.5401 -0.3358 -0.0631 0.9838 0.9298 0.3123
+GCN+LReLU+DropEdge -0.7497 0.7185 -0.7939 -0.7974 -0.4619 0.9720 0.5782 0.2515
+GCN+Tanh+DropEdge 0.2319 0.8388 0.5872 0.8887 0.8763 0.9934 0.9558 0.2096
+GCN+LReLU+Residual -0.0857 -0.3072 -0.0559 -0.1611 -0.1656 -0.2466 -0.1296 1.0046
+GCN+Tanh+Residual -0.6751 -0.7406 -0.3031 -0.4019 -0.6043 -0.5425 -0.4897 1.0122
+Average correlation -0.3801 0.3901 -0.2525 0.2616 0.2946 0.6638 0.6280
+Table 4: Additional correlation coefficient results on GCNs with different network components.
+Appendix G Metric Behaviour Examples on Cora, Citeseer and Pubmed
+We extend Figure 2 to additional datasets and network components. The experimental setting is consistent with that of Section 6.
+GCN + LReLU $r^{*}_{\mathrm{ER}}=1.450100,\quad r^{*}_{\mathrm{NR}}=1.207918$ GCN + Tanh $r^{*}_{\mathrm{ER}}=1.690520,\quad r^{*}_{\mathrm{NR}}=1.007135$
          [Uncaptioned image]          [Uncaptioned image]
-LReLU + GCN + Bias $r^{*}_{\mathrm{ER}}=1.454981,\quad r^{*}_{\mathrm{NR}}=1.020319$ Tanh + GCN + Bias $r^{*}_{\mathrm{ER}}=1.858602,\quad r^{*}_{\mathrm{NR}}=1.021634$
+GCN + LReLU + Bias $r^{*}_{\mathrm{ER}}=1.454981,\quad r^{*}_{\mathrm{NR}}=1.020319$ GCN + Tanh + Bias $r^{*}_{\mathrm{ER}}=1.858602,\quad r^{*}_{\mathrm{NR}}=1.021634$
          [Uncaptioned image]          [Uncaptioned image]
-LReLU + GCN + LayerNorm $r^{*}_{\mathrm{ER}}=1.017424,\quad r^{*}_{\mathrm{NR}}=1.000101$ Tanh + GCN + LayerNorm $r^{*}_{\mathrm{ER}}=1.528688,\quad r^{*}_{\mathrm{NR}}=1.003682$
+GCN + LReLU + LayerNorm $r^{*}_{\mathrm{ER}}=1.017424,\quad r^{*}_{\mathrm{NR}}=1.000101$ GCN + Tanh + LayerNorm $r^{*}_{\mathrm{ER}}=1.528688,\quad r^{*}_{\mathrm{NR}}=1.003682$
          [Uncaptioned image]          [Uncaptioned image]
-LReLU + GAT $r^{*}_{\mathrm{ER}}=1.295623,\quad r^{*}_{\mathrm{NR}}=1.079250$ Tanh + GAT $r^{*}_{\mathrm{ER}}=1.002191,\quad r^{*}_{\mathrm{NR}}=1.000006$
+GAT + LReLU $r^{*}_{\mathrm{ER}}=1.295623,\quad r^{*}_{\mathrm{NR}}=1.079250$ GAT + Tanh $r^{*}_{\mathrm{ER}}=1.002191,\quad r^{*}_{\mathrm{NR}}=1.000006$
          [Uncaptioned image]          [Uncaptioned image]
 Table 5: The table showcases the behaviour of different metrics and the classification accuracies for 8 GNNs separately trained on Cora Dataset. This table is an extension of figure 2
-LReLU + GCN $r^{*}_{\mathrm{ER}}=1.750563,\quad r^{*}_{\mathrm{NR}}=1.142120$ Tanh + GCN $r^{*}_{\mathrm{ER}}=1.436012,\quad r^{*}_{\mathrm{NR}}=1.003035$
+GCN + LReLU $r^{*}_{\mathrm{ER}}=1.750563,\quad r^{*}_{\mathrm{NR}}=1.142120$ GCN + Tanh $r^{*}_{\mathrm{ER}}=1.436012,\quad r^{*}_{\mathrm{NR}}=1.003035$
          [Uncaptioned image]          [Uncaptioned image]
-LReLU + GCN + Bias $r^{*}_{\mathrm{ER}}=1.526297,\quad r^{*}_{\mathrm{NR}}=1.045451$ Tanh + GCN + Bias $r^{*}_{\mathrm{ER}}=1.752850,\quad r^{*}_{\mathrm{NR}}=1.007417$
+GCN + LReLU + Bias $r^{*}_{\mathrm{ER}}=1.526297,\quad r^{*}_{\mathrm{NR}}=1.045451$ GCN + Tanh + Bias $r^{*}_{\mathrm{ER}}=1.752850,\quad r^{*}_{\mathrm{NR}}=1.007417$
          [Uncaptioned image]          [Uncaptioned image]
-LReLU + GCN + LayerNorm $r^{*}_{\mathrm{ER}}=1.006405,\quad r^{*}_{\mathrm{NR}}=1.000009$ Tanh + GCN + LayerNorm $r^{*}_{\mathrm{ER}}=1.473417,\quad r^{*}_{\mathrm{NR}}=1.002555$
+GCN + LReLU + LayerNorm $r^{*}_{\mathrm{ER}}=1.006405,\quad r^{*}_{\mathrm{NR}}=1.000009$ GCN + Tanh + LayerNorm $r^{*}_{\mathrm{ER}}=1.473417,\quad r^{*}_{\mathrm{NR}}=1.002555$
          [Uncaptioned image]          [Uncaptioned image]
-LReLU + GAT $r^{*}_{\mathrm{ER}}=1.541711,\quad r^{*}_{\mathrm{NR}}=1.148837$ Tanh + GAT $r^{*}_{\mathrm{ER}}=1.003797,\quad r^{*}_{\mathrm{NR}}=1.000005$
+GAT + LReLU $r^{*}_{\mathrm{ER}}=1.541711,\quad r^{*}_{\mathrm{NR}}=1.148837$ GAT + Tanh $r^{*}_{\mathrm{ER}}=1.003797,\quad r^{*}_{\mathrm{NR}}=1.000005$
          [Uncaptioned image]          [Uncaptioned image]
 Table 6: The table showcases the behaviour of different metrics and the classification accuracies for 8 GNNs separately trained on Citeseer Dataset.
-LReLU + GCN $r^{*}_{\mathrm{ER}}=1.420913,\quad r^{*}_{\mathrm{NR}}=1.085279$ Tanh + GCN $r^{*}_{\mathrm{ER}}=1.838982,\quad r^{*}_{\mathrm{NR}}=1.014851$
+GCN + LReLU $r^{*}_{\mathrm{ER}}=1.420913,\quad r^{*}_{\mathrm{NR}}=1.085279$ GCN + Tanh $r^{*}_{\mathrm{ER}}=1.838982,\quad r^{*}_{\mathrm{NR}}=1.014851$
          [Uncaptioned image]          [Uncaptioned image]
-LReLU + GCN + Bias $r^{*}_{\mathrm{ER}}=1.450100,\quad r^{*}_{\mathrm{NR}}=1.207918$ Tanh + GCN + Bias $r^{*}_{\mathrm{ER}}=1.443576,\quad r^{*}_{\mathrm{NR}}=1.018361$
+GCN + LReLU + Bias $r^{*}_{\mathrm{ER}}=1.450100,\quad r^{*}_{\mathrm{NR}}=1.207918$ GCN + Tanh + Bias $r^{*}_{\mathrm{ER}}=1.443576,\quad r^{*}_{\mathrm{NR}}=1.018361$
          [Uncaptioned image]          [Uncaptioned image]
-LReLU + GCN + LayerNorm $r^{*}_{\mathrm{ER}}=1.002088,\quad r^{*}_{\mathrm{NR}}=1.000074$ Tanh + GCN + LayerNorm $r^{*}_{\mathrm{ER}}=1.668289,\quad r^{*}_{\mathrm{NR}}=1.009275$
+GCN + LReLU + LayerNorm $r^{*}_{\mathrm{ER}}=1.002088,\quad r^{*}_{\mathrm{NR}}=1.000074$ GCN + Tanh + LayerNorm $r^{*}_{\mathrm{ER}}=1.668289,\quad r^{*}_{\mathrm{NR}}=1.009275$
          [Uncaptioned image]          [Uncaptioned image]
-LReLU + GAT $r^{*}_{\mathrm{ER}}=1.577163\quad r^{*}_{\mathrm{NR}}=1.119213$ Tanh + GAT $r^{*}_{\mathrm{ER}}=1.016096,\quad r^{*}_{\mathrm{NR}}=1.000019$
+GAT + LReLU $r^{*}_{\mathrm{ER}}=1.577163\quad r^{*}_{\mathrm{NR}}=1.119213$ GAT + Tanh $r^{*}_{\mathrm{ER}}=1.016096,\quad r^{*}_{\mathrm{NR}}=1.000019$
          [Uncaptioned image]          [Uncaptioned image]

### v4

@@
-Rethinking Oversmoothing in Graph Neural Networks: A Rank-Based Perspective
+Are We Measuring Oversmoothing in Graph Neural Networks Correctly?
 Abstract
@@
 Traditionally, oversmoothing has been quantified using metrics that measure the similarity of neighbouring node features, such as the Dirichlet energy.
-While these metrics are related to oversmoothing, we argue they have critical limitations and fail to reliably capture oversmoothing in realistic scenarios.
-For instance, they provide meaningful insights only for very deep networks and under somewhat strict conditions on the norm of network weights and feature representations.
+We argue that these metrics have critical limitations and fail to reliably capture oversmoothing in realistic scenarios.
+For instance, they provide meaningful insights only for very deep networks, while typical GNNs show a performance drop already with as few as 10 layers.
 As an alternative, we propose measuring oversmoothing by examining the numerical or effective rank of the feature representations.
-We provide theoretical support for this approach, demonstrating that the numerical rank of feature representations converges to one for a broad family of nonlinear activation functions under the assumption of nonnegative trained weights.
-To the best of our knowledge, this is the first result that proves the occurrence of oversmoothing in the nonlinear setting without assumptions on the boundedness of the weight matrices.
-Along with the theoretical findings, we provide extensive numerical evaluation across diverse graph architectures.
-Our results show that rank-based metrics consistently capture oversmoothing, whereas energy-based metrics often fail.
-Notably, we reveal that a significant drop in the rank aligns closely with performance degradation, even in scenarios where energy metrics remain unchanged.
+We provide extensive numerical evaluation across diverse graph architectures and datasets to show that rank-based metrics consistently capture oversmoothing, whereas energy-based metrics often fail.
+Notably, we reveal that drops in the rank align closely with performance degradation, even in scenarios where energy metrics remain unchanged.
+Along with the experimental evaluation, we provide theoretical support for this approach, clarifying why Dirichlet-like measures may fail to capture performance drop and proving that the numerical rank of feature representations collapses to one for a broad family of GNN architectures.
 1 Introduction
-Graph neural networks (GNNs) have emerged as a powerful framework for learning representations from graph-structured data, with applications spanning knowledge retrieval and reasoning [25, 39], personalised recommendation systems [7, 26], social network analysis [10], and 3D mesh classification [37].
+Graph neural networks (GNNs) have emerged as a powerful framework for learning representations from graph-structured data, with applications spanning knowledge retrieval and reasoning (Tian et al., 2022; Peng et al., 2023), personalised recommendation systems Peng et al.
+(2022); Damianou et al.
+(2024), social network analysis Fan et al.
+(2019), and 3D mesh classification Shi and Rajkumar (2020).
 Central to most GNN architectures is the message-passing paradigm, where node features are iteratively aggregated from their neighbours and transformed using learned functions, such as multi-layer perceptrons or graph-attention mechanisms.
 However, the performance of message-passing-based GNNs is known to deteriorate after only a few layers, essentially placing a limit on their depth.
-This issue, often linked to the increasingly similar learned features as GNNs deepen, is known as oversmoothing [19, 23].
-Over the years, oversmoothing in GNNs, as well as methods to alleviate it, have been studied based on the decay of some node feature similarity metrics, such as the Dirichlet energy and its variants [24, 4, 2, 21, 8, 45, 30].
+This issue, often linked to the increasingly similar learned features as GNNs deepen, is known as oversmoothing Li et al.
+(2018); Nt and Maehara (2019); Wu et al.
+(2022); Rusch et al.
+(2023a); Zhao et al.
+(2024); Arnaiz-Rodriguez and Errica (2025).
+In recent years, oversmoothing in GNNs, as well as methods to alleviate it, have been studied based on the decay of some node feature similarity metrics, such as the Dirichlet energy and its variants Oono and Suzuki (2019); Cai and Wang (2020); Bodnar et al.
+(2022); Nguyen et al.
+(2022); Di Giovanni et al.
+(2023); Wu et al.
+(2023); Roth and Liebig (2023).
 At a high level, most of these metrics directly measure the norm of the absolute deviation from the dominant eigenspace of the message-passing matrix.
@@
 However, when nonlinear activation functions or biases are used, the dominant eigenspace may change, causing these oversmoothing metrics to fail and give false negative signals about the oversmoothing state of the learned features.
-Therefore, these metrics are often considered as sufficient but not necessary evidences for oversmoothing [33].
-Despite this, there is a considerable body of literature using these unreliable metrics as their evidences for non-occurrence of oversmoothing in GNNs [47, 6, 35, 44, 20, 22, 34, 9, 36, 43, 29, 42].
-As we show in Section 6, the performance degradation of GNNs trained on real datasets often happens well before any noticeable decay in these oversmoothing metrics can be observed.
-Most empirical studies in the literature that observe the decay of the Dirichlet-energy-like metrics are conducted over the layers of the same very deep untrained (with randomly sampled weights) or effectively untrained [^1] Deep networks (with, say, over 100 layers) that are trained but whose loss and accuracy remain far from acceptable.
-GNNs [44, 35, 34, 45, 29, 42], where the decay of the metrics is likely driven by the small weight initializations reducing all features to zero.
-Instead, we observe that when GNNs of different depths are separately trained, these metrics do not correlate well with their performance degradation.
-Furthermore, we note that these metrics can only indicate oversmoothing when their values converge exactly to zero, corresponding to either an exact alignment to the dominant eigenspace or to the feature representation matrix collapsing to the all-zero matrix.
+While these metrics are often considered as sufficient but not necessary evidence for oversmoothing Rusch et al.
+(2023a), there is a considerable body of literature using these unreliable metrics as their evidence for non-occurrence of oversmoothing in GNNs Zhou et al.
+(2021); Chen et al.
+(2022); Rusch et al.
+(2022); Wang et al.
+(2022); Maskey et al.
+(2023); Nguyen et al.
+(2023); Rusch et al.
+(2023b); Epping et al.
+(2024); Scholkemper et al.
+(2024); Wang and Cho (2024); Roth (2024); Wang et al.
+(2025); Arnaiz-Rodriguez and Errica (2025).
+However, as we show in Section˜6, the performance degradation of GNNs trained on real datasets often happens well before any noticeable decay in these oversmoothing metrics can be observed.
+The vast majority of empirical studies in the literature that observe the decay of Dirichlet-like energy metrics are conducted over the layers of very deep but untrained (with randomly sampled weights) or effectively untrained [^1] Deep networks (with, say, over 100 layers) that are trained but whose loss and accuracy remain far from acceptable.
+GNNs Wang et al.
+(2022); Rusch et al.
+(2022; 2023b); Wu et al.
+(2023); Roth (2024); Wang et al.
+(2025), where the decay of the metrics is only driven by the small weight initialization.
+Instead, we show that when GNNs of different depths are trained with proper weight initialization, these metrics do not correlate with the model’s performance degradation.
+Furthermore, we argue that these metrics can only indicate oversmoothing when their values converge exactly to zero, corresponding to either an exact alignment to the dominant eigenspace or to the feature representation matrix collapsing to the all-zero matrix.
 This double implication presents an issue: in realistic settings with a large but not excessively large number of layers, we may observe the decay of the oversmoothing metric by, say, two orders of magnitude while still being far from zero.
@@
 As an alternative to address these shortcomings, we advocate for the use of a continuous approximation of the rank of the network’s feature representations to measure oversmoothing.
-Our experimental evaluation across various GNN architectures trained for node classification demonstrates that continuous rank relaxations, such as the numerical rank and the effective rank, correlate strongly with performance degradation in independently trained GNNs—even in settings where popular energy-like metrics show little to no correlation.
+Collapse of the rank of the feature representations was already considered as a cause of oversmoothing, e.g.
+in Guo et al.
+(2023b); Roth and Liebig (2023), however rank rank-based measures were never explicitly studied, i.e.
+they were never compared with other oversmoothing measures and their capacity in measuring oversmoothing was never quantified, theoretically and/or numerically.
+Our work fills this gap.
+Indeed, our experimental evaluation across various GNN architectures trained for node classification demonstrates that continuous rank relaxations, such as the numerical rank and the effective rank, correlate strongly with performance degradation in independently trained GNNs—even in settings where popular energy-like metrics show little to no correlation.
 Overall, the main contributions of this paper are as follows:
   • We review popular oversmoothing metrics in the current literature and simplify their theoretical analysis from a novel perspective of nonlinear activation eigenvectors.
-  • We propose the rank as a better metric for quantifying oversmoothing, and thereby re-defining oversmoothing in GNNs as the convergence towards a low-rank matrix rather than to a matrix of exactly rank one.
-  • We prove the convergence of the numerical rank towards one for linear GNNs and nonlinear GNNs where the eigenvector of the message-passing matrix is also the eigenvector of the nonlinear activation function under the assumption of non-negative weights.
-To the best of our knowledge, this is the first theoretical result proving that oversmoothing may occur independently of the weights’ magnitude with nonlinear activation functions.
-  • We provide extensive numerical evidence that continuous rank relaxation functions better capture oversmoothing than commonly used Dirichlet-like metrics.
+  • We notice that the rank can be a better metric for quantifying oversmoothing, and thereby we redefine oversmoothing in GNNs as the convergence towards a low-rank matrix rather than to a matrix of exactly rank one.
+  • We provide extensive numerical evidence that continuous rank relaxation functions provide a much more compelling measure of oversmoothing than commonly used Dirichlet-like metrics
+Additionally, we investigate theoretically the causes of decay of the numerical rank.
+In particular, we show that both the aggregation matrices and the nonlinear activation functions can contribute to the decay.
+Our theoretical study is restricted to linear GNNs and nonlinear and non-negative GNNs where the eigenvector of the message-passing matrix is also the eigenvector of the nonlinear activation function.
+For these models, we prove that the numerical rank of the features converges exactly to one.
+Such results provide theoretical support to our empirical evidence that oversmoothing may occur independently of the weights’ magnitude and align with our perspective on oversmoothing from the point of view of nonlinear activation functions.
 2 Background
@@
 The rows of the feature matrix $X\in\mathbb{R}^{N\times d}$ are the concatenation of the $d$-dimensional feature vectors of all nodes in the graph.
-At each layer $l$, the node feature update of Graph Convolutional Network (GCN) [17] follows $X^{(l+1)}=\sigma(AX^{(l)}W^{(l)})$ where $\sigma$ is a nonlinear activation function, applied component-wise, and $W^{(l)}$ is a trainable weight matrix.
+At each layer $l$, the node feature update of Graph Convolutional Network (GCN) Kipf and Welling (2016) follows $X^{(l+1)}=\sigma(AX^{(l)}W^{(l)})$ where $\sigma$ is a nonlinear activation function, applied component-wise, and $W^{(l)}$ is a trainable weight matrix.
 2.2 Graph Attention Network
-Graph Attention Networks (GATs) [41, 3] perform graph convolution via a layer-dependent message-passing matrix $A^{(l)}$ learned through an attention mechanism $A^{(l)}_{ij}=\text{softmax}_{j}(\sigma_{a}(p_{1}^{(l)\top}W^{(l)\top}X_{i,:}+p_{2}^{(l)\top}W^{(l)\top}X_{j,:}))$ where $p_{i}^{(l)}$ are learnable parameter vectors, $X_{i,:},X_{j,:}$ denote the feature of the $i$-th and $j$th nodes respectively, the activation $\sigma_{a}$ is typically chosen to be LeakyReLU, and $\text{softmax}_{j}$ corresponds to the row-wise normalization $\text{softmax}_{j}(A_{ij})=\exp(A_{ij})/\sum_{j^{\prime}}\exp(A_{ij^{\prime}})$.
+Graph Attention Networks (GATs) Veličković et al.
+(2017); Brody et al.
+(2021) perform graph convolution via a layer-dependent message-passing matrix $A^{(l)}$ learned through an attention mechanism $A^{(l)}_{ij}=\text{softmax}_{j}(\sigma_{a}(p_{1}^{(l)\top}W^{(l)\top}X_{i,:}+p_{2}^{(l)\top}W^{(l)\top}X_{j,:}))$ where $p_{i}^{(l)}$ are learnable parameter vectors, $X_{i,:},X_{j,:}$ denote the feature of the $i$-th and $j$th nodes respectively, the activation $\sigma_{a}$ is typically chosen to be LeakyReLU, and $\text{softmax}_{j}$ corresponds to the row-wise normalization $\text{softmax}_{j}(A_{ij})=\exp(A_{ij})/\sum_{j^{\prime}}\exp(A_{ij^{\prime}})$.
 The corresponding feature update is $X^{(l+1)}=\sigma(A^{(l)}X^{(l)}W^{(l)})\,.$
@@
 Indeed, if $\widetilde{A}$ is the adjacency matrix of a fully connected graph, $A$ will have spectral radius equal to $1$ with multiplicity $1$, and $A^{l}$ will converge toward the eigenspace spanned by the dominant eigenvector.
-Precisely, $A^{l}\to uv^{\top}$ as $l\rightarrow\infty$, where $Au=u$ and $A^{\top}v=v$ [40].
-As a consequence, if the product of the weight matrices $W^{(0)}\cdots W^{(l)}$ converges in the limit $l\rightarrow\infty$, then the features degenerate to a matrix having rank at most one, where all the features are aligned with the dominant eigenvector $u$.
+Precisely, $A^{l}\to uv^{\top}$ as $l\rightarrow\infty$, where $Au=u$ and $A^{\top}v=v$, see e.g.
+tudiscoComplexPowerNonnegative2015.
+As a consequence, if the product of the weight matrices $W^{(0)}\cdots W^{(l)}$ does not diverge in the limit $l\rightarrow\infty$, then the features degenerate to a matrix having rank at most one, where all the features are aligned with the dominant eigenvector $u$.
 Mathematically, if we assume $u$ to be such that $\|u\|=1$, this alignment can be expressed by stating that the difference between the features and their projection onto $u$, given by $\|X^{(l)}-uu^{\top}X^{(l)}\|$, converges to zero.
@@
 In particular, in most cases, it is measured exactly by the alignment of the features with the dominant eigenvector of the matrix $A$.
-The most prominent metric that has been used to quantify oversmoothing is the Dirichlet energy, which measures the norm of the difference between the degree-normalized neighbouring node features [4, 33]
+The most prominent metric that has been used to quantify oversmoothing is the Dirichlet energy, which measures the norm of the difference between the degree-normalized neighbouring node features Cai and Wang (2020); Rusch et al.
+(2023a)
 $E_{\mathrm{Dir}}(X)=\sum_{(i,j)\in\mathcal{E}}\left\|\frac{X_{i,:}}{u_{i}}-\frac{X_{j,:}}{u_{j}}\right\|^{2}_{2},$ (1)
@@
 This intuition suggests that a similar behaviour may occur for “smooth-enough” nonlinearities.
-In particular, in the case of a GCN, the dominant eigenvector $u$ is defined by $u_{i}=\sqrt{1+d_{i}}$ and [4] have proved that, using LeakyReLU activation functions, it holds $E_{\mathrm{Dir}}(X^{(l+1)})\leq s_{l}\bar{\lambda}E_{\mathrm{Dir}}(X^{(l)})$, where $s_{l}=\|W^{(l)}\|_{2}$ is the largest singular value of the weight matrix $W^{(l)}$, and $\bar{\lambda}=(1-\min_{i}\lambda_{i})^{2}$, where $\lambda_{i}\in(0,2]$ varies among the nonzero eigenvalues of the normalized graph Laplacian $\widetilde{\Delta}=I-A=I-\widetilde{D}^{-\frac{1}{2}}\widetilde{A}\widetilde{D}^{-\frac{1}{2}}$.
+In particular, in the case of a GCN, the dominant eigenvector $u$ is defined by $u_{i}=\sqrt{1+d_{i}}$ and Cai and Wang (2020) have proved that, using LeakyReLU activation functions, it holds $E_{\mathrm{Dir}}(X^{(l+1)})\leq s_{l}\bar{\lambda}E_{\mathrm{Dir}}(X^{(l)})$, where $s_{l}=\|W^{(l)}\|_{2}$ is the largest singular value of the weight matrix $W^{(l)}$, and $\bar{\lambda}=(1-\min_{i}\lambda_{i})^{2}$, where $\lambda_{i}\in(0,2]$ varies among the nonzero eigenvalues of the normalized graph Laplacian $\widetilde{\Delta}=I-A=I-\widetilde{D}^{-\frac{1}{2}}\widetilde{A}\widetilde{D}^{-\frac{1}{2}}$.
 Similarly, in the case of GATs, the matrices $A_{i}$ are all row stochastic, meaning that $u_{i}=1$ for all $i$.
-In this case, it has been proved that whenever the product of the entry-wise absolute value of the weights is bounded, that is $\|\Pi_{k=1}^{\infty}|W^{(k)}|\|<\infty$, then the following variant of the Dirichlet energy decays to zero [45]
+In this case, it has been proved that whenever the product of the entry-wise absolute value of the weights is bounded, that is $\|\Pi_{k=1}^{\infty}|W^{(k)}|\|<\infty$, then the following variant of the Dirichlet energy decays to zero Wu et al.
+(2023)
 $E_{\mathrm{Proj}}(X)=\|X-\mathcal{P}X\|^{2}_{F}$ (2)
 where $\mathcal{P}=uu^{\top}$ is the projection matrix on the space spanned by the dominant eigenvector $u$ of the matrices $A^{(l)}$.
-In particular, this metric can be used only if the dominant eigenvector $u$ is the same for all $l$; this is, for example, the case with row stochastic matrices or when $A^{(l)}=A$ for all $l$.
+Note that both these metrics, $E_{\mathrm{Proj}}$ and $E_{\mathrm{Dir}}$, can be used only if the dominant eigenvector of $A^{(l)}$ is the same for all $l$; this is, for example, the case with row stochastic matrices or when $A^{(l)}=A$ for all $l$.
+Moreover, both these metrics essentially measure the deviation of the feature representations from the dominant eigenspace of the aggregation matrices $A^{(l)}$.
+So we expect them to perform very similarly in capturing oversmoothing.
+In particular, it is not difficult to show that they are equivalent metrics from a mathematical point of view, i.e.
+there exist constants $C_{1},C_{2}>0$ such that $C_{1}E_{\mathrm{Dir}}(X)\leq E_{\mathrm{Proj}}(X)\leq C_{2}E_{\mathrm{Dir}}(X)$, see Lemma˜A.1.
 3.2 A Unifying Perspective Based on the Eigenvectors of Nonlinear Activations
-In the following discussion, we present a unified and more general perspective of the necessary conditions to have oversmoothing in the sense of the classical metrics, based on the eigenvectors of a nonlinear activation function.
-In the interest of space, longer proofs for this and the subsequent sections are moved to Appendix A.
+We present here a unifying and more general perspective of the necessary conditions to have oversmoothing in the sense that $E_{\mathrm{Proj}}$ and $E_{\mathrm{Dir}}$ decay to zero, based on the concept of eigenvectors for a nonlinear activation function.
+In the interest of space, longer proofs for this and the subsequent sections are moved to Appendix˜A.
 Definition 3.1.
 We say that a vector $u\in\mathbb{R}^{N}\setminus\{0\}$ is an eigenvector of the (nonlinear) activation function $\sigma:\mathbb{R}^{N}\to\mathbb{R}^{n}$ if for any $t\in\mathbb{R}\setminus\{0\}$, there exists $\mu_{t}\in\mathbb{R}$ such that $\sigma(tu)=\mu_{t}u$.
-With this definition, we can now provide a unifying characterization of message-passing operators $A^{(l)}$ and activation functions $\sigma$ that guarantee the convergence of the Dirichlet-like energy metric $E_{\mathrm{Proj}}$ to zero for the feature representation sequence defined by $X^{(l+1)}=\sigma(A^{(l)}X^{(l)}W^{(l)})$.
+With this definition, we can now provide a unifying characterization of message-passing operators $A^{(l)}$ and activation functions $\sigma$ that guarantee the convergence of the Dirichlet-like energy metrics $E_{\mathrm{Proj}}$ and $E_{\mathrm{Dir}}$ to zero for the feature representation sequence defined by $X^{(l+1)}=\sigma(A^{(l)}X^{(l)}W^{(l)})$.
 Specifically, Theorem 3.2 shows that this holds provided all matrices $A^{(l)}$ share a common dominant eigenvector $u$, which is also an eigenvector of $\sigma$.
-This assumption recurs throughout our theoretical analysis, aligning with existing results in the literature.
-For example, (a) the proof by [4] applies to GCNs with LeakyReLU, where the dominant eigenvector of $A$ is nonnegative by the Perron-Frobenius theorem and, therefore, an eigenvector of LeakyReLU; and (b) the proof by [45] holds for stochastic message-passing matrices $A^{(l)}$, which inherently share a common dominant eigenvector with constant entries, that is an eigenvector for any pointwise nonlinear activation function.
 Theorem 3.2.
-Let $X^{(l+1)}=\sigma(A^{(l)}X^{(l)}W^{(l)})$, $l=1,\dots,L$, be a GNN such that $u$ is the dominant eigenvector of $A_{l}$ for any $l$ and is also an eigenvector of the activation function $\sigma$.
+Let $X^{(l+1)}=\sigma(A^{(l)}X^{(l)}W^{(l)})$, $l=1,\dots,L$, be a GNN such that $u$ is the dominant eigenvector of $A^{(l)}$ for any $l$ and also an eigenvector of the activation $\sigma$.
 If $\sigma$ is $1$-Lipschitz, namely $\|\sigma(x)-\sigma(y)\|\leq\|x-y\|$ for any $x,y$, and $\lim_{L\to\infty}\prod_{l=0}^{L}\|(I-\mathcal{P})A^{(l)}\|_{2}\|W^{(l)}\|_{2}=0$, then
 $E_{\mathrm{Proj}}(X^{(L)})\to 0\quad\text{ as }\quad L\to\infty.$
-Note, in particular, that in the case of GCNs, the matrix $A^{(l)}=A$ is symmetric, and thus $\|I-\mathcal{P}A^{(l)}\|_{2}=\lambda_{2}$.
-Therefore, when $\sigma=\text{LeakyReLU}$, we obtain the result by [4] as convergence to zero is guaranteed if $\|W^{(l)}\|_{2}\leq\lambda_{2}$.
-Recall that, as discussed above, the choice $\sigma=\text{LeakyReLU}$ satisfies our eigenvector assumption since $u\geq 0$ by the Perron-Frobenius theorem, and thus LeakyReLU$(tu)=\alpha tu$ with $\alpha$ depending only on the sign of $t$.
-Similarly, in the case of GATs, the matrices $A^{(l)}$ are stochastic for all $l$, implying that $u=\mathbbm{1}$ is the constant vector with $(u)_{i}=1$ for all $i$.
-If $\sigma=\otimes\psi$ is a nonlinear activation function acting entry-wise through $\psi$, then $\sigma(t\mathbbm{1})=\psi(t)\mathbbm{1}$.
-Therefore, Theorem 3.2 implies that if the weights are sufficiently small, the features align independently of the activation function used.
-This is consistent with the results in [45].
-However, we note that the bounds on the weights required by Theorem 3.2 and those in [45] on the weights $W^{(l)}$ are not identical, and it is unclear which of the two is more significant.
-Nonetheless, in both cases, having bounded weights along with any $1$-Lipschitz pointwise activation function is a sufficient condition for observing oversmoothing via $E_{\mathrm{Proj}}$ in GATs.
-In addition to offering a different and unifying theoretical perspective on the results in [4, 45], we highlight the simplicity of our eigenvector-based proof, which we hope provides added clarity on this phenomenon.
+[…]
 4 Energy-like Metrics: What Can Go Wrong
@@
 While the former is a symptom of oversmoothing, the latter does not necessarily imply oversmoothing.
-Moreover, this convergence property requires the weights to be bounded.
-However, in practical cases, performance degradation is observed even in relatively shallow networks, far from being infinitely deep, and with arbitrarily large (or small) weight magnitudes.
+Moreover, this convergence property requires the weights to be strongly bounded.
+However, in most practical cases, performance degradation is observed even in relatively shallow networks, far from being infinitely deep, and with weight magnitudes arbitrarily larger than what is prescribed by Cai and Wang (2020); Wu et al.
+(2023) or Theorem˜3.2.
 This aligns with our intuition and what occurs in the linear case.
@@
 Theorem 4.1.
-[…]
+Let $X^{(l+1)}=AX^{(l)}W^{(l)}$ be a linear GCN.
+Let $\lambda_{1},\lambda_{2}$ be the largest and second-largest eigenvalues (in modulus) of $A$, respectively.
+Assume the weights $\{W^{(l)}\}_{l=1}^{\infty}$ are randomly sampled from i.i.d.
+random variables with distribution $\nu$ such that $\int\log^{+}(\|W\|)d\nu+\int\log^{+}(\|W^{-1}\|)d\nu<\infty$, with $\log^{+}(t)=\max\{\log(t),\,0\}$.
+If $|\lambda_{2}/\lambda_{1}|<1$, then almost surely
 $\lim_{l\to\infty}\frac{\|(I-\mathcal{P})X^{(l)}\|_{F}}{\|\mathcal{P}X^{(l)}\|_{F}}=0$
 with a linear rate of convergence $|\lambda_{2}/\lambda_{1}|$.
-In particular, the theorem above implies that for a large spectral gap $|\lambda_{2}/\lambda_{1}|\ll 1$, $X^{(l)}$ is predominantly of rank one, namely $X^{(l)}=\lambda_{1}^{l}\Big{(}uv^{\top}+R(l)\Big{)}$ for some $v$, with $R(l)\sim O(|\lambda_{2}/\lambda_{1}|)^{l}$ and thus converging to 00 as $l\to\infty$.
+In particular, the theorem above implies that $X^{(l)}=\lambda_{1}^{l}\big(uv^{\top}+R(l)\big)$ for some $v$, with $R(l)\sim O(|\lambda_{2}/\lambda_{1}|)^{l}$ and thus, when the spectral gap is large $|\lambda_{2}/\lambda_{1}|\ll 1$, $X^{(l)}$ is predominantly of rank one, even for moderate values of $l$.
 This results in weakly expressive feature representations, independently of the magnitude of the feature weights.
-This phenomenon can be effectively captured by measuring the rank of $X^{(l)}$, whereas Dirichlet-like energy measures may fail to detect it, this is e.g.
-the case when $\lambda_{1}>1$ having $X^{(l)}\rightarrow\infty$.
-Another important limitation of these metrics is their dependence on a specific, known dominant eigenspace, which must either be explicitly known or computed in advance.
+This phenomenon can be effectively captured by measuring the rank of $X^{(l)}$, whereas Dirichlet-like energy measures may fail to detect it, as it would, for example, be the case when $\lambda_{1}>1$, having $X^{(l)}\rightarrow\infty$.
+Another important limitation of Dirichlet-like metrics is their dependence on a specific dominant eigenspace, which must either be explicitly known or computed in advance.
 Consequently, their applicability is strongly tied to the specific architecture of the network.
 In particular, the dominant eigenvector $u$ of $A^{(l)}$ must be known and remain the same for all $l$.
-This requirement excludes their use in most cases where $A^{(l)}$ varies with $l$, such as when $A^{(l)}$ is the standard weighted adjacency matrix of the graph.
+This requirement excludes their use in cases where $A^{(l)}$ varies with $l$.
 5 The Rank as a Measure of Oversmoothing
-Inspired by the behaviour observed in the linear case, we argue that measuring the rank of feature representations provides a more effective way to quantify oversmoothing, in alignment with recent work on oversmoothing [14].
+Inspired by the behaviour observed in the linear case, we argue that measuring the rank of feature representations provides a more effective way to quantify oversmoothing, in alignment with recent work on oversmoothing Guo et al.
+(2023b); Roth and Liebig (2023).
 However, since the rank of a matrix is defined as the number of nonzero singular values, it is a discrete function and thus not suitable as a measure.
 A viable alternative is to use a continuous relaxation that closely approximates the rank itself.
-Examples of possible continuous approximations of the rank include the numerical rank, the stable rank, and the effective rank [31, 32, 1].
-The stable rank is defined as $\text{StableRank}(X)=\|X\|_{*}^{2}/\|X\|_{F}^{2}$, where $\|X\|_{*}=\sum_{i}\sigma_{i}$ is the nuclear norm.
-The numerical rank is given by $\text{NumRank}(X)=\|X\|_{F}^{2}/\|X\|_{2}^{2}$.
-Finally, given the singular values $\sigma_{1}>\sigma_{2}>\dots>\sigma_{\min\{N,d\}}$ of $X$, the effective rank is defined as
-$\text{Erank}(X)=\exp\left(-\textstyle{\sum_{k}p_{k}\log p_{k}}\right),$
-where $p_{k}=\sigma_{k}/\sum_{i}\sigma_{i}$ is the $k$-th normalized singular value. These rank relaxation measures exhibit similar empirical behaviour as shown in Section 6.
-In practice, measuring oversmoothing in terms of a continuous approximation of the rank is a reasonable approach that helps addressing the limitations of Dirichlet-like measures.
+Examples of possible continuous approximations of the rank include the numerical rank, the stable rank, and the effective rank royEffectiveRankMeasure2007a; rudelsonSamplingLargeMatrices2006; Arora et al.
+(2019), defined as follows
+$\text{StabRank}(X)=\frac{\|X\|_{*}^{2}}{\|X\|_{F}^{2}},\quad\text{NumRank}(X)=\frac{\|X\|_{F}^{2}}{\|X\|_{2}^{2}},\quad\text{Erank}(X)=\exp\left(-\textstyle{\sum_{k}p_{k}\log p_{k}}\right)$
+where $\|X\|_{*}=\sum_{i}\sigma_{i}$ is the nuclear norm, and given the singular values $\sigma_{1}>\sigma_{2}>\dots>\sigma_{\min\{N,d\}}$ of $X$, $p_{k}=\sigma_{k}/\sum_{i}\sigma_{i}$ is defined as the $k$-th normalized singular value.
+These rank relaxation measures exhibit similar empirical behaviour as shown in Section 6.
+In practice, measuring oversmoothing in terms of a continuous approximation of the rank helps to address the limitations of Dirichlet-like measures.
 Specifically, it offers the following advantages: (a) it is scale-invariant, meaning it remains informative even when the feature matrix converges to zero or explodes to infinity; (b) it does not rely on a fixed, predetermined eigenspace but instead captures convergence of the feature matrix toward an arbitrary lower-dimensional subspace; (c) it allows for the detection of oversmoothing in shallow networks without requiring exact convergence to rank one.
 A small value of the effective rank directly implies that the feature representations are low-rank, suggesting a potentially suboptimal network architecture.
-In Figure 1, we present a toy example illustrating that classical oversmoothing metrics fail to correctly capture oversmoothing unless the features are perfectly aligned.
+In Figure˜1, we present a toy example illustrating that classical oversmoothing metrics fail to correctly capture oversmoothing unless the features are perfectly aligned.
 This observation implies that these metrics can quantify oversmoothing only when the rank of the feature matrix converges exactly to one.
@@
 $E_{\mathrm{Proj}}$ can hardly differentiate between # 3 and # 4, and is thus not robust in quantifying oversmoothing.
-To compute $E_{\mathrm{Proj}}$ and $E_{\mathrm{Dir}}$, the first feature was considered in place of $u$ in (1) and (2).
+$E_{\mathrm{Proj}}$ and $E_{\mathrm{Dir}}$ where computed using the first feature in place of $u$ in (1) and (2).
 5.1 Theoretical Analysis of Rank Decay
-In this section, we provide an analytical study proving the decrease of the numerical rank for a broad class of graph neural network architectures under the assumption that the weight matrices are entry-wise nonnegative.
-While this is a somewhat restrictive setting, our result is the first theoretical proof that oversmoothing occurs independently of the weight (and thus feature) magnitude.
+In this section, we provide an analytical study proving the decrease of the numerical rank for a broad class of graph neural network architectures under the assumption of linear models or nonlinear models with weight matrices that are entry-wise nonnegative.
+Our results rigorously show that in these settings, oversmoothing can occur independently of the weight (and thus feature) magnitude and shed light on the possible causes of rank decay.
 We begin with several useful observations.
@@
 Thus, we obtain the following bound:
-$\text{NumRank}(X)=\frac{\|\mathcal{P}X+(I-\mathcal{P})X\|_{F}^{2}}{\|X\|_{2}^{2}}=\frac{\|\mathcal{P}X\|_{F}^{2}+\|(I-\mathcal{P})X\|_{F}^{2}}{\|X\|_{2}^{2}}\leq 1+\frac{\|(I-\mathcal{P})X\|_{F}^{2}}{\|X\|_{2}^{2}},$ (4)
-The above inequality, together with Theorem 4.1, allows us to establish the convergence of the numerical rank for linear networks.
+$\text{NumRank}(X)=\frac{\|\mathcal{P}X+(I-\mathcal{P})X\|_{F}^{2}}{\|X\|_{2}^{2}}=\frac{\|\mathcal{P}X\|_{F}^{2}+\|(I-\mathcal{P})X\|_{F}^{2}}{\|X\|_{2}^{2}}\leq 1+\frac{\|(I-\mathcal{P})X\|_{F}^{2}}{\|X\|_{2}^{2}}.$ (4)
+The above inequality, together with Theorem˜4.1, allows us to establish the convergence of the numerical rank for linear networks.
 The Linear Case
@@
 This occurs whenever the features grow faster in the direction of the dominant eigenvector than in any other direction.
-As established in Theorem 4.1, this is almost surely the case in linear GNNs.
+As established in Theorem˜4.1, this is almost surely the case in linear GNNs.
 As a direct consequence, we obtain the following result:
 Theorem 5.1.
-Let $X^{(l+1)}=AX^{(l)}W^{(l)}$ be a linear GCN. Under the same assumptions as in Theorem 4.1, the following identity holds almost surely:
+Let $X^{(l+1)}=AX^{(l)}W^{(l)}$ be a linear GCN. Under the same assumptions as in Theorem˜4.1, the following identity holds almost surely:
 $\lim_{l\rightarrow\infty}\mathrm{NumRank}(X^{(l)})=1.$
 Extending the result above to general GNNs with nonlinear activation functions is highly nontrivial.
-In the next subsection, we present our main theoretical result, which generalizes this analysis to a broader class of GNNs under the assumption of nonnegative weights.
-5.1.1 The Nonlinear Case
-To study the case of networks with nonlinear activations, we make use of tools from the nonlinear Perron-Frobenius theory; we refer to [18, 12] and the reference therein for further details.
-Assume all the intermediate features of the network to be in the positive open cone $\mathcal{K}:=\mathbb{R}_{+}^{N}=\{x\in\mathbb{R}^{N}\>|\>x_{i}>0\;\forall i=1,\dots,N\}$.
-On $\mathcal{K}$, it is possible to introduce the partial ordering
-$x\leq_{\mathcal{K}}y\>(x\ll_{\mathcal{K}}y)\quad\text{iff}\quad y-x\in\bar{\mathcal{K}}\>(y-x\in\mathcal{K})$
-where $\bar{\mathcal{K}}$ denotes the closed cone of nonnegative vectors.
-Given two points $x,y\in\mathcal{K}$ we write $M(x/y)=\max_{i}x_{i}/y_{i}$ and $m(x/y)=\min_{i}x_{i}/y_{i}$.
-Then, we can define the following Hilbert distance between any two points $x,y\in\mathcal{K}$, $d_{H}(x,y)=\log(M(x/y)/m(x/y)).$ Note that $d_{H}$ is not a distance on $\mathcal{K}$, indeed $d_{H}(\alpha x,\beta y)=d_{H}(x,y)$ for any $x,y\in\mathcal{K}$ and $\alpha,\beta>0$.
-However it is a distance up to scaling; that is, it becomes a concrete distance whenever we restrict ourselves to a slice of the cone.
-It is well-known that any matrix $A$ such that $Ax\in\mathcal{K}$ for any $x\in\mathcal{K}$ is non-expansive with respect to the Hilbert distance, i.e.
-$d_{H}(Ax,Ay)\leq\beta d_{H}(x,y)\qquad\forall x,y\in\mathcal{K}$ (5)
-with $0\leq\beta\leq 1$.
-The last inequality is a consequence of the fact that $A$ preserves the ordering induced by the cone, i.e.
-if $x\geq_{\mathcal{K}}y$, then $Ax\geq_{\mathcal{K}}Ay$.
-In particular, we recall that (5) always hold for $\beta<1$ when $A$ is entry-wise strictly positive, as positive matrices map the whole $\bar{\mathcal{K}}$ in its interior, i.e.
-$x\geq_{\mathcal{K}}y$ implies $Ax>Ay$.
-In the next definition, we include all the nonnegative matrices that have the same behaviour as the positive matrices but only on a neighbourhood of a positive vector in $\mathcal{K}$.
-To this end, we recall that any irreducible nonnegative matrix $A$ has a positive eigenvector $u$ corresponding to its dominant eigenvalue.
-Definition 5.2.
-A family of nonnegative irreducible matrices $\{A^{(l)}\}$ with the same dominant eigenvector $u>0$ is uniformly contractive with respect to $u$ if for any $C>0$ there exists some $\beta_{C}<1$ such that $d_{H}(A^{(l)}x,u)\leq\beta_{C}d_{H}(x,u)$ for all $l$ and $x\in\mathcal{K}$ such that $d_{H}(x,u)\leq C$
-Note that the above definition is significantly weaker than asking for $A$ to be strictly positive. We show this with an illustrative example in Section A.3.
-Consider now first the case of a linear function that can be represented in the form $F(X)=AXW$, where $X$ is a positive $N\times d$ matrix, $A$ and $W$ are nonnegative $N\times N$ and $d\times d$ matrices, respectively.
-Then, under mild assumptions, we can prove that the columns of $F(X)$ are closer to the dominant eigenvector $u$ of $A$ as compared to the columns of $X$.
+However, a simplified setting to study is the one where the weights are nonnegative.
+Indeed, exactly as in the linear case, while generally the rank decreases only on average, considering nonnegative weights yields a monotone decrease.
+The Nonnegative Nonlinear Case
+To study the case of networks with nonlinear activations, we make use of tools from the nonlinear Perron-Frobenius theory; we refer to lemmensNonlinearPerronFrobeniusTheory2012; gautier2013nonlinear and the reference therein for further details.
+We assume all the intermediate features of the network to be in the positive open cone $\mathcal{K}:=\mathbb{R}_{+}^{N}=\{x\in\mathbb{R}^{N}\>|\>x_{i}>0\;\forall i=1,\dots,N\}$.
+Here, we consider the partial ordering $x\leq_{\mathcal{K}}y\>(x\ll_{\mathcal{K}}y)$ if and only if $y-x\geq 0\>(y-x>0)$, where the inequalities have to be understood entrywise.
+Given two points $x,y\in\mathcal{K}$, let
+$d_{H}(x,y)=\log\Big(\max_{i}\frac{x_{i}}{y_{i}}\cdot\max_{i}\frac{y_{i}}{x_{i}}\Big)$ (5)
+denote the Hilbert distance.
+Note that $d_{H}$ is not a distance on $\mathcal{K}$, indeed $d_{H}(\alpha x,\beta y)=d_{H}(x,y)$ for any $x,y\in\mathcal{K}$ and $\alpha,\beta>0$.
+However, it is a distance up to scaling; that is, it becomes a distance whenever we restrict ourselves to a slice of the cone.
+Because of this property, $d_{H}$ is particularly useful for studying the behavior of the rank of the features, which is a scale-invariant function.
+Indeed, the next result shows that, under mild assumptions, nonnegative weights generate layers that are nonexpansive in Hilbert distance.
 Lemma 5.2.
-Let $F(X)=AXW$ with $A$ nonnegative and irreducible with dominant eigenvector $u\in\mathcal{K}$.
-Assume also $X$ to be strictly positive and $W$ nonnegative with $\min_{j}\max_{i}W_{ij}>0$.
-If $A$ is contractive with respect to $u$ and $C=\max_{i}d_{H}(X_{i},u)$ then
-$\max_{i}d_{H}(F(X)_{:,i},u)\leq\beta_{C}\max_{i}d_{H}(X_{:,i},u),$
-where $Y_{:,i}$ denotes the $i$-th column of $Y$, and $\beta_{C}<1$.
-Nonlinear Perron-Frobenius theory extends some of the results about nonnegative matrices to particular classes of nonlinear functions.
-A function $\sigma\in C(\mathcal{K},\mathcal{K})$ is order preserving if given any $x,y\in\mathcal{K}$ with $x\geq_{\mathcal{K}}y$ then $\sigma(x)\geq_{\mathcal{K}}\sigma(y)$.
-Moreover $\sigma$ is subhomogenenous if $\sigma(\lambda x)\leq_{\mathcal{K}}\lambda\sigma(x)$ for all $x\in\mathcal{K}$ and any $\lambda>1$.
-In particular, it is strictly subhomogeneous if $\sigma(\lambda x)\ll_{\mathcal{K}}\lambda\sigma(x)$ for all $x\in\mathcal{K}$ and $\lambda>1$ and homogeneous if $\sigma(\lambda x)=\lambda\sigma(x)$ for all $x\in\mathcal{K}$ and $\lambda\geq 0$
-Subhomogeneity is a useful property and of practical utility.
-In fact, as discussed in  [38, 27], a broad range of activation functions commonly used in deep learning is subhomogeneous on $\mathcal{K}$.
-For this family of activation functions, Lemma 5.3 combined with arguments from nonlinear Perron–Frobenius theory yield the following main result
+Let $A$ be a nonnegative and irreducible matrix with dominant eigenvector $u\in\mathcal{K}$.
+Assume $X$ to be strictly positive, $W$ nonnegative with $\min_{j}\max_{i}W_{ij}>0$, and $\sigma$ a continuous (nonlinear) function that is order preserving, subhomogeneous, and such that $u$ is also an eigenvector of $\sigma$.
+Then
+$\max_{i}d_{H}\left(\sigma\big((AXW)_{:,i}\big),u\right)\leq\beta\max_{i}d_{H}(X_{:,i},u),$ (6)
+with $0\leq\beta\leq 1$. Where $Y_{:,i}$ denotes the $i$-th column of $Y$. In particular, if $A$ is contractive in the Hilbert distance or $\sigma$ is strictly sub-homogeneous, then $\beta<1$.
+In the above statement, order-preserving means that given any $x,y\in\mathcal{K}$ with $x\geq_{\mathcal{K}}y$, it holds $\sigma(x)\geq_{\mathcal{K}}\sigma(y)$, while (strictly) subhomogenenous means that $\sigma(\lambda x)(\ll_{\mathcal{K}})\leq_{\mathcal{K}}\lambda\sigma(x)$ for all $x\in\mathcal{K}$ and any $\lambda>1$.
+We recall that, as discussed in Sittoni and Tudisco (2024); piotrowski2024fixed, a broad range of activation functions commonly used in deep learning is subhomogeneous and order-preserving on $\mathcal{K}$.
+In particular, whenever an activation function is monotone increasing on $\mathbb{R_{+}}$, it is trivially also order-preserving.
+Additionally, as we prove in Proposition˜A.6, if the activation function is homogeneous, e.g.
+LeakyReLU, then any nonnegative vector is an eigenvector of $\sigma$.
+By contrast, if $\sigma$ is strictly subhomogeneous, e.g.
+$\tanh$, then the only strictly positive eigenvector is the entrywise constant one.
+Next, we prove that for neural networks with nonnegative feature representations, the numerical rank goes to 1 as the depth grows to infinity.
 Theorem 5.3.
-Consider a GNN of the form $X^{(l+1)}=\sigma(A^{(l)}X^{(l)}W^{(l)})$ with $X^{(l)}_{:,i}\in\mathcal{K}$ for any $i=1,\dots,d$ and $A^{(l)}$ and $W^{(l)}$ nonnegative, for any $l$.
-Assume also that $u\in\mathcal{K}$ is the dominant eigenvector of all of the matrices $A^{(l)}$ and that it is also an eigenvector of $\sigma$.
-Then, if the matrices $\{A^{(l)}\}$ are uniformly contractive with respect to $u$ and $\min_{j}\max_{i}W^{(l)}_{ij}>0$ for any $l$, it holds
+Consider a GNN of the form $X^{(l+1)}=\sigma(A^{(l)}X^{(l)}W^{(l)})$ with $X^{(l)}_{:,i}\in\mathcal{K}$ for any $i=1,\dots,d$.
+If there exists $u\in\mathcal{K}$ such that $\lim_{l\rightarrow\infty}\max_{i}d_{H}(X_{:,i}^{(l)},u)=0$, then
 $\lim_{l\rightarrow\infty}\mathrm{NumRank}(X^{(l)})=1.$
-We conclude with a formal investigation of the eigenpairs of activation functions that are entry-wise subhomogeneous.
-Let $\sigma=\otimes^{N}\psi$ with $\psi\in C(\mathbb{R},\mathbb{R})$ that is subhomogeneous on $\mathbb{R}_{+}$.
-Then one can easily show that $\sigma$ is itself subhomogeneous on +$\mathbb{R}^{N}_{+}$.
-We have,
-Proposition 5.5.
-Let $\sigma=\otimes^{N}\psi$ with $\psi\in C(\mathbb{R}_{+},\mathbb{R}_{+})$ be order preserving.
-Then: 1) If $\sigma$ is homogeneous, any positive vector is an eigenvector of $\sigma$.
-2) If $\sigma$ is strictly subhomogeneous, the only eigenvector of $\sigma$ in $\mathcal{K}$ is the constant vector.
-As a consequence of the above result, we find that the numerical rank of the features collapses to one both for GCNs with LeakyReLU activation function (the setting of [4]) as well as for GATs with any kind of subhomogeneous activation function (the setting of [45]), as the constant vector is always an eigenvector of an entry-wise nonlinear map.
-For GCNs with any other subhomogeneous activation functions, the convergence of the numerical rank to 1 as stated in Theorem 5.1 may not hold as discussed in Appendix D.
+Theorem˜5.3 requires that the Hilbert distance between the feature representation and a fixed vector $u$ goes to zero.
+Note that this is implied by the relative metrics $E_{\mathrm{Dir}}(X)/\|X\|$ or $E_{\mathrm{Proj}}(X)/\|X\|$ going to zero, but is actually quite weaker.
+Note also that the bound (6) in Lemma˜5.2 directly provides guidance on situations where the hypotheses of Theorem˜5.3 are satisfied.
+We discuss several such situations along with some alternative and possibly weaker assumptions for Theorem˜5.3 in detail in Section˜A.6.
+Finally, we recall that, because of Lemma˜5.2, the last result applies either to GCNs with homogeneous activation function or GATs with any kind of activation function.
+The convergence of the numerical rank to $1$ may not hold, as discussed in Appendix˜D.
 Dataset Model $E_{\mathrm{Dir}}$ $E_{\mathrm{Proj}}$ MAD Erank NumRank Accuracy ratio
@@
 Average correlation -0.7179 0.4036 -0.7571 -0.4504 0.1601 0.9103 0.8374
-Table 1: Correlation between the classification accuracy and the logarithm of metric values on GNNs with LeakyReLU and of depths from 2 to 24 layers, separately trained on different homophilic (Cora, Citeseer, Pubmed), heterophilic (Squirrel, Chameleon, Amazon Ratings) and large-scale (OGB-Arxiv) datasets.
+Table 1: Correlation between the classification accuracy and the logarithm of metric values on GNNs with LeakyReLU and depths ranging from 2 to 24 layers, separately trained on different homophilic (Cora, Citeseer, Pubmed), heterophilic (Squirrel, Chameleon, Amazon Ratings) and large-scale (OGB-Arxiv) datasets.
 For Erank and NumRank, we subtract 1 so that both metrics approach zero.
@@
 Some heterophilic datasets may be more resilient to the increasing network depth, in-line with observations from the literature, e.g.
-[13].
-Additional results on other datasets and activation functions are presented in Appendix E, and with additional network components in Appendix F.
+Guo et al.
+(2023a).
+Additional results on other datasets, activation functions and additional network components are presented in Appendix˜E and F.
 6 Experiments
-Empirical studies on the evolution of oversmoothing measures often use untrained, hundred-layer-deep GNNs [35, 44, 34, 45].
-We emphasize that this is an overly simplified setting.
-In more realistic settings, as the ones considered in this section, a trained GNN may suffer from significant performance degradation after only few-layers, at which stage the convergence patterns of most of the metrics are difficult to observe.
-The experiments that we present in this section validate the robustness of the relaxed rank metrics in quantifying oversmoothing in GNNs against the other metrics.
+The vast majority of empirical studies in the literature that observe the decay of Dirichlet-like energy metrics are conducted over the layers of deep but untrained networks Wang et al.
+(2022); Rusch et al.
+(2022; 2023b); Wu et al.
+(2023); Roth (2024); Wang et al.
+(2025).
+Moreover, the measurements are done by looking at the layers of a single network, rather than different networks of increasing depth.
+We emphasize that this is an overly simplified and unrealistic setting.
+In this section, we perform an extensive numerical investigation on the behaviour of different oversmoothing metrics when measured on networks of different depths $l=2,\dots,L$, trained in isolation at different depths.
+Our analysis shows that GNN suffer from significant performance degradation after only few-layers, at which stage the convergence patterns of Dirichlet-like metrics are difficult to observe, while relaxed rank metric already show a significant decrease, well-correlating with the performance drop.
 In particular, we compare how different oversmoothing metrics behave compared to the classification accuracy, varying the GNN architectures for node classification on real-world graph data.
 In our experiments, we consider the following metrics:
-  • The Dirichlet Energy $E_{\mathrm{Dir}}$ [4, 33] and its variant $E_{\mathrm{Proj}}$ [45]. Both are discussed in Section 3.1, see in particular (1) and (2).
+  • The Dirichlet Energy $E_{\mathrm{Dir}}$ Cai and Wang (2020); Rusch et al.
+(2023a) and its variant $E_{\mathrm{Proj}}$ Wu et al.
+(2023).
+Both are discussed in Section˜3.1, see in particular (1) and (2).
   • Normalized versions of Dirichlet energy and its variant, $E_{\mathrm{Dir}}(X)/\|X\|_{F}^{2}$ and $E_{\mathrm{Proj}}(X)/\|X\|_{F}$.
 Indeed, from our previous discussion, a robust oversmoothing measure should be scale invariant with respect to the features.
-Metrics with global normalization like the ones we consider here have also been proposed in [8, 30, 20].
-  • The Mean Average Distance (MAD) [5]
+Metrics with global normalization like the ones we consider here have also been proposed in Di Giovanni et al.
+(2023); Roth and Liebig (2023); Maskey et al.
+(2023).
+  • The Mean Average Distance (MAD) Chen et al. (2020)
     $\textstyle\text{MAD}(X)=\frac{1}{|\mathcal{E}|}\sum_{(i,j)\in\mathcal{E}}\Big(1-\frac{X_{i,:}^{\top}X_{j,:}}{|X_{i,:}||X_{j,:}|}\Big).$
@@
   • Relaxed rank metrics: We consider the Numerical Rank and Effective Rank.
-Both are discussed in Section 5.
+Both are discussed in Section˜5.
 We point out that from our theoretical investigation, in particular from (4), the numerical rank decays to $1$ faster than the decay of the normalized $E_{\mathrm{Proj}}$ energy to zero.
-This further supports the use of the Numerical Rank as an improved measure of oversmoothing with respect to $E_{\mathrm{Proj}}$.
-Figure 2: Four examples of the metric behaviours computed at the last hidden layer of separately trained GCNs of increasing depths.
-For Erank and Numrank, we measure Erank$(X)-r^{*}_{\mathrm{ER}}$ and $\mathrm{NumRank}(X)-r^{*}_{\mathrm{NR}}$ for some $r^{*}>1$.
-In these particular cases, $r^{*}_{\mathrm{ER}}<1.85$, $r^{*}_{\mathrm{NR}}<1.3$.
-Note that the effective rank and numerical rank of the input features $X^{(0)}$ are about 1084 and 13.6, respectively.
-Additional results are attached in Appendix G.
-In Table 1 and Figure 2, we train GNNs of a fixed hidden dimension equal to 32 on homophilic, heterophilic and large-scale datasets in their default splits.
-We follow the standard setups of GCN and GAT as stated in Sections 2.1 and 2.2, and use homogeneous LeakyReLU (LReLU) as the activation function.
+This further supports the use of the Numerical Rank as an improved measure of oversmoothing.
+[…]
+In Table˜1 and Figure˜2, we train GNNs of a fixed hidden dimension equal to 32 on homophilic, heterophilic and large-scale datasets in their default splits.
+We follow the standard setups of GCN and GAT as stated in Sections˜2.1 and 2.2, and use homogeneous LeakyReLU (LReLU) as the activation function.
 For each configuration, GNNs of eight different depths ranging from 2 to 24 are trained.
@@
 The oversmoothing metrics are computed at the last hidden layer before the output layer.
-In Figure 2 and in Appendix G, we plot the behaviour of the different oversmoothing measures, the norm of the features, and the accuracy of the trained GNNs with increasing depth.
+In Figure˜2 and in Appendix˜G, we plot the behaviour of the different oversmoothing measures, the norm of the features, and the accuracy of the trained GNNs with increasing depth.
 These figures clearly show that the network suffers a significant drop in accuracy, which is not matched by any visible change in standard oversmoothing metrics.
 By contrast, the rank of the feature representations decreases drastically, following closely the behaviour of the network’s accuracy.
-These findings are further sustained by the results shown in Table 1, where we compute the Pearson correlation coefficient between the logarithm of every measure and the classification accuracy of every GNN model.
+These findings are further supported by the results shown in Table˜1, where we compute the Pearson correlation coefficient between the logarithm of every measure and the classification accuracy of every GNN model.
 The use of a logarithmic transformation is based on the understanding that oversmoothing grows exponentially with the length of the network.
-Extensions of these results are provided in Appendices E and F.
-In Appendix D, we perform an asymptotic ablation study on very deep (300-layer) synthetic networks with randomly sampled, untrained weights.
+Extensions of these results are provided in Appendices˜E and F.
+In Appendix˜D, we perform an asymptotic ablation study on very deep (300-layer) synthetic networks with randomly sampled, untrained weights.
 This study serves to validate our theoretical findings on the convergence of relaxed rank metrics and to demonstrate that such untrained settings offer little insight into the ability of existing metrics to quantify oversmoothing in realistic, trained networks.
@@
 We provide extensive experiments to validate the robustness of the effective rank against the classical measures.
-In addition, we have proved theoretically the decay of the rank of the features for message-passing GNNs.
-Acknowledgement
+In addition, we have analysed theoretically the decay of the rank of the features for message-passing GNNs.
+Notably, our work opens new possible research directions that investigate the use of rank-based metrics in novel architectures that can alleviate the oversmoothing phenomenon.
+Indeed, Dirichlet-like metrics have been employed not only to measure oversmoothing but also to design new architectures able to mitigate this phenomenon, e.g.
+by considering residual architectures where the magnitude of the residual is regulated according to the Dirichlet energy of neighboring nodes Rusch et al.
+(2023b); Jin and Zhu (2024).
+In light of our results and the superiority of rank-based metrics in capturing oversmoothing, it is natural to wonder whether the use of rank-based metrics in place of Dirichlet-like ones could better mitigate oversmoothing in similar kind of architectures.
+We consider this as a promising future research line.
+8 Acknowledgement
 KZ was supported by the EPSRC Centre for Doctoral Training in Mathematical Modelling, Analysis and Computation (MAC-MIGS) funded by the UK Engineering and Physical Sciences Research Council (grant EP/S023291/1), Heriot-Watt University and the University of Edinburgh.
 PD and FT are members of the Gruppo Nazionale Calcolo Scientifico - Istituto Nazionale di Alta Matematica (GNCS-INdAM).
-PD was supported by the MUR-PRO3 grant STANDS.
-DJH was supported by a Fellowship from the Leverhulme Trust.
-FT was partially funded by the PRIN-MUR project MOLE, code 2022ZK5ME7, and PRIN-PNRR project FIN4GEO within the European Union’s Next Generation EU framework, Mission 4, Component 2, CUP P2022BNB97.
+PD was supported by the INdAM-GNCS project “NLA4ML—Numerical Linear Algebra Techniques for Machine Learning”.
+FT is partially funded by the PRIN-MUR project MOLE (code 2022ZK5ME7) and the PRIN-PNRR project FIN4GEO (code P2022BNB97).
+DJH was supported by the Advanced Grant “Numerical Analysis for Stable AI” 101198795 from the European Research Council.
 Appendix A Proofs of the main results
-A.1 Proof of Theorem 3.2
+A.1 Equivalence of $E_{\mathrm{Proj}}$ and $E_{\mathrm{Dir}}$
+Lemma A.1.
+Assume the graph $\mathcal{G}$ to be connected and the eigenvector $u$ to be strictly positive $u_{i}>0$ for all $i$ and such that $\|u\|_{2}=1$. Then there exist $C_{1}>0$ and $C_{2}>0$ such that
+$C_{1}E_{\mathrm{Dir}}(X)\leq E_{\mathrm{Proj}}(X)\leq C_{2}E_{\mathrm{Dir}}(X)\qquad\forall X\in\mathbb{R}^{N\times d}.$
+Proof.
+First we show that the Dirichlet energy is equivalent to the modified Dirichlet energy where all the pairs of nodes $(i,j)$ are considered
+$E_{\mathrm{Dir}}(X)=\sum_{(i,j)\in\mathcal{E}}\left\|\frac{X_{i,:}}{u_{i}}-\frac{X_{j,:}}{u_{j}}\right\|^{2}_{2}\leq\sum_{i\in\mathcal{V}}\sum_{j\in\mathcal{V}}\left\|\frac{X_{i,:}}{u_{i}}-\frac{X_{j,:}}{u_{j}}\right\|^{2}_{2}=\tilde{E_{\mathrm{Dir}}}(X).$ (7)
+Second observe that if $i,j\in\mathcal{V}$, since the graph is connected there exists some path $i_{1}=1,i_{2},\dots,i_{n+1}=j$ such that $(i_{k},i_{k+1})\in\mathcal{E}$ for all $k$.
+Thus, using the traingular inequality we get
+$\left\|\frac{X_{i,:}}{u_{i}}-\frac{X_{j,:}}{u_{j}}\right\|^{2}_{2}\leq n\sum_{k=1}^{n}\left\|\frac{X_{i_{k},:}}{u_{i_{k}}}-\frac{X_{i_{k+1},:}}{u_{i_{k+1}}}\right\|^{2}_{2}\leq nE_{\mathrm{Dir}}(X).$ (8)
+repeating the same argument for any pair of nodes $(i,j)$ we observe that for some constant $C>1$
+$\tilde{E_{\mathrm{Dir}}}(X)\leq\widetilde{C}E_{\mathrm{Dir}}(X).$ (9)
+So, to prove the equivalence of $E_{\mathrm{Dir}}$ and $E_{\mathrm{Proj}}$ it is sufficient to prove the equivalence between $E_{\mathrm{Proj}}$ and $\tilde{E_{\mathrm{Dir}}}$.
+If we make explicit the expression of the norms in $\tilde{E_{\mathrm{Dir}}}$ we get
+$\tilde{E_{\mathrm{Dir}}}(X)=$ $\sum_{i\in\mathcal{V}}\sum_{j\in\mathcal{V}}\sum_{k=1}^{d}\left|\frac{X_{i,k}}{u_{i}}-\frac{X_{j,k}}{u_{j}}\right|^{2}=\sum_{k=1}^{d}\sum_{i\in\mathcal{V}}\sum_{j\in\mathcal{V}}\left|\frac{X_{i,k}}{u_{i}}-\frac{X_{j,k}}{u_{j}}\pm u^{T}X_{:,k}\right|^{2}$ (10)
+$\leq\sum_{k=1}^{d}\sum_{i\in\mathcal{V}}\sum_{j\in\mathcal{V}}\frac{2}{u_{i}}\left|X_{i,k}-(u^{T}X_{:,k})u_{i}\right|^{2}+\frac{2}{u_{j}}\left|X_{j,k}-(u^{T}X_{:,k})u_{j}\right|^{2}$
+$\leq\frac{4|\mathcal{V}|}{\min_{i}\{u_{i}\}}\sum_{k=1}^{d}\left\|X_{:,k}-uu^{T}X_{:,k}\right\|^{2}=\frac{4|\mathcal{V}|}{\min_{i}\{u_{i}\}}E_{\mathrm{Proj}}(X).$
+To prove the opposite observe the following
+$E_{\mathrm{Proj}}(X)=$ $\sum_{k=1}^{d}\sum_{i\in\mathcal{V}}u_{i}\left|\frac{X_{i,k}}{u_{i}}-(u^{T}X_{:,k})\right|^{2}=\sum_{k=1}^{d}\sum_{i\in\mathcal{V}}u_{i}\left|\frac{X_{i,k}}{u_{i}}-\left(\sum_{h}u_{h}^{2}\frac{X_{h,k}}{u_{h}}\right)\right|^{2}$ (11)
+$\leq\max_{j}\{u_{j}\}\sum_{k=1}^{d}\sum_{i\in\mathcal{V}}\max_{j}\left|\frac{X_{i,k}}{u_{i}}-\frac{X_{j,k}}{u_{j}}\right|^{2}$
+$\leq\max_{j}\{u_{j}\}\sum_{k=1}^{d}\sum_{i\in\mathcal{V}}\sum_{j\in\mathcal{V}}\left|\frac{X_{i,k}}{u_{i}}-\frac{X_{j,k}}{u_{j}}\right|^{2}=\tilde{E_{\mathrm{Dir}}}(X)$
+where in the first inequality we have used that, since $\|u\|_{2}=1$, $\sum_{h}u_{h}^{2}\frac{X_{h,k}}{u_{h}}$ is a convex combination of $\left\{\frac{X_{h,k}}{u_{h}}\right\}_{h}$.
+In particular, the last inequality concludes the proof.
+A.2 Proof of Theorem˜3.2
 We start proving that
@@
 So the thesis follows from the hypothesis about the product $\Pi_{l=0}^{L-1}\|(I-\mathcal{P})A^{(l)}\|_{2}\|W^{(l)}\|_{2}$.
-A.2 Proof for Theorem 4.1
+A.3 Proof for Theorem˜4.1
 Start by studying the norm of $(I-\mathcal{P})X^{(l)}$.
@@
 $\geq|\lambda_{1}|^{l}\Big(\|v_{1}^{\top}X^{(0)}W^{(0)}\dots W^{(l-1)}\|_{F}-\Big\|u^{\top}\widetilde{M}O\Big({l\choose N}\Big(\frac{\lambda_{2}}{\lambda_{1}}\Big)^{l}\widetilde{M^{\prime}}X^{(0)}W^{(0)}\dots W^{(l-1)}\Big\|_{F}\Big)\geq$
-$\geq|\lambda_{1}|^{l}\|v_{1}^{\top}X^{(0)}W^{(0)}\dots W^{(l-1)}\|_{F}\Big{(}1-O\Big{(}{l\choose N}\Big{|}\frac{\lambda_{2}}{\lambda_{1}}\Big{|}^{l}\frac{\|X^{(0)}W^{(0)}\dots W^{(l-1)}\|_{F}}{\|v_{1}^{\top}X^{(0)}W^{(0)}\dots W^{(l-1)}\|_{F}}\Big{)}$
-Now observe that under the randomness hypothesis from [11] and more generally from the Oseledets ergodic multiplicative theorem, we have that for almost any the limit $w\in\mathbb{R}^{d}$ $\lim_{l\rightarrow\infty}\frac{1}{l}\log\|w^{\top}W^{(0)}\dots W^{(l-1)}\|=c(\nu)$ exists and is equal to the maximal Lyapunov exponent of the system.
+$\geq|\lambda_{1}|^{l}\|v_{1}^{\top}X^{(0)}W^{(0)}\dots W^{(l-1)}\|_{F}\Big(1-O\Big({l\choose N}\Big|\frac{\lambda_{2}}{\lambda_{1}}\Big|^{l}\frac{\|X^{(0)}W^{(0)}\dots W^{(l-1)}\|_{F}}{\|v_{1}^{\top}X^{(0)}W^{(0)}\dots W^{(l-1)}\|_{F}}\Big)\$
+Now observe that under the randomness hypothesis from furstenbergRandomMatrixProducts1983 and more generally from the Oseledets ergodic multiplicative theorem, we have that for almost any the limit $w\in\mathbb{R}^{d}$ $\lim_{l\rightarrow\infty}\frac{1}{l}\log\|w^{\top}W^{(0)}\dots W^{(l-1)}\|=c(\nu)$ exists and is equal to the maximal Lyapunov exponent of the system.
 In particular for any $w$ and $\epsilon>0$ there exists $l_{w,\epsilon}$ sufficiently large such that for any $l>l_{w,\epsilon}$
@@
 i.e.
-$e^{l(c(\nu)-\epsilon)}\leq\|w^{\top}W^{(0)}\dots W^{(l-1)}\|<e^{l(c(\nu)+\epsilon)}\quad\forall l\geq l_{w,\epsilon}$ (14)
-Now take as vector $w$ first the rows of $X^{(0)}$ and then the vector $v_{1}^{\top}X^{(0)}$, than almost surely for any $\epsilon$ there exists $l_{\epsilon}$ such that for any $l>l_{\epsilon}$
+$e^{l(c(\nu)-\epsilon)}\leq\|w^{\top}W^{(0)}\dots W^{(l-1)}\|<e^{l(c(\nu)+\epsilon)}\quad\forall l\geq l_{w,\epsilon}.$ (20)
+Now take as vector $w$ first the rows of $X^{(0)}$ and then the vector $v_{1}^{\top}X^{(0)}$, then almost surely for any $\epsilon$ there exists $l_{\epsilon}$ such that for any $l>l_{\epsilon}$
 $e^{l(c(\nu)-\epsilon)}\leq\|w^{\top}W^{(0)}\dots W^{(l-1)}\|<e^{l(c(\nu)+\epsilon)},$ (21)
@@
 In particular for any $\epsilon$, there exists $l$ sufficiently large such that
-($\left({l\choose N}\Big{|}\frac{\lambda_{2}}{\lambda_{1}}\Big{|}^{l}\frac{\|X^{(0)}W^{(0)}\dots W^{(l-1)}\|_{F}}{\|v_{1}^{\top}X^{(0)}W^{(0)}\dots W^{(l-1)}\|_{F}}\right)\leq\left({l\choose N}\Big{|}\frac{\lambda_{2}}{\lambda_{1}}\Big{|}^{l}e^{2l\epsilon}\right)$ (17)
+$\left({l\choose N}\Big|\frac{\lambda_{2}}{\lambda_{1}}\Big|^{l}\frac{\|X^{(0)}W^{(0)}\dots W^{(l-1)}\|_{F}}{\|v_{1}^{\top}X^{(0)}W^{(0)}\dots W^{(l-1)}\|_{F}}\right)\leq\left({l\choose N}\Big|\frac{\lambda_{2}}{\lambda_{1}}\Big|^{l}e^{2l\epsilon}\right)$ (23)
 and thus, since $|\lambda_{2}|<|\lambda_{1}|$ and we can choose $\epsilon$ arbitrarily small, almost surely it has limit equal to zero. In particular we can write
@@
 where we have used the same argument as before to state that the limit is zero.
-A.3 An illustrative example
-Consider the stochastic nonnegative primitive matrix
-$A:=\begin{pmatrix}0&1\\[4.30554pt] 1/2&1/2\end{pmatrix}.$ (19)
-Since $A$ is row stochastic, the dominant eigenvector is given by the constant vector $u=(1,1)$. In particular given a vector $x=(x_{1},x_{2})\in\mathcal{K}$ we have that
-$d_{H}(x,u)=\log\Big{(}\frac{\max_{i=1,2}\{x_{i}\}}{\min_{i=1,2}\{x_{i}\}}\Big{)}=\begin{cases}\log\big{(}x_{1}/x_{2}\big{)}\quad\text{if}\;x_{1}\geq x_{2}\\ \log\big{(}x_{2}/x_{1}\big{)}\quad\text{if}\;x_{2}\geq x_{1}\end{cases}.$ (20)
-On the other hand $Ax=\big{(}x_{2}\,,\,(x_{1}+x_{2})/2\big{)}$, thus
-$d_{H}(Ax,u)=\begin{cases}\log\big{(}(x_{1}+x_{2})/2x_{2}\big{)}\quad\text{if}\;x_{1}\geq x_{2}\\ \log\big{(}2x_{2}/(x_{1}+x_{2})\big{)}\quad\text{if}\;x_{2}\geq x_{1}\end{cases}.$ (21)
-Now observe that the set of vectors $x$ such that $d_{H}(x,u)\leq C$ is given by the points $\{x\,|\;x_{1}\leq x_{2}\leq C^{*}x_{1}\}\cup\{x\,|\;x_{2}\leq x_{1}\leq C^{*}x_{2}\}$ with $C^{*}=e^{C}>1$.
-We want to prove the existence of some $\beta_{C}=1-\epsilon_{C}\in(0,1)$ such that $d_{H}(Ax,u)\leq\beta_{C}d_{H}(x,u)$ for any $x$ with $d_{H}(x,u)\leq C$.
-To this end consider first the case of $x\in\Omega_{1}=\{x\,|\;x_{1}\leq x_{2}\leq C^{*}x_{1}\}$. In this case
-$d_{H}(Ax,u)=\log\Big{(}\frac{2x_{2}}{x_{1}+x_{2}}\Big{)}\leq(1-\epsilon_{C})\log\Big{(}\frac{x_{2}}{x_{1}}\Big{)}=(1-\epsilon_{C})d_{H}(x,u)\qquad\forall x\in\Omega_{1}$ (22)
-$\Longleftrightarrow\quad\epsilon_{C}\log\Big{(}\frac{x_{2}}{x_{1}}\Big{)}\leq\log\Big{(}\frac{x_{2}(x_{1}+x_{2})}{2x_{1}x_{2}}\Big{)}\qquad\forall x\in\Omega_{1}$
-$\Longleftrightarrow\quad\epsilon_{C}\log(t)\leq\log\Big{(}\frac{1+t}{2}\Big{)}\qquad\forall t\in[1,C^{*}]$
-$\Longleftrightarrow t^{\epsilon_{C}}\leq\frac{1+t}{2}\qquad\forall t\in[1,C^{*}]$
-where we have written $t=x_{2}/x_{1}$. Second we consider the case of $x\in\Omega_{2}=\{x\,|\;x_{2}\leq x_{1}\leq C^{*}x_{2}\}$. In this case
-$d_{H}(Ax,u)=\log\Big{(}\frac{x_{1}+x_{2}}{2x_{2}}\Big{)}\leq(1-\epsilon_{C})\log\Big{(}\frac{x_{1}}{x_{2}}\Big{)}=(1-\epsilon_{C})d_{H}(x,u)\qquad\forall x\in\Omega_{2}$ (23)
-$\Longleftrightarrow\quad\epsilon_{C}\log\Big{(}\frac{x_{1}}{x_{2}}\Big{)}\leq\log\Big{(}\frac{2x_{1}x_{2}}{x_{2}(x_{1}+x_{2})}\Big{)}\qquad\forall x\in\Omega_{2}$
-$\Longleftrightarrow\quad\epsilon_{C}\log(t)\leq\log\Big{(}\frac{2t}{t+1}\Big{)}\qquad\forall t\in[1,C^{*}]$
-$\Longleftrightarrow t^{\epsilon_{C}}\leq\frac{2t}{t+1}\qquad\forall t\in[1,C^{*}]$
-where we have written $t=x_{1}/x_{2}$.
-Both the functions $t\rightarrow(1+t)/2$ and $t\rightarrow 2t/(t+1)$ are monotonically increasing for $t\in[1,\infty)$ with derivative $1/2$ in $t=1$, thus for any $C<\infty$ there exists some $10<\epsilon_{C}<$ 1 such that both (22) and (22) are satisfied in the interval $t\in[1,C^{*}]$.
-This shows that $A$ is contractive with respect to $u$ even if $A$ is not strictly positive.
-A.4 proof of Lemma 5.3
-By the contraction properties of the matrix $A$ we know that if $\max_{i}d_{H}\big{(}(AX)_{:,i},u\big{)}\leq C$, then
+A.4 Proof of Lemma˜5.2
+Since $A$ is nonnegative, then from Perron Frobenius theory lemmensNonlinearPerronFrobeniusTheory2012, we know that
 $d_{H}\Big((AX)_{:,i},u\Big)=d_{H}\Big((AX)_{:,i},\lambda_{1}(A)u\Big)=d_{H}\Big((AX)_{:,i},Au\Big)\leq\beta d_{H}\big(X_{:,i},u\big)\qquad\forall i.$ (25)
-for some $\beta<1$, where we have used $\lambda_{1}(A)>0$ and the scaling invariant property of the Hilbert distance.
-Then note that, for any $i$, we can write $F(X)_{:,j}$ as follows
-$\big{(}F(X)\big{)}_{:,j}=\sum_{j}W_{ij}(AX)_{:,j}\,.$ (25)
+for some $\beta\leq 1$, where we have used $\lambda_{1}(A)>0$ and the scaling invariant property of the Hilbert distance. In particular if $A$ is contractive in Hilbert distance $\beta<1$.
+Then note that, for any $i$, we can write $(AXW)_{:,j}$ as follows
+$\big(AXW\big)_{:,j}=\sum_{j}W_{ij}(AX)_{:,j}\,.$ (26)
 Thus we CLAIM that given $x_{1},x_{2},y\in\mathcal{K}$ then
 $d_{H}(x_{1}+x_{2},y)\leq\max\{d_{H}(x_{1},y),d_{H}(x_{2},y)\}.$ (27)
-Observe that if the claim holds we have conlcuded the proof, indeed by induction it can trivially be extended from $2$ to $d$ points yielding
-$d_{H}\Big{(}F(X)_{:,j},u\Big{)}\leq\max_{i}d_{H}\Big{(}W_{ij}(AX)_{:,i},u\Big{)}\leq\max_{j}d_{H}\Big{(}(AX)_{:,j},u\Big{)}\leq\beta\max_{j}d_{H}\Big{(}X_{:,j},u\Big{)},$ (27)
+Observe that if the claim holds, by induction it can trivially be extended from $2$ to $d$ points yielding
+$d_{H}\Big((AXW)_{:,j},u\Big)\leq\max_{i}d_{H}\Big(W_{ij}(AX)_{:,i},u\Big)\leq\max_{j}d_{H}\Big((AX)_{:,j},u\Big)\leq\beta\max_{j}d_{H}\Big(X_{:,j},u\Big),$ (28)
 where we have used the scale-invariance property of the Hilbert distance and the fact that $\max_{i}W_{ij}>0$ for all $j$.
-We only miss to prove the claim. To this end, exploiting the expression of the Hilbert distance we write
+We miss to prove the claim. To this end, exploiting the expression of the Hilbert distance we write
 $d_{H}(x_{1}+x_{2},y)$ $=\log\bigg(\sup_{j}\sup_{i}\frac{(x_{1})_{i}+(x_{2})_{i}}{(y)_{i}}\frac{(y)_{j}}{(x_{1})_{j}+(x_{2})_{j}}\bigg)$ (29)
@@
 concluding the proof.
-A.5 proof of Theorem 5.4
-To prove the theorem we start proving that, a continuous subhomogeneous and order-preserving function $\sigma$ with an eigenvector $u$ in the cone, is not nonexpansive in Hilbert distance with respect to $u$.
+Next we prove that, a continuous subhomogeneous and order-preserving function $\sigma$ with eigenvector $u$ in the cone, is not nonexpansive in Hilbert distance with respect to $u$.
 Formally we claim that
 $d_{H}\big(\sigma(y),u\big)\leq d_{H}\big(y,u\big)\qquad\forall y\in\mathcal{K}.$ (30)
-To prove it let $y\in\mathcal{K}$ and assume w.l.o.g. that $\|y\|_{1}=t>0$ and $\|u\|_{1}=1$, then
+To prove it, let $y\in\mathcal{K}$ and assume w.l.o.g. that $\|y\|_{1}=t>0$ and $\|u\|_{1}=1$, then
 $M(y/tu)=\max_{i=1,\dots,N}\frac{y_{i}}{t(u)_{i})}\geq\frac{\|y\|_{1}}{t\|u\|}=1\qquad m(y/tu)=\min_{i=1,\dots,N}\frac{y_{i}}{t(u)_{i}}\leq\frac{\|y\|_{1}}{t\|u\|_{1}}=1.$ (31)
@@
 Thus we use the subhomogeneity of $\sigma$ and the fact that $u$ is an eigenvector of $\sigma$ to get the following inequalities:
-$m(y/tu)\lambda_{t}tu\leq_{\mathcal{K}}\sigma\big{(}m(y/tu)tu\big{)}\leq_{\mathcal{K}}f(y)\leq_{\mathcal{K}}f\big{(}M(y/tu)tu\big{)}\leq_{\mathcal{K}}M(y/x_{c})\lambda_{t}tu.$ (31)
-In particular we have $m(f(y)/tu)\geq\lambda_{t}m(y/tu)$ and $M(f(y)/tu)\leq\lambda_{t}M(y/tu)$. Finally the last inequalities and the scale invariance property of $d_{H}$ yield the thesis:
+$m(y/tu)\lambda_{t}tu\leq_{\mathcal{K}}\sigma\big(m(y/tu)tu\big)\leq_{\mathcal{K}}f(y)\leq_{\mathcal{K}}f\big(M(y/tu)tu\big)\leq_{\mathcal{K}}M(y/x_{c})\lambda_{t}tu,$ (32)
+where the inequalities are strict if $\sigma$ is strictly subhomogeneous.
+In particular we have $m(f(y)/tu)\geq\lambda_{t}m(y/tu)$ and $M(f(y)/tu)\leq\lambda_{t}M(y/tu)$.
+Finally the last inequalities and the scale invariance property of $d_{H}$ yield the thesis:
 $d_{H}(f(y),u)=d_{H}(f(y),tu)=\log\Big(\frac{M(f(y)/tu)}{m(f(y)/tu)}\Big)\leq\log\Big(\frac{M(y/tu)}{m(y/tu)}\Big)=d_{H}(y,tu)=d_{H}(y,u)\,.$ (33)
-Now using the claim above and Lemma 5.3, we can easily conclude that
-$\lim_{l\rightarrow\infty}\max_{i}d_{H}(X_{:,i}^{(l)},u)=0.$ (33)
-Indeed by the uniform contractivity of $\{A^{(l)}\}$, if $C=\max_{i}d_{H}(X^{(0)}_{:,i},u)$, then there exists some $\beta_{C}<1$ such that
-$\max_{i}d_{H}(X^{(l)}_{:,i},u)\leq\beta_{C}^{l}\max_{i}d_{H}(X^{(0)}_{:,i},u).$ (34)
-Finally to conclude we prove that, as a consequence of (33),
-$\lim_{l\rightarrow\infty}\frac{\|(I-\mathcal{P})X^{(l)}\|_{F}}{\|\mathcal{P}X^{(l)}\|_{F}}=0.$ (35)
-As a consequence of the inequality:
-$1\leq\mathrm{NumRank}(X^{(l)})\leq 1+\frac{\|(I-\mathcal{P})X^{(l)}\|_{F}^{2}}{\|X^{(l)}\|_{2}^{2}}\leq 1+\frac{\|(I-\mathcal{P})X^{(l)}\|_{F}^{2}}{\|\mathcal{P}X^{(l)}\|_{F}^{2}},$ (36)
-we have that (35) finally yields the thesis. To prove (35), we recall from Lemma 2.5.1 in [18] that for any $w$ such that $u^{\top}w=c$
+with the inequality that is strict if $\sigma$ is strictly subhomogeneous.
+Then the thesis of the Lemma follows by applying (30) to (28).
+A.5 proof of Theorem˜5.3
+We will prove that, as a consequence of the hypothesis $\lim_{l\rightarrow\infty}\max_{i}d_{H}(X_{:,i}^{(l)},u)=0$,
+$\lim_{l\rightarrow\infty}\frac{\|(I-\mathcal{P})X^{(l)}\|_{F}}{\|\mathcal{P}X^{(l)}\|_{F}}=0,$ (34)
+where $\mathcal{P}=uu^{T}/\|u\|_{2}^{2}$. Indeed (34) is equivalent to proving that the numerical rank goes to 1 as $l$ goes to $\infty$:
+$1\leq\mathrm{NumRank}(X^{(l)})\leq 1+\frac{\|(I-\mathcal{P})X^{(l)}\|_{F}^{2}}{\|X^{(l)}\|_{2}^{2}}\leq 1+\frac{\|(I-\mathcal{P})X^{(l)}\|_{F}^{2}}{\|\mathcal{P}X^{(l)}\|_{F}^{2}}.$ (35)
+To prove (34), we recall from Lemma 2.5.1 in lemmensNonlinearPerronFrobeniusTheory2012 that for any $w$ such that $u^{\top}w=c$
 $\|w-\mathcal{P}w\|_{u}\leq\|\mathcal{P}w\|_{u}(e^{d_{T}(w,\mathcal{P}w)}-1),$ (36)
-where $d_{T}(x,y)=\log(\max\{M(x/y),m(x/y)^{-1}\})$ and where since the dual cone of +$\mathbb{R}^{n}_{+}$ is +$\mathbb{R}^{n}_{+}$ itself, we are considering the norm induced by $u$ on the cone, i.e.
+where $d_{T}(x,y)=\log(\max\{M(x/y),m(x/y)^{-1}\})$ and where since the dual cone of $\mathbb{R}^{n}_{+}$ is $\mathbb{R}^{n}_{+}$ itself, we are considering the norm induced by $u$ on the cone, i.e.
 $\|x\|_{u}=u^{\top}x$ for any $x$ in the cone.
@@
 $\|(I-\mathcal{P})X^{(l)}\|_{F}^{2}\leq C\|\mathcal{P}X^{(l)}\|_{F}^{2}\big(e^{\max_{i}\{d_{H}(X^{(l)}_{:,i},\mathcal{P}X^{(l)}_{:,i})\}}-1\big)^{2}.$ (39)
-the proof is concluded using (33) and observing that by the scale invariance property of the Hilbert distance $d_{H}(X^{(l)}_{:,i},\mathcal{P}X^{(l)}_{:,i})=d_{H}(X^{(l)}_{:,i},u)$, yielding:
+the proof is concluded using the hypothesis and observing that by the scale invariance property of the Hilbert distance $d_{H}(X^{(l)}_{:,i},\mathcal{P}X^{(l)}_{:,i})=d_{H}(X^{(l)}_{:,i},u)$, yielding:
 $1\leq\big(\mathrm{NumRank}(X^{(l)})\big)\leq 1+\frac{\|(I-\mathcal{P})X^{(l)}\|_{F}^{2}}{\|\mathcal{P}X^{(l)}\|_{F}^{2}}\leq 1+C\big(e^{\max_{i}\{d_{H}(X^{(l)}_{:,i},u)\}}-1\big)^{2}$ (40)
 and concluding the proof.
-A.6 proof of Proposition 5.5
+A.6 Some situations where $\lim_{l\rightarrow\infty}\max_{i}d_{H}(X_{:,i}^{(l)},u)=0$
+In this section we explore three different situations where the Hilbert distance of the features from the dominant eigenvector $u$ of the thea aggregation matrices is guaranteed to converge to $1$.
+1st situation
+A first situation where the Hilbert distance of the features from the dominant eigenvector $u$ converges to zero is the case of all matrices $A^{(l)}$ are contractive.
+Lemma A.2.
+Let $A^{(l)}$ be nonnegative and irreducible matrices with dominant eigenvector $u\in\mathcal{K}$.
+Assume also $X^{(0)}$ to be strictly positive, $W^{(l)}$ nonnegative with $\min_{j}\max_{i}W_{ij}^{(l)}>0$ and $\sigma\in C(\mathbb{R}^{N}_{+},\mathbb{R}^{N}_{+})$ a nonlinear function that is order preserving, subhomogeneous and such that $u$ is also an eigenvector of $\sigma$.
+Any matrix $A^{(l)}$ is known to satisfy $d_{H}(A^{(l)}x,A^{(l)}y)\leq d_{H}(x,y)$ for all $x,y\in\mathcal{K}$, then if $\lim_{k\rightarrow\infty}\prod_{l=1}^{k}\beta_{l}=0$
+$\lim_{l\rightarrow\infty}\max_{i}d_{H}(X_{:,i}^{(l)},u)=0.$ (41)
+Proof.
+The proof is a trivial consequence of Lemma˜5.2. Indeed iterating the result in the thesis we know that
+$\max_{i}d_{H}\left(\sigma\big(X^{(l)}_{:,i}\big),u\right)\leq\prod_{i=1}^{l-1}\beta_{i}\max_{i}d_{H}(X_{:,i}^{(0)},u),$ (42)
+concluding the proof ∎
+In particular we remind that from the Perron-Frobenius theory any strictly positive matrix $A$ is known to be contractive of a parameter $\beta<1$, see lemmensNonlinearPerronFrobeniusTheory2012.
+2nd situation
+A second situation where the Hilbert distance of the features from the dominant eigenvector $u$ converges to zero is the case of a strictly subhomogeneous activation function.
+Lemma A.3.
+Let $A^{(l)}$ be nonnegative and irreducible matrices with dominant eigenvector $u\in\mathcal{K}$.
+Assume also $X^{(0)}$ to be strictly positive, $W^{(l)}$ nonnegative with $\min_{j}\max_{i}W_{ij}^{(l)}>0$ and $\sigma\in C(\mathbb{R}^{N}_{+},\mathbb{R}^{N}_{+})$ a nonlinear function that is order preserving, strongly subhomogeneous and such that $u$ is also an eigenvector of $\sigma$.
+Since $\sigma$ is strongly subhomogeneous, for any $l$ there exists $\beta_{l}<1$ such that $\max_{i}d_{H}(\sigma(A^{(l)}X^{(l)}W^{(l)})_{:,i},u)\leq\beta_{l}d_{H}((A^{(l-1)}X^{(l-1)}W^{(l-1)})_{:,i},u)$, then if $\lim_{k\rightarrow\infty}\prod_{l=1}^{k}\beta_{l}=0$
+$\lim_{l\rightarrow\infty}\max_{i}d_{H}(X_{:,i}^{(l)},u)=0.$ (43)
+Proof.
+The proof is again a trivial consequence of Lemma˜5.2. Indeed iterating the result in the thesis we know that
+$\max_{i}d_{H}\left(\sigma\big(X^{(l)}_{:,i}\big),u\right)\leq\prod_{i=1}^{l-1}\beta_{i}\max_{i}d_{H}(X_{:,i}^{(0)},u),$ (44)
+concluding the proof. ∎
+Note that whenever we have a strongly concave activation function $\sigma$ on $\mathbb{R}_{+}$, e.g. tanh, then it is strongly subhomogeneous and so the result above applies.
+3rd situation
+Here we discuss a third situation, possibly weaker than the previous ones.
+This is essentially an adaptation of the ergodic theorem proved in Nussbaum (1990).
+Consider the cone $\mathcal{K}^{\prime}=\mathbb{R}_{+}^{N\times d}$ and introduce the set $\Pi=\{Y\in\mathcal{K}^{\prime}\text{ s.t.
+}Y_{:,i}=\alpha_{i}u\;\forall i=1,\dots,d\}$.
+Moreover let $\psi\in\mathrm{Int}(\mathcal{K}^{\prime})$ and define $\Sigma_{\psi}=\{Y\in\mathcal{K}^{\prime}\text{ s.t.
+}\psi(Y)=1\}$.
+Finally given $X\in\mathcal{K}^{\prime}$ define
+$\pi(X):=\operatorname*{arg\,min}_{Y\in\Sigma_{\psi}}d_{H}(X,Y).$ (45)
+Definition A.4 (Hypotheses H).
+We say that a GNN $X^{(l+1)}=f^{(l)}(X^{(l)})=\sigma(A^{(l)}X^{(l)}W^{(l)})$ with $X^{(0)}\in\mathcal{K}^{\prime}$ satisfies hypotheses $[H]$ if the following conditions are verified:
+  1. H1
+    $A^{(l)}$ is nonnegative and $\min_{j}\max_{i}W^{(l)}_{ij}>0$ for any $l$.
+  2. H2
+    The function $\sigma$ is subhomogeneous and differentiable in $\mathbb{R}_{+}$.
+  3. H3
+    $u\in\mathcal{K}$ is the dominant eigenvector of all of the matrices $A^{(l)}$ and it is also an eigenvector of $\sigma$.
+  4. H4
+    There exists an integer $p>0$ and a sequence of $dN\times dN$ strictly positive matrices $\{B^{(l)}\}$ such that $\forall X$ with $m\big(X^{(lp)},\pi(X^{(lp)})\big)\pi(X^{(lp)})\leq X\leq M\big(X^{(lp)},\pi(X^{(lp)})\big)\pi(X^{(lp)})$
+    $\partial_{X}g^{(l)}(X)\geq B^{(l)}\quad\forall l\geq 1$
+    where $g^{(l)}(X):=f^{((l+1)p-1)}\circ\dots\circ f^{(lp)}(X)$.
+  5. H5
+    $\forall l\geq 0$ there exists $\eta^{(l)}>0$ s.t. $B^{(l)}\pi(X^{(lp)})\geq\eta^{(l)}g^{(l)}\big(\pi(X^{(lp)})\big)$.
+  6. H6
+    $\lim_{M\rightarrow\infty}\sum_{l=0}^{M}\eta^{(l)}exp\big(-\Delta(B^{(l)})\big)=\infty$ where $\Delta(B)=\sup_{x,y\in\mathcal{K}^{\prime}}d_{H}(Bx,By)<\infty$ is the projective diameter of the matrix $B$.
+Theorem A.5.
+Let $X^{(l+1)}=f^{(l)}(X^{(l)})=\sigma(A^{(l)}X^{(l)}W^{(l)})$, with $X^{(0)}\in\mathcal{K}^{\prime}$, be a GNN satisfying hypotheses $[H]$. Then
+$\lim_{l\rightarrow\infty}\max_{i}\big(d_{H}(X^{(l)}_{:,i},u)\big)=0.$
+Proof.
+We claim that under hypotheses [H]
+$\lim_{l\rightarrow\infty}d_{H}(X^{(l)},\Pi)=0.$ (46)
+As a consequence of this it is easy to note that $\lim_{l\rightarrow\infty}\max_{i}\big(d_{H}(X^{(l)}_{:,i},u)\big)=0$. Indeed it is not difficult to check that
+$exp\Big(d_{H}\big(X^{(l)},\pi\big(X^{(l)}\big)\Big)$ $=\frac{M(X^{(l)},\pi\big(X^{(l)}\big)}{m(X^{(l)},\pi\big(X^{(l)}\big)}=\max_{ji}\max_{hk}\frac{X^{(l)}_{ji}\pi\big(X^{(l)}\big)_{hk}}{X^{(l)}_{hk}\pi\big(X^{(l)}\big)_{ij}}\geq$ (47)
+$\geq\max_{i}\max_{j}\max_{h}\frac{X^{(l)}_{ji}\pi\big(X^{(l)}\big)_{hi}}{X^{(l)}_{hi}\pi\big(X^{(l)}\big)_{ji}}=\max_{i}exp\Big(d_{H}(X^{(l)}_{:,i},u)\Big).$
+where we have used that by definition $\pi\big(X^{(l)}\big)_{:,i}=\alpha_{i}u$ for all $i$, where necessarily $\alpha_{i}>0$ for all $i$.
+Otherwise, since $X^{(l)}\in\mathcal{K}^{\prime}$, we would have $d_{H}\big(X^{(l)},\pi\big(X^{(l)})\big)=\infty$, against the minimality.
+Next we prove the claim.
+The proof is adapted for our scopes from the proof of the weak ergodic theorem 2.1 proved in Nussbaum (1990) for homogeneous mappings.
+To simplify the notation we denote by $X:=X^{(lp)}$, $\pi:=\pi(X^{(lp)})$, $g:=g^{(l)}$, $m:=m(X^{(lp)},\pi(X^{(lp)}))$, $B:=B^{(l)}$, $\eta:=\eta^{(l)}$ and $M:=M(X^{(lp)},\pi(X^{(lp)}))$.
+Then consider $z^{1}(t)=(1-t)m\pi+tX$ and $z^{2}(t)=(1-t)X+tM\pi$.
+Then from hypothesis [H4] we have the following inequalities:
+$g(X)-g(m\pi)=\int_{0}^{1}\partial_{X}g(z^{1}(t))\big(X-m\pi\big)\geq B\big(X-m\pi\big)$ (48)
+$g(M\pi)-g(x)=\int_{0}^{1}\partial_{X}g(z^{2}(t))\big(M\pi-X\big)\geq B\big(M\pi-X\big).$
+In particular, since $g$ is subhomogeneous and by definition of $\pi$ $m\leq 1$ and $M\geq 1$ we have:
+$mg(\pi)+B(X-m\pi)$ $\leq g(m\pi)+B(X-m\pi)\leq g(X)\leq$ (49)
+$\leq g(M\pi)-B(M\pi-X)\leq Mg(\pi)-B(M\pi-X)$
+Since $B(X-m\pi)+B(M\pi-X)=(M-m)B\pi$, we can use Lemma 2.2 in Nussbaum (1990) we know that
+$B(X-m\pi)\geq\gamma(M-m)B\pi\qquad\text{or}\qquad B(M\pi-X)\geq\gamma(M-m)B\pi,$ (50)
+where $\gamma=(1/2)exp(-\Delta(B))$. Without loss of generality assume that $B(X-m\pi)\geq\gamma(M-m)B\pi$ the second case can be handled analogously. Then, since $B(M\pi-X)$, from (51), we have
+$mg(\pi)+\gamma\eta(M-m)g(\pi)$ $\leq mg(\pi)+\gamma(M-m)B\pi\leq g(X)\leq Mg(\pi).$ (51)
+In particular
+$d_{H}(g(X),g(\pi))\leq\log(\frac{M}{m+\eta\gamma(M-m})$ $=\log\Big(\frac{M}{m+\eta\gamma(M-m)}\Big)=$ (52)
+$=\log(M/m)\frac{\log\Big(\frac{M}{m}\frac{1}{1+\eta\gamma(M/m-1)}\Big)}{\log(M/m)}\leq$
+$\leq\log(M/m)(1-\eta\gamma)=(1-\eta\gamma)d_{H}(X,\pi).$
+In the last we have used the following fact: if $\xi_{1}(s)=\log\big(s(1+\eta\gamma(s-1))^{-1}\big)$ and $\xi_{2}(s)=\log(s)$ with $s>1$ then since $\xi_{1,2}(1)=0$ and $1+\eta\gamma(s^{\prime}-1)>1$ for $s^{\prime}>1$, for any $s>1$ there exists some $s^{\prime}>1$ such that
+$\xi_{1}(s)/\xi_{2}(s)=\xi^{\prime}_{1}(s^{\prime})/\xi^{\prime}_{2}(s^{\prime})=(1-\eta\gamma)(1+\eta\gamma(s^{\prime}-1))^{-1}\leq(1-\eta\gamma).$ (53)
+Now note that $d_{H}(X,\pi)=d_{H}(X,\Pi)$ by the minimality of $\pi$ and the fact that $d_{H}(X,\alpha Y)=d_{H}(X,Y)$ for any $\alpha>0$.
+Second, using [H1] and [H3] it is very easy to observe that $g(\pi)\in\Pi$.
+So $d_{H}(g(X),\Pi)\leq d_{H}(g(X),g(\pi))$.
+In conclusion we have proved that
+$d_{H}\big(X^{(lp)},\Pi\big)=d_{H}\big(g^{(l)}(X^{(l)}\big),\Pi)\leq\left(1-\eta^{(l)}\frac{exp\big(-\Delta(B^{(l)}\big)}{2}\right)d_{H}(X^{(l)},\Pi).$ (54)
+In particular iterating, using (47) and recalling that $\max_{i}d_{H}(X^{(l_{1})}_{i},u)\leq\max_{i}d_{H}(X^{(l_{2})}_{i},u)$ if $l_{2}>l_{1}$ (see Lemma˜5.2) we have that for any $L>lp$:
+$\max_{i}d_{H}(X^{(L)}_{i},u)$ $\leq\max_{i}d_{H}(X^{(lp)}_{i},u)\leq d_{H}\big(X^{(lp)}_{i},\Pi\big)\leq$ (55)
+$\leq\prod_{j=0}^{l}\left(1-\eta^{(j)}\frac{exp\big(-\Delta(B^{(j)}\big)}{2}\right)d_{H}(X^{(0)},\Pi).$
+Moreover, we have that
+$\lim_{l\rightarrow\infty}\prod_{j=0}^{l}\left(1-\eta^{(j)}\frac{exp\big(-\Delta(B^{(j)}\big)}{2}\right)$ $=0\iff$ (56)
+$\lim_{l\rightarrow\infty}\sum_{j=0}^{l}-\log\left(1-\eta^{(j)}\frac{exp\big(-\Delta(B^{(j)}\big)}{2}\right)$ $=\infty\iff$
+$\lim_{l\rightarrow\infty}\sum_{j=0}^{l}\left(\eta^{(j)}\frac{exp\big(-\Delta(B^{(j)}\big)}{2}\right)$ $=\infty,$
+which concludes the proof.
+Next, we discuss briefly the hypothesis H4 which is the most technical one.
+In particular, such hypothesis recalls the property of a primitive matrix.
+Actually we note that, if the matrices $A^{(l)}$ are primitive, under few additional mild assumptions, it is possible to prove that the hypothesis H4 is always satisfied.
+Assume that there exists an index $p$ such that:
+  1. 1.
+    The matrices $A^{(l)}$ with $l=0,\dots,p-1$ satisfy $A^{(l)}\geq\alpha_{1}A^{*}$ where $A^{*}$ is the unweighted adjacency matrix of the graph; i.e.
+$A^{*}_{ij}=1$ if the edge $(i,j)$ belongs to the graph, $A^{*}_{ij}=0$ otherwise.
+And assume that $A^{*}$ is primitive of index smaller then $p$, i.e.
+$\big((A^{*})^{p}\big)_{ij}>1$ for all $i,j$
+  2. 2.
+    $\partial_{x}\sigma(X_{ij})\geq\alpha_{2}$ for all $X$ with $X=f^{(l)}\circ\dots\circ f^{(0)}(X^{*})$ varying $l=0,\dots,p-1$ and $X^{*}$ among the matrices that satisfy $m\big(X^{(0)},\pi(X^{(0)}\big)\pi(X^{(0)})\leq X\leq M\big(X^{(0)},\pi(X^{(0)})\big)\pi(X^{(0)})$.
+  3. 3.
+    $\big(W^{(0)}\cdot\dots\cdot W^{(p-1)}\big)_{ij}>\alpha_{3}>0$ for all $i,j$.
+Then for any matrix $X$ satisfying $m\big(X^{(0)},\pi(X^{(0)})\big)\pi(X^{(0)})\leq X\leq M\big(X^{(0)},\pi(X^{(0)})\big)\pi(X^{(0)})$ we have
+$\Big(\partial_{X}g(X)\Big)_{ij}\geq\alpha_{1}^{p}\alpha_{2}^{p}\alpha_{3}\quad\forall ij$
+where $g(X):=f^{(p-1)}\circ\dots\circ f^{(0)}(X)$.
+Indeed, vectorizing the matrix $X$ we can write $\partial_{X}f^{(i)}(X)=\mathrm{diag}\Big(\sigma^{\prime}\big((W^{T}\otimes A)vec(X)\big)\Big)(W^{T}\otimes A)\geq\alpha_{2}(W^{T}\otimes A).$ Then we can apply the chain rule to the function $g$ to get
+$\Big(\partial_{X}g(X^{*})\Big)\geq\alpha_{2}^{p}\big({W^{(p-1)}}^{T}\otimes A^{(p-1)}\big)\dots\big({W^{(0)}}^{T}\otimes A^{(0)}\big).$ (57)
+Finally we use the properties of the matrices $W^{(i)}$ and $A^{(i)}$ and the properties of the Kronecker product to obtain
+$\Big(\partial_{X}g(X^{*})\Big)\geq\alpha_{2}^{p}\big({W^{(p-1)}}^{T}\dots W^{(0)}{}^{T}\big)\otimes\big(A^{(p-1)}\dots A^{(0)}\big)\geq\alpha_{2}^{p}\alpha_{1}^{p}\alpha_{3}\mathds{1}$ (58)
+where $\mathds{1}$ is the matrix with every entry equal to $1$.
+A.7 Eigenpairs of entry-wise subhomogeneous maps
+We conclude with a formal investigation of the eigenpairs of activation functions that are entry-wise subhomogeneous.
+Let $\sigma=\otimes^{N}\psi$ with $\psi\in C(\mathbb{R},\mathbb{R})$ that is subhomogeneous on $\mathbb{R}_{+}$.
+Then one can easily show that $\sigma$ is itself subhomogeneous on $\mathbb{R}^{N}_{+}$.
+We have the following result,
+Proposition A.6.
+Let $\sigma=\otimes^{N}\psi$ with $\psi\in C(\mathbb{R}_{+},\mathbb{R}_{+})$ be order preserving.
+Then: 1) If $\sigma$ is homogeneous, any positive vector is an eigenvector of $\sigma$, 2) If $\sigma$ is strictly subhomogeneous, the only eigenvector of $\sigma$ in $\mathcal{K}$ is the constant vector.
 We start from the homogenous case.
@@
 Although a small effective rank or numerical rank indicates oversmoothing and can be subsequently linked to the underperformance of GNNs, a large effective rank or numerical rank does not necessarily correspond to a good network performance.
-Prior study has suggested some degree of smoothing can be beneficial [16], and as an extreme example, features sampled from a uniform distribution and with randomly assigned labels would almost surely have a large effective rank, but they cannot be classified accurately due to the lack of any underlying pattern.
+Prior study has suggested some degree of smoothing can be beneficial Keriven (2022), and as an extreme example, features sampled from a uniform distribution and with randomly assigned labels would almost surely have a large effective rank, but they cannot be classified accurately due to the lack of any underlying pattern.
 We note that this limitation is not specific to the two relaxed rank measures, as the same argument is directly applicable to all other oversmoothing metrics.
-Consequently, as shown in Table 4, when additional components are used to (partially) alleviate oversmoothing, particularly when residual connections are used, the accuracy ratio may remain large over the layers, and all oversmoothing metrics correlate poorly with the accuracy of GNNs.
-This in-turn suggests these oversmoothing metrics become less informative as the oversmoothing problem is mitigated or alleviated.
+Consequently, as shown in Table˜4, when additional components are used to (partially) alleviate oversmoothing, particularly when residual connections are used, the accuracy ratio may remain large over the layers, and all oversmoothing metrics correlate poorly with the accuracy of GNNs.
+This in turn suggests these oversmoothing metrics become less informative as the oversmoothing problem is mitigated or alleviated.
 Appendix C Computational Complexity Analysis
@@
 The cost of additionally computing effective or numerical ranks is marginal.
-Moreover, we note that computing the metrics on a subsection of a graph can be sufficiently informative, and as a consequence, most oversmoothing metrics studied in this paper can be computed in less than $10~{}ms$ for a graph (or a subsection of it) with less than 2000 nodes.
+Moreover, we note that computing the metrics on a subsection of a graph can be sufficiently informative, and as a consequence, most oversmoothing metrics studied in this paper can be computed in less than 10ms10~ms for a graph (or a subsection of it) with less than 2000 nodes.
 Appendix D Additional Experimental Results with Synthetic Weights
 In this section, we conduct an asymptotic ablation study using randomly sampled (synthetic) untrained weights. The aim of these experiments is twofold:
-  • to demonstrate that similar untrained asymptotic experiments are inherently unrealistic, despite being extensively used in the literature [44, 35, 34, 45, 29, 42], as they fail to reliably capture oversmoothing in shallower GNNs where the performance degradation occurs.
-  • to examine the convergence properties of different oversmoothing metrics with weight size control and to empirically validate Theorems 5.1 and 5.4.
+  • to demonstrate that similar untrained asymptotic experiments are inherently unrealistic, despite being extensively used in the literature Wang et al.
+(2022); Rusch et al.
+(2022; 2023b); Wu et al.
+(2023); Roth (2024); Wang et al.
+(2025), as they fail to reliably capture oversmoothing in shallower GNNs where the performance degradation occurs.
+  • to examine the convergence properties of different oversmoothing metrics with weight size control and to empirically validate Theorems˜5.1 and 5.3.
 We construct a 10-node Barabasi-Albert graph with each node having 32 features.
@@
 The feature initialization $X^{(0)}$ is sampled from $\mathcal{U}(0,1)$, and is iterated over 300 layers.
-In this asymptotic synthetic setting, as presented in Table 2, the normalized $E_{\mathrm{Dir}}$ and $E_{\mathrm{Proj}}$ exhibit decay patterns similar to those of the effective rank and numerical rank, suggesting these metrics are equally sensitive to asymptotic rank collapse.
-However, this behaviour stands in stark contrast to results on trained networks presented in Tables 1 and 3 and Appendix G, where the normalized $E_{\mathrm{Dir}}$ and $E_{\mathrm{Proj}}$ often fail to detect oversmoothing.
-Moreover, these asymptotic results validate Theorems 5.1 and 5.4, showing that the numerical rank converges to one for GCN + LReLU and GAT + any subhomogeneous activation functions.
+In this asymptotic synthetic setting, as presented in Table˜2, the normalized $E_{\mathrm{Dir}}$ and $E_{\mathrm{Proj}}$ exhibit decay patterns similar to those of the effective rank and numerical rank, suggesting these metrics are equally sensitive to asymptotic rank collapse.
+However, this behaviour stands in stark contrast to results on trained networks presented in Tables˜1 and 3 and Appendix˜G, where the normalized $E_{\mathrm{Dir}}$ and $E_{\mathrm{Proj}}$ often fail to detect oversmoothing.
+Moreover, these asymptotic results validate Theorems˜5.1 and 5.3, showing that the numerical rank converges to one for GCN + LReLU and GAT + any subhomogeneous activation functions.
 Without making any additional assumption on the normalization of the adjacency matrix, the effective rank and numerical rank do not generally decay to one when subhomogeneous activation functions, e.g.
@@
 Appendix E Additional Experimental Results on Activation Functions and Different Datasets
-We extend Table 1 to subhomogenous Tanh activation function and a few additional datasets. The experimental setting is consistent with that of Section 6.
+We extend Table˜1 to subhomogenous Tanh activation function and a few additional datasets. The experimental setting is consistent with that of Section˜6.
 Dataset Architecture $E_{\mathrm{Dir}}$ $E_{\mathrm{Proj}}$ MAD Erank NumRank Accuracy ratio
@@
 Appendix F Additional Experimental Results with Different Network Components
-Prior literature indicates that when adding additional components, such as bias [33] or residual terms [36], the Dirichlet energy does not decay.
-Therefore, we compare the correlation coefficients in Table 4 between existing oversmoothing metrics when additional components are added, such as bias, LayerNorm, BatchNorm, PairNorm [46], DropEdge [28, 15] and residual connections [36].
-All experiments follow the setup described in Section 6. In addition, DropEdge has a probability of 0.5 in removing each edge at each layer. The residual connection is implemented as follows
-$X^{(l+1)}=\sigma(AX^{(l)}W_{1}^{(l)})+X^{(0)}W_{2}^{(l)}$
-Table 4 demonstrates that both the effective rank and numerical rank achieve a higher average correlation with the classification accuracy than alternative metrics.
+Prior literature indicates that when adding additional components, such as bias Rusch et al.
+(2023a) or residual terms Scholkemper et al.
+(2024), the Dirichlet energy does not decay.
+Therefore, we compare the correlation coefficients in Table˜4 between existing oversmoothing metrics when additional components are added, such as bias, LayerNorm, BatchNorm, PairNorm Zhao and Akoglu (2019), DropEdge Rong et al.
+(2019); Huang et al.
+(2020) and residual connections Scholkemper et al.
+(2024).
+All experiments follow the setup described in Section˜6. In addition, DropEdge has a probability of 0.5 in removing each edge at each layer. The residual connection is implemented as follows
+$X^{(l+1)}=\sigma(AX^{(l)}W_{1}^{(l)})+X^{(0)}W_{2}^{(l)}.$
+Table˜4 demonstrates that both the effective rank and numerical rank achieve a higher average correlation with the classification accuracy than alternative metrics.
 This finding confirms their superior consistency in detecting oversmoothing across a variety of architectural variants.
@@
 when residual connection is used, all metrics have a poor correlation with the classification accuracy.
-This generic limitation is discussed in Appendix B.
+This generic limitation is discussed in Appendix˜B.
 Architecture $E_{\mathrm{Dir}}$ $E_{\mathrm{Proj}}$ MAD Erank NumRank Accuracy ratio
@@
 Appendix G Metric Behaviour Examples on Cora, Citeseer and Pubmed
-We extend Figure 2 to additional datasets and network components. The experimental setting is consistent with that of Section 6.
+We extend Figure˜2 to additional datasets and network components. The experimental setting is consistent with that of Section˜6.
 GCN + LReLU $r^{*}_{\mathrm{ER}}=1.450100,\quad r^{*}_{\mathrm{NR}}=1.207918$ GCN + Tanh $r^{*}_{\mathrm{ER}}=1.690520,\quad r^{*}_{\mathrm{NR}}=1.007135$